### v2

Abstract
Recently, introducing Tensor Decomposition (TD) techniques into unsupervised feature selection (UFS) has been an emerging research topic. A tensor structure is beneficial for mining the relations between different modes and helps relieve the computation burden. However, while existing methods exploit TD to preserve the data tensor structure, they do not consider the influence of data orientation and thus have difficulty in handling orientation-specific data such as time series. To solve the above problem, we utilize the orientation-dependent tensor-tensor product from Tensor Singular Value Decomposition based on $\star_{\mathbf{M}}$-product (T-SVDM) and extend the one-dimensional Sparse Principal Component Analysis (SPCA) to a tensor form. The proposed sparse tensor PCA model can constrain sparsity at the specified mode and yield sparse tensor principal components, enhancing flexibility and accuracy in learning feature relations. To ensure fast convergence and a flexible description of feature correlation, we develop a convex version specially designed for general UFS tasks and propose an efficient slice-by-slice algorithm that performs dual optimization in the transform domain. Experimental results on real-world datasets demonstrate the effectiveness and remarkable computational efficiency of the proposed method for tensor data of diverse structures over the state-of-the-arts. With a proper combination of data orientation and transform domain, our method is promising for various applications. The codes related to our proposed methods and the experiments are available at https://github.com/zjj20212035/STPCA.git.
keywords:
[…]
1 Introduction
Feature Selection (FS) is a technique of selecting as few features (variables) as possible to meet the demands of a certain task [1, 2]. It is usually employed as a pre-processing algorithm to relieve the computational burden and enhance interpretability by offering only the necessary features. According to the amount of supervised information during training, FS methods can be classified as supervised [3], semi-supervised [4], and unsupervised [5]. Among them, unsupervised feature selection (UFS) is more challenging due to the unavailability of labels. To overcome the difficulty, most UFS methods adopt learning methods including but not limited to principal component analysis (PCA) [38], spectral clustering [40], and graph-based learning [18]. In this paper, we focus on PCA-based UFS methods.
In most UFS methods, data samples are vectorized, resulting in the curse of dimensionality [6] as well as data structure loss when dealing with tensors. As a natural extension of matrix to store multi-way data samples, tensor is now a common processed object in various communities such as image processing [7], signal processing [8], and economical analysis[9]. A $n$-th order tensor comprises $n$ modes, each containing several dimensions. The tensor structure contains rich information on the relations of modes and dimensions, which is beneficial for analyzing the features. In this paper, we focus on leveraging the tensor structure to reduce computational complexity without compromising UFS performance.
To utilize the structure information, tensor-based UFS methods apply all kinds of Tensor Decomposition (TD) techniques[18, 19, 20]. Tensor decomposition aims to factorize a tensor into a product of several subtensors, which has been considerably studied for decades since the famous Tucker decomposition [21] was proposed in 1964. As different definitions for tensor rank are established, various TD methods have been put forward, such as Higher-order Singular Value Decomposition (HOSVD) [12], CANDECOMP/PARAFAC decomposition (CPD) [22, 23], and Tensor Singular Value Decomposition (T-SVD)[24]. One purpose of TD is to mine the relations between modes and dimensions, which are naturally related to UFS. TD methods preserve the tensor structure of UFS, while UFS helps validate the accuracy and interpretability of the relationships learned by the TD method. By minimizing the reconstruction error between a constructed tensor and its decomposition, tensor-based UFS methods successfully preserve the data structure information. However, most of these methods do not develop sparse TD and thus use the classifier matrix to perform UFS, which is not beneficial for interpretability and utilization of the learned element relations. To address this problem, we start by finding the compact representations of the factor matrices. As is widely known, TD methods are also viewed as ways to achieve Tensor Principal Component Analysis (TPCA) [12, 25]. The principal components have the greatest variance among data samples on a global manifold. Their directions describe the feature necessity during the projection onto the manifold. Therefore, the key to selecting interpretable and discriminative features could be hidden in TPCA.
In terms of performing linear projection, the current TPCA methods can be roughly divided into two types: Tuckers-based TPCA [12, 26, 27] and T-SVDM-based TPCA [28, 29, 30]. Tuckers-based TPCA applies direction-unfolding products defined by Tucker decomposition and its variations (including CPD, which can be viewed as a special case of Tucker’s), while T-SVDM-based TPCA is based on the $\star_{\mathbf{M}}$-product defined by T-SVD and its generalization T-SVDM [31]. The main differences between the above two types of methods lie in the following two aspects: 1) Tucker-based methods conduct tensor-matrix products through tensor matricization in the original domain, using factor matrices, while T-SVDM-based methods perform tensor-tensor products in the transform domain with a factor tensor. By applying factor tensors, T-SVDM-based methods obtain significantly more weights assigned for elements in the data tensor, describing feature relations in a more delicate way. 2) Tucker-based methods obtain identical results regardless of data organization, while T-SVDM-based methods can achieve various decompositions by applying different transform domains and changing tensor orientation. The reorientation works as prior knowledge of mode relations and guides the T-SVDM-based methods to further learn the feature relations in a proper direction, especially in some orientation-specific scenarios such as time series. Moreover, the transform operation provides T-SVDM-based methods with greater flexibility, allowing their application to various scenarios. In a word, T-SVDM-based methods have a unique advantage in handling complex scenarios. We are interested in introducing T-SVDM-based TPCA into UFS. To achieve this goal, an extension to a sparse framework is required to ensure the preservation of only the necessary feature relations. To the best of our knowledge, this framework is still absent.
This paper develops a novel sparse T-SVDM-based TPCA framework and deduces a convex version for UFS. We explore a way of leveraging the tensor structure to reduce computational complexity while preserving satisfying UFS performance. Our contributions are summarized as follows:
  1. [•]
  2. 1.
    We put forward a novel Sparse Tensor PCA model based on $\star_{\mathbf{M}}$-Product (STPCA-MP) that extends the principle of one-dimensional Sparse PCA (SPCA) to a tensor form by applying a tensor decomposition technique called T-SVDM. By considering the orientation of the data tensor, the proposed model can better utilize the data organization and learn the feature correlations.
  3. 2.
    We deduce a convex version of STPCA-MP specially designed for unsupervised feature selection tasks by merging the encoder tensor and decoder tensor into the unified sparse reconstruction factor tensor. Our method not only preserves the global information by minimizing the reconstruction error but also demonstrates consistency between TD and UFS by utilizing the reconstruction tensor to perform UFS, fully leveraging learned latent feature relations and gaining more interpretability. We prove that the global optimal solution falls onto the Hermitian Positive Semidefinite Cone.
  4. 3.
    We design an optimization algorithm that exploits the transform domain to solve a dual problem in a slice-by-slice manner. By applying the HPSD projector, we can efficiently solve each subproblem and end up with a score map for feature selection. The convergence of the proposed algorithm is proven theoretically and practically. Compared to non-tensor-based methods, our proposed method reaches a lower computational complexity.
  5. 4.
    Real-world datasets are used to validate the effectiveness and efficiency of the proposed method, especially towards orientation-specific data like time series.
Figure 1: Overall framework of our work
2 Related Work
In this section, we introduce notations used throughout the paper, and research revolving around tensor decomposition, tensor PCA, and UFS.
2.1 Notations
2.1.1 Basic Symbols
In most cases, we use capital letters (e.g. $A$) to denote sets, except the common ones like integer set $\mathbb{Z}$, the real-number set $\mathbb{R}$ and the complex-number set $\mathbb{C}$. In particular, $H^{d}_{+}$ stands for the HPSD Cone for $d\times{d}$ matrices, and $\Phi$ is the empty set. For an integer $n$, $[n]$ represents the set $\left[1,2,3,...,n\right]$. #$\#(A)$ stands for the number of elements in the set $A$. A set with the superscript "o" has ordered elements; for example, $A^{o}=\{1,2,3\}\neq B^{o}=\{1,3,2\}$.
We use lowercase letters (e.g. $a$) to denote scalars, boldface lowercase letters (e.g. $\mathbf{a}$) to denote vectors, boldface capital letters (e.g. $\mathbf{A}$) and Euler script letters (e.g. $\mathcal{A}$) to denote tensors. For a $n^{th}$-order tensor $\mathcal{A}\in\mathbb{C}^{d_{1}\times{d_{2}}\times{\cdots}\times{d_{n}}}$, its $(i_{1},i_{2},\cdots,i_{n})$-th element is denoted by $\mathcal{A}_{i_{1}i_{2}\cdots i_{n}}$. For a matrix, the corresponding element is denoted by $a_{i_{1}i_{2}}$. For a matrix $\mathbf{A}$, its row vectors are denoted by $\mathbf{A}_{j,:}$ or boldface lowercase letters $\mathbf{a}^{j}$. The column vectors are denoted by $\mathbf{A}_{:,j}$ or boldface lowercase letters $\mathbf{a}_{j}$.
There are two types of common sections for third-order tensors: fibers and slices. The vector along the $n$-th mode is referred to as an $n$-mode fiber. A third-order tensor $\mathcal{A}\in\mathbb{C}^{d_{1}\times{d_{2}}\times{d_{3}}}$ has column (1-mode) fibers $\mathcal{A}_{:i_{2}i_{3}}$, row (2-mode) fibers $\mathcal{A}_{i_{1}:i_{3}}$, and tube (3-mode) fibers $\mathcal{A}_{i_{1}i_{2}:}$. A slice is a matrix obtained by fixing all but two indices in a tensor. For a third-order tensor, slices can be divided into frontal slices $\mathcal{A}_{::i_{3}}$, lateral slices $\mathcal{A}_{:i_{2}:}$ and horizontal slices $\mathcal{A}_{i_{1}::}$.
The transpose, conjugation, and conjugate transpose of a matrix are denoted by $\mathbf{A}^{T}$, $\mathbf{A}^{*}$, and $\mathbf{A}^{H}$, respectively. We denote the trace of a square matrix by ${\rm Tr}(\mathbf{A})$. We denote Frobenius norm, nuclear norm and $\ell_{2,1}$ norm by ${\|\mathcal{A}\|}_{F}$,${\|\mathbf{A}\|}_{*}$, and ${\|\mathbf{A}\|}_{2,1}$ respectively. The nuclear norm of a matrix is defined as the sum of its singular values. The $\ell_{2,1}$-norm is obtained by calculating the $\ell_{2}$-norm of each column vector in a matrix and summing the results. The Kronecker product is denoted by $\bigotimes$.
2.1.2 Tensor Operations
In this section, we introduce several tensor operations.
a) Tensor unfolding and tensor-matrix product
We use $n$-mode unfolding to perform tensor matricization. A $m$-th order tensor $\mathcal{A}^{d_{1}\times{d_{2}}\dots\times{d_{m}}}$ can be transformed into $\mathbf{X}_{(n)}$ along the $n$-th mode where
[…]
Using the unfolding technique, the tensor-matrix product is defined as
$\mathcal{Y}=\mathcal{X}\times_{n}\mathbf{U}\iff\mathbf{Y}_{(n)}=\mathbf{U}\mathbf{X}_{(n)},$ (2)
which is also known as the $n$-mode product.
b) $\star_{\mathbf{M}}$-product
In the T-SVDM framework, third-order tensor products are achieved through the $\star_{\mathbf{M}}$-product. Given a third-order tensor $\mathcal{A}\in\mathbb{C}^{d\times{m}\times{n}}$, we first denote a tensor in the transform domain specified by an invertible matrix $\mathbf{M}\in\mathbb{C}^{n\times{n}}$:
$\hat{\mathcal{A}}:=\mathcal{A}\times_{3}\mathbf{M}.$ (3)
Then, given another tensor $\mathcal{B}\in\mathbb{C}^{m\times{l}\times{n}}$, $\mathcal{A}\star_{\mathbf{M}}\mathcal{B}=\mathcal{C}$ is defined as
$\hat{\mathcal{C}}_{::i}$ $=\hat{\mathcal{A}}_{::i}\hat{\mathcal{B}}_{::i},\quad i=1,2,\cdots,n,$ (4)
$\mathcal{C}$ $=\hat{\mathcal{C}}\times_{3}\mathbf{M}^{-1}\in{\mathbb{C}^{d\times{l}\times{n}}}.$
Hereafter, we use $\hat{\mathcal{A}}$ to denote the result of $\mathcal{A}\times_{3}\mathbf{M}$. The $\star_{\mathbf{M}}$-product is a slice-by-slice operation and is thus orientation-dependent. Different orientations of the tensor will lead to different results.
The identity tensor $\mathcal{I}$ under $\star_{\mathbf{M}}$ satisfies the following condition:
$\mathcal{A}\star_{\mathbf{M}}\mathcal{I}=\mathcal{I}\star_{\mathbf{M}}\mathcal{A}^{H}=\mathcal{A}.$ (5)
Each frontal slice of $\hat{\mathcal{I}}$ is an identity matrix $\mathbf{I}$.
The conjugate transpose of $\mathcal{A}\in\mathbb{C}^{d\times{m}\times{n}}$ under $\star_{\mathbf{M}}$ is defined as:
$\left(\hat{\mathcal{A}}^{H}\right)_{::i}=\left(\hat{\mathcal{A}}_{::i}\right)^{H},\quad i=1,\cdots,n,$ (6)
which makes sure that $\mathcal{A}^{H}\star_{\mathbf{M}}\mathcal{B}^{H}=\left(\mathcal{B}\star_{\mathbf{M}}\mathcal{A}\right)^{H}$. We can further define a $\star_{\mathbf{M}}$-orthogonal tensor $\mathcal{A}\in{\mathbb{C}^{m\times{m}\times{n}}}$ if
$\mathcal{A}^{H}\star_{\mathbf{M}}\mathcal{A}=\mathcal{A}\star_{\mathbf{M}}\mathcal{A}^{H}=\mathcal{I}.$ (7)
2.2 T-SVDM
Using the $\star_{\mathbf{M}}$ product, T-SVDM [31] decomposes a third-order tensor $\mathcal{A}\in\mathbb{C}^{d_{1}\times{d_{2}}\times{d_{3}}}$ into:
$\mathcal{A}=\mathcal{U}\star_{\mathbf{M}}\mathcal{S}\star_{\mathbf{M}}\mathcal{V}^{H},$ (8)
where $\mathcal{U}\in\mathbb{C}^{d_{1}\times d_{1}\times d_{3}}$, $\mathcal{V}\in\mathbb{C}^{d_{2}\times d_{2}\times d_{3}}$ are $\star_{\mathbf{M}}-unitary$, and $\mathcal{S}\in\mathbb{C}^{d_{1}\times{d_{2}}\times{d_{3}}}$ is a f-diagonal tensor whose frontal slices are diagonal. The decomposition is called T-SVD [24] when $\mathbf{M}$ is the discrete Fourier transform (DFT) matrix and T-SVDM [31] in a usual case.
[28] was the first to establish the relationship between T-SVD and tensor PCA. It applies T-SVD to compress the human-face images with a pre-rotation of the data tensor. [31] further extends this idea to T-SVDM. [29] highlights the importance of leveraging the orientation-dependence of T-SVDs to extract information from different directions. Inspired by these works, we derive a sparse model for tensors combining T-SVDM and SPCA.
2.3 PCA and Sparse PCA (SPCA)
Principal Component Analysis (PCA) [47] is a well-known unsupervised feature extraction method. It aims to find an orthogonal linear transformation matrix that projects the original data vectors into a lower-dimensional manifold, maximizing the variance of the projected samples. The optimization problem of PCA can be formulated as the following self-contained regression form:
$\min\limits_{\mathbf{U}^{T}\mathbf{U}=\mathbf{I}_{k}}{\|{\mathbf{X}-\mathbf{U}\mathbf{U}^{T}\mathbf{X}}\|}^{2}_{F},$ (9)
where $\mathbf{X}=\left[\mathbf{x}_{1},\mathbf{x}_{2},\mathbf{x}_{3},\cdots,\mathbf{x}_{n}\right]\in\mathbb{R}^{d\times{n}}$ is the data matrix, with each column representing a sample. $\mathbf{U}\in\mathbb{R}^{d\times{k}}$ ($k\ll d$) is an orthogonal transformation matrix.
To obtain sparse weights and improve interpretability, Sparse PCA (SPCA) [32] was proposed by adding ridge and lasso constraints to PCA:
$\min\limits_{\mathbf{U},\mathbf{Q}}$ ${\|{\mathbf{X}-\mathbf{U}\mathbf{Q}^{T}\mathbf{X}}\|}^{2}_{F}+\alpha\sum\limits_{j=1}^{k}{\|{\mathbf{q}_{j}}\|}^{2}_{2}+\beta\sum\limits_{j=1}^{k}{\|{\mathbf{q}_{j}}\|}_{1}$ (10)
$s.t.$ ${\mathbf{U}^{T}\mathbf{U}=\mathbf{I}_{k}},$
where $\alpha,\beta>0$ are regularization parameters.
The self-contained regression term in SPCA incorporates an encoder-decoder architecture, selectively transforming a sparse subset of input features into encoded latent representations and reconstructing them through an orthogonal decoder, as Figure 2 shows. This approach gains flexibility by relaxing the transformation matrix $\mathbf{Q}$. When $\mathbf{Q}=\mathbf{U}$, SPCA reduces to PCA.
Figure 2: The encoder-decoder principle of one-dimensional SPCA
By minimizing reconstruction errors, SPCA preserves the global data structure in an unsupervised manner, maintaining essential geometric relationships within the data manifold. The lasso constraint enables SPCA to assign weights only to the necessary features. The orthogonality constraint ensures that the latent features are complementary and reduces redundancy. Therefore, SPCA is well-suited for UFS tasks, particularly in balancing the clustering performance and feature interpretability.
2.4 Tensor-based unsupervised feature selection
Tensor-based unsupervised feature selection remains an emerging research area. Most of the existing methods rely on domain knowledge and are limited to certain application scenarios. For example, [20] proposes an image-transformation-based feature selection method, aiming at vehicle detection. [34] constructs a multi-omics data tensor and exploits feature selection criterion designed using the knowledge in bioinformatics. Recently, some methods designed for general cases have emerged. GRLTR [35] combines graph regularization with tensor tubal rank constraint to simultaneously preserve global and local information. However, it still requires vectorization. CPUFS [18] derives a model based on CPD and effectively preserves the tensor structure by minimizing the reconstruction error. It represents a promising approach to tensor-based UFS. However, CPUFS does not derive sparse factor matrices. Instead, it applies spectral clustering and utilizes a sparse linear classifier matrix to perform UFS, leading to insufficient utilization of the information in factor matrices and reduced interpretability of selected features. Additionally, the constructed graph in CPUFS needs the number of classes as prior knowledge and can be unreliable due to noise features.
Recently, developing sparse PCA models for UFS appears to be a new trend [36, 37, 38, 39]. PCA-like methods do not rely on the construction of graphs. By global manifold learning, they can preserve the data structure and achieve superior average performance. What’s more, they utilize the factor matrix that maximizes the variance of data samples in the projected space, whose sparse weights demonstrate a clearer relevance of the original features and have better interpretability. Unfortunately, existing PCA-like methods are non-tensor-based. Motivated by the above considerations, this study develops a sparse TPCA-based methodology tailored for UFS tasks.
3 Methodology
In this section, we extend the framework of SPCA to tensor form by utilizing techniques from T-SVDM and derive a model called Sparse Tensor PCA based on $\star_{\mathbf{M}}$ product (STPCA-MP). Then, we deduce a convex version of STPCA-MP for UFS tasks by merging the encoder and decoder tensors into a single sparse reconstruction tensor. We propose an efficient slice-by-slice optimization algorithm that individually solves each dual subproblem in the transform domain. Eventually, by considering both the transform matrix $\mathbf{M}$ and the learned reconstruction tensor, we compute a feature score map and perform feature selection. The overall framework is illustrated in Figure 1.
3.1 Model formulation of STPCA-MP
We formulate the model in the third-order tensor scenario for the simplicity of deduction and algorithm design. The principle can be extended to high-order scenarios in a recursive fashion. See Section 4 for details.
Given a data tensor $\mathcal{X}\in\mathbb{C}^{d_{1}\times d_{2}\times n}$ containing $n$ samples of size $d_{1}\times d_{2}$, we can reorient it with an order set $D^{o}$ such that the second mode represents the sample dimension. We denote the rotated data tensor by $\mathcal{X}_{D^{o}}\in\mathbb{C}^{q\times n\times p}$. Based on (8), the row dimensionality of each frontal slice in $\mathcal{X}_{D^{o}}$ can be reduced with $\mathcal{U}^{H}\in\mathbb{C}^{k\times q\times p}(k<d)$, followed by reconstruction with $\mathcal{U}$:
$\mathcal{Y}$ $=\mathcal{U}^{H}\star_{\mathbf{M}}\mathcal{X}_{D^{o}},$ (11)
$\tilde{\mathcal{X}}_{D^{o}}$ $=\mathcal{U}\star_{\mathbf{M}}\mathcal{Y},$
where $\mathcal{Y}\in\mathbb{C}^{k\times n\times p}$, and $\tilde{\mathcal{X}}_{D^{o}}\in\mathbb{C}^{q\times n\times p}$ is the reconstructed tensor.
Intuitively, we can first formulate a tensor PCA model following the idea of one-dimensional PCA in (9):
$\underset{\mathcal{U}}{\min}$ ${\|\mathcal{X}_{D^{o}}-\mathcal{U}\star_{\mathbf{M}}\mathcal{U}^{H}\star_{\mathbf{M}}\mathcal{X}_{D^{o}}\|^{2}_{F}}$ (12)
$s.t.$ $\quad\mathcal{U}^{H}\star_{\mathbf{M}}\mathcal{U}=\mathcal{I}.$
Then, inspired by the encoder-decoder structure of SPCA, we use another tensor $\mathcal{Q}\in\mathbb{C}^{d\times k\times p}$ to perform feature extraction instead of $\mathcal{U}$. By doing so, we obtain a more flexible reconstruction framework that yields a better latent feature representation:
$\underset{\mathcal{U}}{\min}$ ${\|\mathcal{X}_{D^{o}}-\mathcal{U}\star_{\mathbf{M}}\mathcal{Q}^{H}\star_{\mathbf{M}}\mathcal{X}_{D^{o}}\|^{2}_{F}}$ (13)
$s.t.$ $\quad\mathcal{U}^{H}\star_{\mathbf{M}}\mathcal{U}=\mathcal{I}.$
Specifically, $\mathcal{Q}$ contains the weights that measure the importance of each original feature during the projection into the latent feature space. To encode only the necessary original features, $\mathcal{Q}$ is constrained to have sparse loadings. Meanwhile, for highly relevant features, we hope the model can treat them equally, instead of preferring certain features, especially when $dm\gg n$. Therefore, lasso regression and ridge regression terms are introduced as constraints for $\mathcal{Q}$, resulting in a sparse Tensor PCA model:
$\underset{\mathcal{U},\mathcal{Q}}{\min}$ ${\|\mathcal{X}_{D^{o}}-\mathcal{U}\star_{\mathbf{M}}\mathcal{Q}^{H}\star_{\mathbf{M}}\mathcal{X}_{D^{o}}\|^{2}_{F}}+\alpha\sum\limits^{pn}_{j=1}{\|\left(\mathbf{Q}_{(1)}\right)_{:j}\|^{2}_{2}}+\beta\sum\limits^{pn}_{j=1}{\|\left(\mathbf{Q}_{(1)}\right)_{:j}\|_{1}}$ (14)
$s.t.$ $\quad\mathcal{U}^{H}\star_{\mathbf{M}}\mathcal{U}=\mathcal{I}.$
By obtaining the optimal $\mathcal{Q}$, the sparse tensor principal components can be extracted from the data tensor.
Discussion: To this end, the proposed model is directly applicable to tasks traditionally addressed by SPCA (e.g., image compression, visualization, and denoising) while preserving the tensor structure. However, the encoder tensor alone is suboptimal for UFS tasks. In the following section, we reformulate the model to meet UFS-specific requirements.
3.2 Convex version of STPCA-MP for UFS
As previously discussed, the self-contained regression term has an encoder-decoder structure. Therefore, the importance of each original feature is not only expressed by the weights contained in $\mathcal{Q}$ (encoder) but also described by $\mathcal{U}$ (decoder). Comprehensively considering their relationship is essential for adapting the model to UFS.
Dual Problem: Since the transformation between tensor $\mathcal{A}$ and its counterpart $\hat{\mathcal{A}}$ is uniquely defined by $\hat{\mathcal{A}}=\mathcal{A}\times_{3}\mathbf{M}$, Problem (14) has its dual problem in the transform domain, which means they can be solved simultaneously. To facilitate optimization, we can rewrite Problem (14) in the transform domain where matrix algebra can be applied:
$\underset{\hat{\mathcal{U}},\hat{\mathcal{Q}}}{\min}$ $\sum\limits^{p}_{i=1}{\|\left(\hat{\mathcal{X}}_{D^{o}}\right)_{::i}-\hat{\mathcal{U}}_{::i}\hat{\mathcal{Q}}^{H}_{::i}\left(\hat{\mathcal{X}}_{D^{o}}\right)_{::i}\|^{2}_{F}}+\lambda\sum\limits^{p}_{j=1}{\|\hat{\mathcal{Q}}_{::i}\|_{2,1}}$ (15)
$s.t.\quad\hat{\mathcal{U}}_{::i}^{H}\left(\hat{\mathcal{U}}_{::i}\right)=\mathbf{I},i\in[p].$
The $\ell_{2,1}$ norm is employed to simultaneously achieve the goals of lasso regression and ridge regression.
Variable Merging: Notably, there exists a tensor $\mathcal{A}\in\mathbb{C}^{d\times d}$ whose $i$-th frontal slice in the transform domain $\hat{\mathcal{A}}_{::i}=\hat{\mathcal{U}}_{::i}\hat{\mathcal{Q}}^{H}_{::i}$. Each row vector in $\hat{\mathcal{A}}_{::i}$ contains reconstruction weights that combine elements from both $\hat{\mathcal{U}}_{::i}$ and $\hat{\mathcal{Q}}^{H}_{::i}$. Therefore, $\mathcal{A}$ is a more suitable feature selection tensor than $\mathcal{Q}$. Guided by $D^{o}$, each frontal slice in $\hat{\mathcal{A}}$ reconstructs the fibers of a certain mode in the data tensor, utilizing the structural information hidden in the mode organization. Remarkably, it can be proven that the optimal $\hat{\mathcal{A}}$ is a Hermitian positive semidefinite (HPSD) matrix, detailed by Theorem 1.
Theorem 1.
Given the optimal solution $\hat{\mathcal{U}}^{opt}_{::i}$ and $\hat{\mathcal{Q}}^{opt}_{::i}$ to Problem (15), then $\hat{\mathcal{A}}^{opt}_{::i}=\hat{\mathcal{U}}^{opt}_{::i}(\hat{\mathcal{Q}}^{opt}_{::i})^{H}\in H^{q}_{+}$.
[…]
pfProof of Theorem 1 {pf} See Appendix.
Norm Simplification: Since $\hat{\mathcal{A}}_{::i}=\hat{\mathcal{U}}_{::i}\hat{\mathcal{Q}}^{H}_{::i}$ and $d\gg k$, $\hat{\mathcal{A}}_{::i}$ is also a low-rank matrix. It is widely known that nuclear norm $\|\cdot\|_{*}$ is often used as the convex approximation to the low-rank matrix constraint. However, its high computational complexity poses a challenge. Fortunately, in our case, $\|\hat{\mathcal{A}}_{::i}\|_{*}={\rm Tr}\left(\hat{\mathcal{A}}_{::i}\right)$ due to the HPSD constraint, significantly reducing computational complexity.
Meanwhile, provided that $\hat{\mathcal{U}}_{::i}$ is orthogonal, it is easy to prove that
${\|\hat{\mathcal{Q}}_{::i}\|}_{2,1}={\|\hat{\mathcal{U}}_{::i}\hat{\mathcal{Q}}^{H}_{::i}\|}_{2,1}={\|\hat{\mathcal{A}}_{::i}\|}_{2,1}={\|\hat{\mathcal{A}}^{H}_{::i}\|}_{2,1}.$ (16)
Therefore we can use ${\|\hat{\mathcal{A}}_{::i}\|}_{2,1}$ to simplify optimization. Moreover, minimizing ${\|\hat{\mathcal{A}}_{::i}\|}_{2,1}$ and ${\rm Tr}\left(\hat{\mathcal{A}}_{::i}\right)$ is equivalent to minimizing ${\|\mathcal{A}_{::i}\|}_{2,1}$ and ${\rm Tr}\left(\mathcal{A}_{::i}\right)$ when $\mathbf{M}$ is orthogonal. This implies that the elements in a column vector of $\mathcal{A}_{::i}$ (representing different weights for the same feature) are minimized toward zero for redundant features, thereby fulfilling the feature selection task.
Finally, a convex version of STPCA-MP for UFS is formulated as follows:
$\underset{\mathcal{A}}{\min}$ ${\|\mathcal{X}_{D^{o}}-\mathcal{A}\star_{\mathbf{M}}\mathcal{X}_{D^{o}}\|^{2}_{F}}+\lambda\sum\limits^{p}_{i=1}{\|\mathcal{A}}_{::i}\|_{2,1}+\eta\sum\limits^{p}_{i=1}{\rm Tr}\left(\mathcal{A}_{::i}\right)$ (17)
$s.t.$ $\hat{\mathcal{A}}_{::i}\in H^{q}_{+},i\in[p].$
Discussion: The convex version of STPCA-MP exhibits a compact model expression. By merging the sparse encoder and orthogonal decoder, the reconstruction tensor simultaneously assigns weights to original features and their latent representations, comprehensively measuring feature importance.
3.3 Optimization Algorithm
The key to solving Problem (17) lies in leveraging the tensor structure to handle the HPSD constraint with reduced computational complexity.
Multiple Dual Subproblems: The proposed model in the original domain (Problem (17)) has a dual problem in the transform domain, since their solutions are uniquely connected via $\mathbf{M}$. As demonstrated below, the model is significantly easier to solve in the transform domain. First, we rewrite Problem (17) in the transform domain:
$\underset{\hat{\mathcal{A}}}{\min}$ $\sum\limits^{p}_{i=1}\left({\|\left(\hat{\mathcal{X}}_{D^{o}}\right)_{::i}-\hat{\mathcal{A}}_{::i}\left(\hat{\mathcal{X}}_{D^{o}}\right)_{::i}\|^{2}_{F}}+\lambda{\|\hat{\mathcal{A}}_{::i}\|}_{2,1}+\eta{\rm Tr}\left(\hat{\mathcal{A}}_{::i}\right)\right)$ (18)
$s.t.$ $\hat{\mathcal{A}}_{::i}\in H^{q}_{+},i\in[p],$
which can be decomposed into $p$ independent subproblems. We can divide and conquer each subproblem and gather the results to construct $\mathcal{A}$.
Derivative: For the $i$-th subproblem, we calculate the derivative concerning $\hat{\mathcal{A}}_{::i}$ and set the value to 0, yielding:
$\hat{\mathcal{A}}_{::i}=\left(\mathbf{S}_{i}-\frac{\eta_{i}}{2}\mathbf{I}_{q}\right)\left(\mathbf{S}_{i}+\lambda_{i}{\mathbf{W}_{i}}+\epsilon_{2}\mathbf{I}_{q}\right)^{-1},$ (19)
where $\mathbf{S_{i}}=\left(\hat{\mathcal{X}}_{D^{o}}\right)_{::i}\left(\hat{\mathcal{X}}_{D^{o}}\right)^{T}_{::i}$ and $\mathbf{W}_{i}\in{\mathbb{R}^{q\times q}}$ is a diagonal matrix whose $j$-th diagonal element is calculated by:
$\mathbf{w}_{jj}=\left(1/\left(2\sqrt{\hat{\mathbf{a}}_{j}^{H}\hat{\mathbf{a}}_{j}+\epsilon_{1}}\right)\right).$ (20)
HPSD Projection: The next step is to project $\hat{\mathcal{A}}_{::i}$ onto the HPSD cone. This is achieved using the HPSD projection operator $P_{H^{d}_{+}}(\cdot)$ [55] which, for an arbitrary square matrix $\mathbf{A}\in\mathbb{C}^{b\times b}$, is defined as:
$P_{H^{b}_{+}}(\mathbf{A})$ $=\Pi_{H^{b}_{+}}\left(\Pi_{H^{b}}(\mathbf{A})\right)$ (21)
$=\Pi_{H^{b}_{+}}\left(\frac{1}{2}(\mathbf{A}+\mathbf{A}^{H})\right)$
$=\sum\limits_{i=1}^{b}max\left\{\sigma_{i},0\right\}\mathbf{u}_{i}^{H}\mathbf{u}_{i},$
where $\sum\nolimits_{i=1}^{b}\sigma_{i}\mathbf{u}_{i}^{H}\mathbf{u}_{i}$ is the eigenvalue decomposition of $\frac{1}{2}(\mathbf{A}+\mathbf{A}^{H})$.
Based on the above techniques, we solve the $i$-th subproblem of (18) with the following two-stage algorithm:
Step 1: Fix $\mathbf{W}_{i}$ and Update $\hat{\mathcal{A}}_{::i}$.
When $\mathbf{W}_{i}$ is fixed, the optimal $\hat{\mathcal{A}}_{::i}$ is obtained at:
$\hat{\mathcal{A}}_{::i}=P_{H^{q}_{+}}\left(\left(\mathbf{S}-\frac{\eta_{i}}{2}\mathbf{I}_{q}\right)\left(\mathbf{S}+\lambda_{i}{\mathbf{W}_{i}}+\epsilon_{2}\mathbf{I}_{q}\right)^{-1}\right).$ (22)
Step 2: Fix $\hat{\mathcal{A}}_{::i}$ and Update $\mathbf{W}_{i}$.
When $\hat{\mathcal{A}}_{::i}$ is fixed, $\mathbf{W}_{i}$ can computed directly by (20).
Discussion: Formulating a series of convex subproblems simplifies the optimization algorithm design. Although the HPSD projection requires eigenvalue decomposition with a computational complexity of $O(x^{3})$, the exploitation of tensor structures ensures that $x\ll\#{\rm Feature}$. The efficiency of the proposed method is demonstrated in the experimental results.
3.4 Feature selection
We assess the importance of an original feature by its contribution to the reconstruction. Given the presence of the transform domain, the influence of the transform matrix $\mathbf{M}$ must be considered.
Suppose we reorient the data tensor to obtain $\mathcal{X}_{D^{o}}\in\mathbb{C}^{p\times n\times q}$. For the $ij$-th feature $f_{ij}$, its transform-domain $\hat{f}_{ij}$ counterpart and reconstructed version $\tilde{f}_{ij}$ are expressed as:
$\hat{f}_{ij}$ $=\sum\limits^{q}_{h=1}m_{jh}f_{ih},\quad\text{and}$ (23)
$\tilde{f}_{ij}$ $=\sum\limits^{p}_{l=1}\hat{a}_{ilj}\hat{f}_{lj}=\sum\limits^{p}_{l=1}\left(\sum\limits^{p}_{k=1}m_{ik}a_{ilk}\right)\left(\sum\limits^{q}_{h=1}m_{jh}f_{lh}\right)=\sum\limits^{p}_{l=1}\sum\limits^{q}_{h=1}\left(\sum\limits^{p}_{k=1}m_{jh}m_{ik}a_{ilj}\right)f_{lh},$
which indicates that the contribution of $f_{lh}$ to the reconstruction of $f_{ij}$ can be described by $\sum^{p}_{k=1}m_{jh}m_{ik}a_{ilj}$. Therefore, the contribution matrix of $f_{lh}$ is defined as:
$\left[Contribution\right]_{ij}=\sum\limits^{p}_{l=1}\sum\limits^{q}_{h=1}\left(\sum\limits^{p}_{k=1}m_{jh}m_{ik}a_{ilj}\right).$ (24)
Finally, the score for $f_{lh}$ is computed as $\|Contribution\|^{2}_{F}$, which comprehensively captures its influence on the reconstruction of all features and leverages the global information. We summarize the whole algorithm in Algorithm 1.
Discussion: Equation (24) shows that an appropriate choice of $\mathbf{M}$ integrates feature correlation information from other modes into the mode targeted for slice-by-slice optimization. In our proposed method, each feature is scored by globally modeling its dependencies with other features without requiring sample vectorization, thus preserving the data organization and reducing computational complexity.
Algorithm 1 The optimization algorithm for STPCA-MP
0:  A third-order centralized data tensor $\mathcal{X}\in\mathbb{C}^{d_{1}\times d_{2}\times n}$, the order set $D^{o}$ determining $q$, an invertible matrices $\mathbf{M}\in\mathbb{C}^{q\times q}$, regularization parameters $\lambda,\eta$.
0:  $h$ selected features.
1:  Randomly initialize $\left(\tilde{\mathcal{A}}\right)_{::i}\in{H^{q}_{+}}$, transform $\mathcal{X}$ into $\tilde{\mathcal{X}}$.
2:  for $i=1,2,\cdots,q$ do
3:     revolve $\tilde{\mathcal{X}}$ with $D^{o}$ and obtain $\tilde{\mathcal{X}}_{D^{o}}$.
4:     input $\left(\tilde{\mathcal{X}}_{D^{o}}\right)_{::i}$ and alternatively update $\hat{\mathcal{A}}_{::i}$ and $\mathbf{W}_{i}$ according to (22) until converge.
5:  end for
6:  Compute $\mathcal{A}=\hat{\mathcal{A}}\times_{3}\mathbf{M}^{-1}$ and $\|Contribution\|^{2}_{F}$ for each score according to (24).
7:  Output a feature score map $Score$.
8:  Sort $Score$ in descending order, and select features of the $h$ greatest scores. For slice-wise tensor, the score for each dimension will be obtained by summation.
4 Discussion
In this section, we discuss the convergence, computational complexity, and extension to higher-order tensor data of STPCA-MP.
4.1 Convergence analysis
First, we prove the convergence of the proposed algorithm. Let us introduce a lemma [40]:
Lemma 1.
For any nonzero vectors $\mathbf{a}$, $\mathbf{b}\in{\mathbb{R}^{c\times{1}}}$, the following inequality holds:
${\|\mathbf{a}\|_{2}}-\frac{{\|\mathbf{a}\|_{2}^{2}}}{2{\|\mathbf{b}\|_{2}}}\leq{\|\mathbf{b}\|_{2}}-\frac{{\|\mathbf{b}\|_{2}^{2}}}{2{\|\mathbf{b}\|_{2}}}.$ (25)
Based on Lemma 1, we have the following theorem:
Theorem 2.
the objective function value of each subproblem in Problem (18) is non-increasing in each iteration of Algorithm 1 and can converge to a global minimum.
[…]
pf2Proof of 2 {pf2} For simplicity, we temporally use $\mathbf{A}$ and $\mathbf{X}$ to replace $\hat{\mathcal{A}}_{::i}$ and $\left(\hat{\mathcal{X}}_{D^{o}}\right)_{::i}$. Denote the updated $\mathbf{A}$ as $\mathbf{\tilde{A}}$ and fix $\mathbf{W}$. By rewriting the $i$-th subproblem in (18) with the trace function, we have the following inequality
${\rm Tr}\left(\mathbf{\tilde{A}}\mathbf{S}\mathbf{\tilde{A}}^{H}\right)$ $-2Tr\left(\mathbf{S}\mathbf{\tilde{A}}\right)+\lambda{\rm Tr}(\mathbf{\tilde{A}}\mathbf{W}\mathbf{\tilde{A}}^{H})+\eta{\rm Tr}(\mathbf{\tilde{A}})$ (26)
$\leq{\rm Tr}(\mathbf{A}\mathbf{S}\mathbf{A}^{H})-2Tr(\mathbf{S}\mathbf{A})+\lambda{\rm Tr}(\mathbf{A}\mathbf{W}\mathbf{A}^{H})+\eta{\rm Tr}(\mathbf{A}).$
Denote $J(\mathbf{A})={\rm Tr}(\mathbf{A}\mathbf{S}\mathbf{A}^{H})-2Tr(\mathbf{S}\mathbf{A})+\lambda{\rm Tr}(\mathbf{A}\mathbf{W}\mathbf{A}^{H})+\eta{\rm Tr}(\mathbf{A})$ and $J(\mathbf{\tilde{A}})$ as the updated $J(\mathbf{A})$. Inequality (26) can be rewritten as
$J(\mathbf{\tilde{A}})+\lambda{\rm Tr}(\mathbf{\tilde{A}}\mathbf{W}\mathbf{\tilde{A}}^{H})\leq J(\mathbf{A})+\lambda{\rm Tr}(\mathbf{A}\mathbf{W}\mathbf{A}^{H}).$ (27)
Adding the same item $\sum\limits_{j=1}^{d}{\lambda\epsilon}/{2\sqrt{\mathbf{a}_{j}^{H}\mathbf{a}_{j}+\epsilon}}$ to both sides of (27), we obtain
$J(\mathbf{\tilde{A}})$ $+\lambda{\rm Tr}(\mathbf{\tilde{A}}\mathbf{W}\mathbf{\tilde{A}}^{H})+\sum\limits_{j=1}^{d}\frac{{\lambda\epsilon}}{{2\sqrt{\mathbf{a}_{j}^{H}\mathbf{a}_{j}+\epsilon}}}\leq J(\mathbf{A})+\lambda{\rm Tr}(\mathbf{A}\mathbf{W}\mathbf{A}^{H})+\sum\limits_{j=1}^{d}\frac{{\lambda\epsilon}}{{2\sqrt{\mathbf{a}_{j}^{H}\mathbf{a}_{j}+\epsilon}}}.$ (28)
In the meantime, we have
${\rm Tr}(\mathbf{A}\mathbf{W}\mathbf{A}^{H})={\rm Tr}(\mathbf{A}^{H}\mathbf{W}\mathbf{A})=\sum\limits_{j=1}^{d}\frac{{\lambda{\mathbf{a}_{j}^{H}\mathbf{a}_{j}}}}{{2\sqrt{\mathbf{a}_{j}^{H}\mathbf{a}_{j}+\epsilon}}}.$ (29)
The same equality holds for ${\rm Tr}(\mathbf{\tilde{A}}\mathbf{W}\mathbf{\tilde{A}}^{H})$. We substituted (29) into (28) and get
$J(\mathbf{\tilde{A}})+\sum\limits_{j=1}^{d}\frac{{\lambda({\mathbf{\tilde{a}}_{j}^{H}\mathbf{\tilde{a}}_{j}}+\epsilon)}}{{2\sqrt{\mathbf{a}_{j}^{H}\mathbf{a}_{j}+\epsilon}}}\leq J(\mathbf{A})+\sum\limits_{j=1}^{d}\frac{{\lambda({\mathbf{a}_{j}^{H}\mathbf{a}_{j}}+\epsilon)}}{{2\sqrt{\mathbf{a}_{j}^{H}\mathbf{a}_{j}+\epsilon}}}.$ (30)
According to Lemma 1, we have
$\sqrt{{\mathbf{\tilde{a}}_{j}^{H}\mathbf{\tilde{a}}_{j}}+\epsilon}-\frac{{{\mathbf{\tilde{a}}_{j}^{H}\mathbf{\tilde{a}}_{j}}+\epsilon}}{{2\sqrt{\mathbf{a}_{j}^{H}\mathbf{a}_{j}+\epsilon}}}\leq\sqrt{{\mathbf{a}_{j}^{H}\mathbf{a}_{j}}+\epsilon}-\frac{{{\mathbf{a}_{j}^{H}\mathbf{a}_{j}}+\epsilon}}{{2\sqrt{\mathbf{a}_{j}^{H}\mathbf{a}_{j}+\epsilon}}}.$ (31)
Further, we can accumulate (30) from $j=1$ to $j=d$. Considering $\lambda>0$, we can get
$\lambda\sum\limits_{j=1}^{d}\sqrt{{\mathbf{\tilde{a}}_{j}^{H}\mathbf{\tilde{a}}_{j}}+\epsilon}-\lambda\sum\limits_{j=1}^{d}\frac{{{\mathbf{\tilde{a}}_{j}^{H}\mathbf{\tilde{a}}_{j}}+\epsilon}}{{2\sqrt{\mathbf{a}_{j}^{H}\mathbf{a}_{j}+\epsilon}}}\leq\lambda\sum\limits_{j=1}^{d}\sqrt{{\mathbf{a}_{j}^{H}\mathbf{a}_{j}}+\epsilon}-\lambda\sum\limits_{j=1}^{d}\frac{{{\mathbf{a}_{j}^{H}\mathbf{a}_{j}}+\epsilon}}{{2\sqrt{\mathbf{a}_{j}^{H}\mathbf{a}_{j}+\epsilon}}}.$ (32)
By summing (30) and (32), we have
$J(\mathbf{\tilde{A}})+\lambda\sum\limits_{j=1}^{d}\sqrt{{\mathbf{\tilde{a}}_{j}^{H}\mathbf{\tilde{a}}_{j}}+\epsilon}\leq J(\mathbf{A})+\lambda\sum\limits_{j=1}^{d}\sqrt{{\mathbf{a}_{j}^{H}\mathbf{a}_{j}}+\epsilon}.$ (33)
Finally, we can get the following inequality
$J(\mathbf{\tilde{A}})+\lambda{\|{\mathbf{\tilde{A}}}\|_{2,1}}\leq J(\mathbf{A})+\lambda{\|{\mathbf{A}}\|_{2,1}}.$ (34)
Each subproblem in (18) is convex and bounded below by 00, ensuring convergence to the global minimum. Since the subproblems are solved independently, Problem (18) is guaranteed to converge to a global solution. When the solution in the transform domain is obtained, the dual problem in the original domain is solved as well.
4.2 Computational complexity
Suppose $\mathcal{X}\in\mathbb{C}^{d_{1}\times d_{2}\times n}$ is rotated into $\mathcal{X}_{D^{o}}\in\mathbb{C}^{q\times n\times p}$ ($n$ is the number of samples, $(p,q)\in\{(d_{1},d_{2}),(d_{2},d_{1})\}$). For each subproblem, it takes $O\left(q^{2}\right)$ to convert data into the transform domain and initialize the data covariance matrix, and $O\left(q^{3}\right)$ to update the reconstruction matrix in each iteration (including matrix inversion and HPSD projection). Therefore we have $p$ subproblems each solved at the computational cost of $O\left(q^{3}t_{i}+q^{2}n\right)$ ($t_{i}$ is the number of iterations for the $i$-th subproblem). The final feature score computation takes $O(p^{2}q)$. Eventually, the total computational complexity is $O\left(\sum\limits^{p}_{i=1}\left(q^{2}n+q^{3}t_{i}\right)+p^{2}q\right)$.
The number of features ($q\times p$) is usually much greater than the number of dimensions in a single mode ($q$ or $p$). Due to the exploitation of tensor structures with separate operations on each slice, STPCA-MP can handle data tensors at a remarkably smaller computation cost than non-tensor-based methods. For example, for image data of size $32\times 32$, non-tensor-based methods requiring matrix inversion or eigenvalue decomposition (such as SOGFS [40] and SPCAFS [38]; see Table 1) have a computational complexity of $O(1024^{3})$, whereas STPCA-MP requires only $O(32\times(32^{3}))$, reducing the complexity by $1024$ times. In addition, the computational complexity is linear to the number of samples, which means the proposed method will still be efficient when facing data tensors with numerous samples.
Table 1 shows the computational complexity of competitive methods in the following experiments. We can see that tensor-based methods reduce the computational burden from $\prod_{i}d_{i}^{x}$ to $d_{i}^{x}$ in each iteration by utilizing the tensor structure. Among tensor-based methods, CPUFS and CPUFSnn are the most recently proposed, which have a computational complexity linear to $d_{i}$ at the price of squared to $n$. They are non-convex models and usually take hundreds of iterations to converge, making it a challenge when facing large-scale datasets.
Table 1: Computational complexity of competitive methods on third-order tensors. $d=d_{1}d_{2}$ is the number of features. $t_{i}$ stands for the number of iterations. $c$ stands for the number of classes.
Method Input Computational complexity
SOGFS [40] matrix $O\left(dn^{2}+t(d^{3}t_{1}+n^{2}c)\right)$
RNE [50] matrix $O\left(dn^{2}+t(t_{1}(n+s)d^{2}+t_{2}n))\right)$
SPCAFS [39] matrix $O\left(d^{2}n+d^{3}t\right)$
NLGMS [56] matrix $O\left(t\left(d^{3}+n^{3}+nd+n^{2}d+d\log{d}\right)\right)$
CPUFS [18] tensor $O\left(d_{1}nc^{2}+d_{2}nc^{2}+d_{1}d_{2}nc+n^{2}c)t\right)$
CPUFSnn [18] tensor $O\left((d_{1}nc^{2}+d_{2}nc^{2}+d_{1}d_{2}nc+(n_{1}c)^{2}+(n_{2}c)^{2}+n^{2}c)t\right)$
MSPCA [33] tensor $O\left(t_{2}\sum_{i=1}^{2}(t_{1}+n)d_{i}^{3}\right)$
STPCA-MP tensor $O\left(\sum^{p}_{i=1}\left(q^{2}n+q^{3}t_{i}\right)+p^{2}q\right)(q\in\{d_{1},d_{2}\})$
4.3 Extension to higher-order tensor data
By conducting the $\star_{\mathbf{M}}$-product in a recursive fashion, the proposed method can be extended to higher-order tensors. Given two $n$-th order tensors $\mathcal{A}\in\mathbb{C}^{d_{1}\times d_{2}\cdots\times d_{n-1}\times d_{n}}$ and $\mathcal{B}\in\mathbb{C}^{d_{2}\times h\cdots\times d_{n}}$, along with a series of invertible matrices $\mathbf{M}_{i}\in\mathbb{C}^{d_{i+2}\times d_{i+2}}$, the $\star_{\mathbf{M}}$-product for higher-order tensors is defined as shown in Algorithm 2. Using this technique, STPCA-MP can be formulated for higher-order settings. However, sequential $\star_{M}$-products with different orientations introduce significant optimization. This remains an open problem under active investigation. Therefore, this paper focuses on the third-order case.
Algorithm 2 $\star_{\mathbf{M}}$-product for higher-order tensor data
0:  $\mathcal{A}\in\mathbb{C}^{d_{1}\times d_{2}\cdots\times d_{n-1}\times d_{n}}$, $\mathcal{B}\in\mathbb{C}^{d_{2}\times h\cdots\times d_{n}}$, $\mathbf{M}_{i}\in\mathbb{C}^{d_{i+2}\times d_{i+2}}$
0:  $\mathcal{C}=\mathcal{A}\star_{M}\mathcal{B}$
1:  Project the input data into the transform domain: $\hat{\mathcal{A}}=\mathcal{A}\times_{3}\mathbf{M}_{1}\times_{4}\mathbf{M}_{2}\cdots\times_{n}\mathbf{M}_{n-2}$, ^=×$\hat{\mathcal{B}}=\mathcal{B}\times_{3}\mathbf{M}_{1}\times_{4}\mathbf{M}_{2}\cdots\times_{n}\mathbf{M}_{n-2}$
2:  for $i=1,2,\cdots,\prod\limits^{n}_{j=3}d_{j}$ do
3:     $\hat{\mathcal{C}}=\hat{\mathcal{A}}_{::i_{3}\cdots i_{n}}\hat{\mathcal{B}}_{::i_{3}\cdots i_{n}}$
4:  end for
5:  Project the result back to the original domain $\mathcal{C}=\hat{\mathcal{C}}\times_{n}\mathbf{M}^{-1}_{n-2}\times_{n-1}\mathbf{M}^{-1}_{n-1}\cdots\times_{3}\mathbf{M}^{-1}_{1}$.
6:  Output $\mathcal{C}$
5 Experiments
In this section, we conduct experiments on real-world data to demonstrate our proposed method’s effectiveness. The codes related to our proposed methods and the experiments are available at https://github.com/zjj20212035/STPCA.git.
5.1 Experimental Settings
5.1.1 Experimental environment
All the experiments are conducted on a personal computer with a 2.20 GHz CPU and 64GB main memory under the environment of the Windows 11 operation system.
5.1.2 Comparative methods
We compare our proposed methods with some state-of-the-art UFS methods as follows:
  1. [•]
  2. 1.
    CPUFS and CPUFSnn [18] (2023): CPUFS is a tensor-based UFS method combining CP decomposition and spectral clustering. It utilizes a sparse classifier to perform UFS. CPUFSnn is a variation with nonnegative constraints on the classifier matrix.
  3. 2.
    MSPCA [33] (2014): MSPCA is a sparse tensor PCA method that utilizes Tucker decomposition to make use of tensor structure. It was originally proposed as a feature extraction method. We use the Kronecker product of its factor matrices to perform UFS, as an example of Tucker-based UFS methods.
  4. 3.
    NLGMS [56] (2024): NLGMS is a non-tensor-based method that introduces a self-representation technique into the graph-based framework to preserve global and local data structure. The self-representation term reconstructs each sample with the linear combination of other samples.
  5. 4.
    SPCAFS [38] (2021): SPCAFS is a non-tensor-based method that introduces $\ell_{2,p}$-norm into the one-dimensional PCA model. It forms a simple yet effective model.
  6. 5.
    RNE [50] (2020): RNE is a non-tensor-based method that makes use of the locally linear embedding algorithm to obtain an optimal projection matrix. It keeps the samples close to their neighbors after being projected onto a local manifold.
  7. 6.
    SOGFS [40] (2019): SOGFS is a non-tensor-based method that optimizes an adaptive graph to preserve the local geometrical structure of data. It updates the similarity matrix during each iteration to obtain a more flexible graph.
We also perform clustering with all features as a baseline method. All the comparative methods are implemented using the codes provided by their authors.
5.1.3 Parameter settings
For all methods, we tune the regularization parameters via grid search over $\{10^{-2},10^{-1},1,10,10^{2}\}$. For graph-based methods, we set the number of k-nearest neighbors to 5. For methods utilizing $\ell_{2,p}$-norm, we set $p=1$. For STPCA-MP, we set $\mathbf{M}=\mathbf{I}$ in most experiments except the one that discusses the choice of $\mathbf{M}$. The direction set $D^{o}$ is also tuned by a grid-search strategy in the range $\{\{1,3,2\},\{2,3,1\}\}$. The other parameters remain the default in the open code provided by the authors. The best results of all methods are presented.
5.1.4 Evaluation methodology and metrics
In clustering experiments, after a UFS method has finished selecting features, we evaluate its performance by first conducting K-means clustering and then mapping the obtained pseudo labels to the real labels by adopting the Kuhn-Munkres algorithm [41]. Due to the K-means’ dependence on initialization, we repeat the clustering 30 times. The clustering metrics we use are as follows:
  1. [•]
  2. 1.
    Accuracy (ACC): ACC refers to the clustering accuracy and is computed according to (35) and (36), where $n$ is the total number of samples, $map(i)$ denotes the clustering label of the $i$-th sample after mapping to the real label, and $label(i)$ represents the ground truth label. A larger ACC indicates better performance.
    $\text{ACC}=\frac{1}{n}\sum\limits_{i=1}^{n}\delta\left(map(i),label(i)\right).$ (35)
    $\delta(p,q)=\left\{\begin{aligned} &1\quad\text{if}\;p=q,\\ &0\quad\text{otherwise}.\end{aligned}\right.$ (36)
  3. 2.
    Normalized Mutual Information (NMI) [42]: NMI is utilized to describe the mutual dependence between clustering results (after being mapped into the real labels) and ground truth labels. Given the clustering results $\mathbf{m}$ and the ground truth labels $\mathbf{l}$ (both including all samples), $I(\mathbf{m},\mathbf{l})$ denotes the mutual information between $\mathbf{m}$ and $\mathbf{l}$. $H(\mathbf{m})$ and $H(\mathbf{l})$ are the entropy of $\mathbf{m}$ and $\mathbf{l}$ respectively. The same as ACC, a larger NMI means better clustering performance.
    $\text{NMI}(\mathbf{m},\mathbf{l})=\frac{I(\mathbf{m},\mathbf{l})}{\sqrt{H(\mathbf{m})H(\mathbf{l})}}.$ (37)
  4. 3.
    Training time: For an algorithm, training time starts from the initialization and ends when the algorithm outputs the selected features.
For orientation-specific data such as multi-channel time series, the temporal dependency of the elements within the same channel makes the clustering metrics unreliable due to the feature independence hypothesis in K-means. Therefore, we define a metric as follows:
  1. [•]
  2. 1.
    Proportion of correctly selected features (POC): We compute POC as:
    $\text{POC}=c/h,$ (38)
    where $c$ is the number of correctly selected features, and $h$ is the total number of selected features at each execution. POC reflects a UFS method’s ability to select discriminative features. In the experiments, we select features with maximal between-class variance (BCV) as the correct ones. We compute POC on slice-wise data instead of ACC and NMI.
5.2 Datasets
Before detailing specific datasets, we first categorize tensor datasets based on their structural organization. As shown in Figure 2, third-order tensor datasets can be divided into two scenarios: tube-wise and slice-wise, according to their organization. Slice-wise data is orientation-specific and exhibits feature relations along the 1-mode, while tube-wise data exhibits feature relations along both the 1-mode and 2-mode.. This distinction provides a basis for evaluating the ability of UFS methods to utilize tensor structure information. In our evaluation, we compare clustering ACC and NMI on tube-wise data and POC on slice-wise data. Experimental results demonstrate the effectiveness of the proposed method in both scenarios.
Table 2: Two scenarios for third-order data tensors. Assume each frontal slice is a sample. Different colors refer to different variables. This paper regards variables as features. Tube-wise scenario: elements within one fiber belong to one variable. Slice-wise scenario: elements within one horizontal slice belong to one variable.
Scenario Visualization Example
Tube-wise [Uncaptioned image] 1. snapshots of array signal [8, 10, 11]. 2. images and video[7, 12, 13].
Slice-wise (orientation-specific) [Uncaptioned image] 1. time series with different channels [14, 15, 16, 9]. 2. multi-way statistical tables [17].
To preliminarily evaluate the UFS ability of our proposed method in the two tensor scenarios, we design two synthetic tensor datasets:
  1. [•]
  2. 1.
    Orbit: As an example of slice-wise data, Orbit comprises 3-dimensional orbit samples with two distinct radii. Each sample includes three channels representing orbit coordinates, with the remaining channels containing random Gaussian noise.
  3. 2.
    Array: As an example of tube-wise data, Array is constructed using a $10\times 10$ synthetic radar unit array that receives a series of echoes from a stationary target. Each echo is represented as a matrix of unit sampling values, treated as a sample. Random measurement errors are introduced to half of the echo matrices to create two distinct classes.
We evaluate our method on nine real-world datasets, including: 1) six tube-wise datasets: COIL20 [54], USPS [52], BreastMNIST [44], Imm40 [51], PIE [13] and JAFFE [43]; 2) four slice-wise datasets: UCIHAR [45] and UCIDSA [46], UMCS [^1] https://www.kaggle.com/c/seizure-detection/data and ASCH [53]. The detailed information of each dataset is shown in Table 3. PIE, Imm40, and JAFFE contain frontal facial images, labeled according to individuals. BreastMNIST is a medical dataset containing breast ultrasound images, from which we construct the data tensor using the test set. USPS is a dataset of handwritten digits. UCIHAR and UCIDSA are human activity datasets collected from multiple subjects using motion sensors. UCIDSA contains multi-channel signals, while UCIHAR comprises extracted statistical features (e.g. mean, std, max, entropy). UMCS and ASCH are electroencephalogram (EEG) datasets collected for disease research, capturing brain activity signals. Due to the large volume of chosen slice-wise data, we retrieve a subset of samples to construct datasets. For UCIDSA, we extract the first $125$ sample points from all $45$ channels covering $45$ activities. For UCIHAR, we select $9$ individuals with all $561$ features sampled during $6$ activities. For UMCS, we collect the first $125$ sample points in all $65$ channels of the first human patient, classified into ictal and interictal segments. For ASCH, we collect the first $640$ sample points in all $16$ channels from healthy groups and groups with symptoms. All datasets are normalized and bounded within $[-1,1]$ for consistency.
Table 3: Statistics of datasets.
Type Dataset #Feature #Samples #Classes
Tube-wise Array $10\times 10$ 100 2
USPS [52] $16\times 16$ 9298 10
BreastMNIST [44] $28\times 28$ 156 2
Imm40 [51] $32\times 32$ 240 40
COIL20 [54] $32\times 32$ 1440 20
PIE [13] $32\times 32$ 1166 53
JAFFE [43] $64\times 64$ 213 7
Slice-wise Orbit $9\times 41$ 100 2
UCIDSA [46] $45\times 125$ 152 19
UCIHAR [45] $561\times 9$ 273 6
UMCS 1 $65\times 125$ 174 2
ASCH [53] $16\times 640$ 84 2
5.3 Synthetic experiment
In this experiment, we conduct four typical UFS methods on synthetic datasets with the grid search strategy. The dataset formulation and the best experimental results of comparative methods are illustrated in Figure 3. On Orbit, we retrieve three top-ranked features in each method’s result and plot their samples in a subset of data. If a method selects the correct features (i.e. the three orbit coordinates), then the scatter plot will contain rings of two different radii. Otherwise, it will be difficult to distinguish the two classes. On Array, we retrieve the feature score map corresponding to the best POC for each method, analyzing the match degree between the top-ranked feature positions and the units with measurement error. We conclude the observations as follows.
  1. [•]
  2. 1.
    Tensor-based methods are generally more capable of analyzing the data structure than non-tensor-based methods. Both CPUFS and STPCA-MP select the orbit coordinates on Orbit. Although CPUFS fails to precisely locate the units with error on Array, its score map clearly shows a grid-like characteristic that connects the rows and columns to analyze feature relations, capturing a roughly correct attention region. A similar pattern also appears in STPCA-MP, where the column-wise relations are learned. In contrast, RNE and NLGMS isolate the features, losing the tensor structure information. Especially in RNE, the scores on Array seem to be randomly distributed.
  3. 2.
    Our proposed method performs the best in selecting discriminative features on the two synthetic datasets. STPCA-MP selects the correct features on both datasets. In particular, the score map on Array is relatively sparse in rows and columns. By taking advantage of data orientation, STPCA-MP can more easily mine the latent feature relations and locate discriminative features.
Figure 3: Synthetic experiments on Orbit and Array. The left part gives the composition of the two synthetic data tensors. The right part shows the feature selection results of comparative methods in a visualized manner.
5.4 Clustering experiment on tube-wise data
In this experiment, we perform clustering on five tube-wise real-world datasets. The regularization parameters of all methods go through a grid search in $\{10^{-2},10^{-1},1,10^{1},10^{2}\}$. We set the selected features to $\{50,100,150,200,250,300\}$ on COIL20, JAFFE, BreastMNIST , Imm40, and $\{10,30,50,70,90,110\}$ on USPS. The best results and the corresponding training time of all methods are given in Table 4. The detailed ACC and NMI curves are illustrated in Figure 4. By analyzing the results, we come out with the following conclusions:
  1. [•]
  2. 1.
    Features selection improves data compactness. On each dataset, one or more methods are capable of achieving performance that approaches or even surpasses the baseline while utilizing only a subset of features. In particular, the proposed STPCA-MP achieves significantly higher clustering metrics than all features on Jaffe, BreastMNIST, and Imm40.
  3. 2.
    STPCA-MP delivers reliable and outstanding performance in clustering tasks. Firstly, compared to graph-based methods like SOGFS, STPCA-MP obtains more global data information by exploiting the reconstruction term, which guarantees the basic clustering performance. We can see the curves of STPCA-MP, on the whole, have a smaller oscillation in amplitude and share better consistency in trend between ACC and NMI. Besides, STPCA-MP reflects a more natural decreasing (or increasing) progression in variation pattern. Secondly, the T-SVDM-based sparse framework is beneficial for extracting tensor structure information, enabling STPCA-MP to describe the feature correlation in a more delicate manner, compared to CPD-based tensor methods (CPUFS and CPUFSnn), Tucker-based tensor method (MSPCA), and non-tensor-based methods (SPCAFS, RNE). Therefore, STPCA-MP demonstrates superior performance, consistently ranking at the forefront on all datasets.
  4. 3.
    STPCA-MP achieves the best performance at a smaller training time cost. On all datasets in Table 4, STPCA-MP takes less than $1$ second to complete training and feature selection, due to its slice-by-slice operation. Note that this is accomplished with reasonable prior knowledge of data organization. As a comparison, NLGMS, also showing remarkable average performance, brings much more computational cost as it overlooks the tensor structure by vectorization operation.
Table 4: Metrics of the comparative methods on real-world datasets. The best ACC, best NMI, and mean POC regarding grid-search range are shown. The training time corresponding to the best result is recorded in seconds. The best result on each dataset is boldfaced, while the second best one is underlined and italic.
Method PIE COIL20 JAFFE BreastMNIST Imm40 USPS UCIDSA UCIHAR UMCS ASCH
All Features ACC 26.21 ACC 58.34 ACC 20.16 ACC 58.25 ACC 48.43 ACC 65.29 POC - POC - POC - POC -
NMI 51.19 NMI 75.74 NMI 4.51 NMI 2.41 NMI 74.21 NMI 61.11
RNE[50] ACC 18.67 ACC 57.83 ACC 26.96 ACC 60.26 ACC 58.54 ACC 55.51 POC TIME 40.00 83.57 POC TIME 0 170.22 POC TIME 40.00 193.22 POC TIME 66.67 205.33
NMI 42.64 NMI 73.81 NMI 12.89 NMI 4.34 NMI 81.11 NMI 50.74
TIME 5.56 TIME 18.17 TIME 63.42 TIME 8.16 TIME 15.44 TIME 1.60
SOGFS[40] ACC 28.22 ACC 57.54 ACC 24.88 ACC 61.45 ACC 57.06 ACC 57.63 POC TIME 37.33 2409.95 POC TIME 1.60 5569.15 POC TIME 17.33 5274.80 POC TIME 23.33 8986.97
NMI 53.01 NMI 72.55 NMI 10.01 NMI 4.55 NMI 79.60 NMI 52.21
TIME 34.28 TIME 55.56 TIME 1794.41 TIME 37.12 TIME 83.77 TIME 391.60
SPCAFS[38] ACC 41.55 ACC 58.10 ACC 29.48 ACC 60.90 ACC 62.86 ACC 64.63 POC TIME 70.67 140.61 POC TIME 23.60 274.54 POC TIME 28.00 435.80 POC TIME 70.00 6569.55
NMI 65.04 NMI 74.02 NMI 12.07 NMI 5.14 NMI 84.09 NMI 61.13
TIME 18.37 TIME 3.11 TIME 59.36 TIME 0.61 TIME 15.76 TIME 0.34
NLGMS[56] ACC 38.40 ACC 61.30 ACC 30.94 ACC 61.86 ACC 59.51 ACC 66.08 POC TIME 40.53 222.98 POC TIME 17.08 410.78 POC TIME 21.87 782.96 POC TIME 46.00 3114.56
NMI 62.69 NMI 76.74 NMI 18.34 NMI 4.29 NMI 82.03 NMI 61.45
TIME 25.99 TIME 25.71 TIME 498.12 TIME 2.98 TIME 19.08 TIME 511.68
CPUFS[18] ACC 31.30 ACC 60.05 ACC 27.82 ACC 61.11 ACC 59.18 ACC 64.99 POC TIME 34.35 51.01 POC TIME 19.16 71.78 POC TIME 21.49 11.47 POC TIME 34.93 5.91
NMI 56.56 NMI 75.48 NMI 13.42 NMI 3.48 NMI 81.41 NMI 59.79
TIME 290.69 TIME 69.89 TIME 11.20 TIME 6.87 TIME 14.01 TIME 1559.53
CPUFSnn[18] ACC 32.51 ACC 61.47 ACC 26.15 ACC 61.54 ACC 57.85 ACC 62.79 POC TIME 34.40 48.77 POC TIME 23.41 17.79 POC TIME 23.75 12.66 POC TIME 34.27 6.67
NMI 32.51 NMI 75.85 NMI 11.86 NMI 4.03 NMI 79.18 NMI 58.37
TIME 101.76 TIME 20.72 TIME 11.68 TIME 7.11 TIME 13.35 TIME 485.50
MSPCA[33] ACC 30.66 ACC 59.74 ACC 27.51 ACC 63.53 ACC 62.54 ACC 63.37 POC TIME 86.67 0.37 POC TIME 24.20 3.22 POC TIME 30.80 0.02 POC TIME 69.00 0.15
NMI 53.81 NMI 75.84 NMI 10.47 NMI 3.46 NMI 83.40 NMI 59.86
TIME 0.33 TIME 0.02 TIME 0.01 TIME 0.01 TIME 0.07 TIME 0.05
STPCA-MP (ours) ACC 42.84 ACC 61.64 ACC 34.74 ACC 64.34 ACC 68.92 ACC 66.77 POC TIME 92.27 0.08 POC TIME 35.20 0.17 POC TIME 41.07 0.04 POC TIME 83.33 0.94
NMI 66.80 NMI 76.29 NMI 23.42 NMI 5.20 NMI 87.81 NMI 61.75
TIME 0.03 TIME 0.04 TIME 0.05 TIME 0.02 TIME 0.02 TIME 0.06
Figure 4: The ACC and NMI curves on tube-wise datasets. The upper row shows the ACC curves, while the bottom row displays the NMI curves.
5.5 POC comparison experiment on slice-wise data
For slice-wise datasets, we view each channel (row vector) as a feature. Therefore, we perform this experiment in the following steps: First, we implement a UFS algorithm and obtain the scores for each element. Second, we sum row-wise scores to evaluate features, using POC to measure the agreement between top-ranked features and those with maximal between-class variance (BCV). We compute BCV after the data are zero-centered.
The BCV of selected features on four slice-wise datasets is shown in Figure 5, where the results corresponding to the greatest POC of typical comparative methods are given. The mean POC of all methods is illustrated in Table 4. We have the following observations:
  1. [•]
  2. 1.
    For each dataset, there exists a small set of features with prominent BCV, judging by the peaks on the BCV curves in Figure 5. Greater BCV usually implies better inter-class discriminability.
  3. 2.
    SPCA-based methods are generally more capable of selecting the features of the greatest BCVs. SPCAFS, MSPCA, and STPCA-MP have a stabler performance in terms of POC, owing to the encoder-decoder framework that minimizes the reconstruction error as well as maximizes the variance among samples on the global manifold. Attributed to the sparse tensor decomposition, MSPCA also achieves relatively stable performance on slice-wise data, referring to its mean POC in Table 4.
  4. 3.
    STPCA-MP has the advantage of handling slice-wise data over comparative methods. Due to the exploitation of data orientation, STPCA-MP can better learn the intrinsic relationship between different channels in a slice-wise tensor, selecting a considerable amount of BCV peaks in Figure 5. Although the best result of NLGMS is compatible to our method on some datasets, STPCA-MP demonstrates enhanced stability in selecting discriminative features, according to the mean POC regarding grid-search range in Table 4. On average, STPCA-MP is more effective toward slice-wise data, compared to the state-of-the-arts.
Figure 5: The BCV of selected features on slice-wise datasets.
5.6 Visualization
In this experiment, we perform a visualization to empirically compare the interpretability of selected features. The best results of three tensor-based methods and one non-tensor-based method in terms of NMI on two image datasets: JAFFE and BreastMNIST, are visualized in Figure 6, where we highlight the selected features to see the attention region of the UFS methods. We have the following observations:
  1. [•]
  2. 1.
    STPCA-MP pays more attention to discriminative features. The exploitation of data orientation enables STPCA-MP to mine essential information with different spatial distributions. On JAFFE, STPCA-MP selects more horizontal features that cover the regions of eyes, eyebrows, mouth, and hair, etc. beneficial for distinguishing facial expressions, while CPUFS and MSPCA are misguided by background pixels. NLGMS, though it selects reasonable features as well, has a more scattered region and fails to select correlated features. On BreastMNIST, the features selected STPCA-MP tend to be grouped at an abnormal shadowed region composed of both horizontally and vertically distributed features, while the results of CPUFS, MSPCA and NLGMS are diverse. We can also observe that compared to CPUFS, MSPCA selects more rows (or columns) containing discriminative features and is less affected by the background, which we believe is attributed to the direct utilization of sparse factor matrices (CPUFS doesn’t apply a sparse tensor decomposition).
  3. 2.
    The score map of STPCA-MP exhibits significantly clearer and more distinct patterns than comparative methods. Thanks to the slice-by-slice operation, STPCA-MP describes feature correlations in a more delicate manner. On JAFFE, the score map of STPCA-MP approximates the shape of a human face. On BreastMNIST, the score map of STPCA-MP demonstrates a smooth score gradient transition radiating from the center of the shadowed region. In contrast, the score maps of CPUFS and MSPCA are less interpretable, due to the vague description of mode relations rooted in CPD and Tucker decomposition. As for NLGMS, its score map adds significant weights to the redundant features in the background. This phenomenon is possibly evoked by the sensitivity to noise features of the constructed graph.
Figure 6: Visualization on JAFFE and BreastMNIST. We show the score maps and highlight the positions of the top-ranked features in the samples. We set the number of selected features to $300$ on JAFFE, and $50$ on BreastMNIST.
5.7 Training time comparison
In this experiment, we compare the computational cost of comparative methods. The training time corresponding to the best clustering results is given in Table 4, while the average training time in terms of regularization parameter combinations is illustrated in Figure 7. We have the following observations:
  1. [•]
  2. 1.
    Tensor-based methods are generally more efficient on high-dimensional data. By leveraging tensor-based techniques, the computational complexity is reduced from quadratic or cubic in the full feature space to quadratic or cubic within a single mode dimension. Therefore CPUFS, CPUFSnn, MSPCA, and our proposed method averagely cost less training time on JAFFE, UMCS, and ASCH compared to other methods.
  3. 2.
    STPCA-MP is efficient on both high-dimensional and large-scale data. CPUFS and CPUFSnn usually take hundreds of iterations to converge because of the non-convexity. Moreover, their computational complexity scales quadratically with #$\#(Sample)$, whereas STPCA-MP exhibits linear complexity relative to #$\#(Sample)$. Therefore, STPCA-MP takes considerably less training time than CPUFS and CPUFSnn on all datasets, especially on large-scale datasets. Although MSPCA is on average more efficient than STPCA-MP due to a smaller weight parameter size requiring computation in Tucker decomposition, it gains no advantage over the proposed method in both clustering tasks and the POC metric. Moreover, the training time associated with optimal performance shows only a marginal discrepancy between MSPCA and STPCA-MP (See Table 4), which is in an acceptable range. On the whole, the proposed method is both effective and efficient.
(a) JAFFE (#Feature=$4096$, #Sample$=213$)
(b) UMCS (#Feature=$8125$, #Sample$=174$)
(c) ASCH (#Feature=$10240$, #Sample$=84$)
(d) USPS (#Feature=$784$, #Sample$=9298$)
Figure 7: Average training time during grid search.
5.8 Convergence analysis
In this experiment, we study the convergence of the proposed method. Figure 8 illustrates the converge curves of objective function value on four real-world datasets. Utilizing the HPSD projector, the proposed two-step algorithm converges fast into the solution space within dozens of iterations.
(a) PIE
(b) JAFFE
(c) BreastMNIST
(d) UCIDSA
Figure 8: Converge curves of STPCA-MP on JAFFE and BreastMNIST. All regularization parameters are set to 1.
5.9 Parameter sensitivity analysis
In this section, we analyze the impact of the parameters $\lambda$ and $\eta$ in STPCA-MP on the performance of feature selection. We vary the parameters in the range of $\{10^{-2},10^{-1},1,10^{1},10^{2}\}$, set the selected features to 100, and record the ACC and NMI of STPCA-MP with all possible parameter combinations. The results on JAFFE and USPS is illustrated in Figure 9. We can see that
  1. [•]
  2. 1.
    On JAFFE, the performance of STPCA-MP is relatively stable when $\lambda$ and $\eta$ vary in a certain range, and degenerates when the parameters become too large. This tendency is consistent between ACC and NMI.
  3. 2.
    On USPS, STPCA-MP is generally not sensitive to the choice of parameters, achieving relatively stable performance with respect to both ACC and NMI.
(a) JAFFE (ACC)
(b) JAFFE (NMI)
(c) USPS (ACC)
(d) USPS (NMI)
Figure 9: Parameter sensitivity analysis of STPCA-MP on JAFFE and USPS.
5.10 Discussion on different choices on data orientation and transform domain.
In this experiment, we discuss the influence of data tensor orientation and transform domain. We conduct STPCA-MP with different data tensor rotations and transform matrices $\mathbf{M}$ on several datasets. The detailed ACC curves on JAFFE and USPS are illustrated in Figure 10. We further provide the best ACC and the mean POC in Table 5. The visualized result with $\mathbf{M}=\mathbf{I}$ on PIE and JAFFE is given in Figure 11. We have the following observations:
  1. [•]
  2. 1.
    Different data orientations can significantly affect STPCA-MP’s clustering performance. We can see significant gaps among certain ACC curves in Figure 10. Like looking at an identical object from different views, decomposing a data tensor in different directions naturally leads to discrepant information. The visualized result in Figure 11 implies that STPCA-MP can capture features with different spatial distributions given different data orientations. A proper choice of data orientation is beneficial for mining feature correlations. The spatial distribution and latent correlation structure of discriminative features are heterogeneous among data tensors. On UCIDSA and UMCS, the overall performance of the Dir1s is better than the Dir2s, since the feature relations in multi-channel time-series data exist along 1-mode. In conclusion, performing slice-by-slice operations along a certain direction benefits UFS at multiple levels on different datasets.
  3. 2.
    A suitable transform domain can benefit the performance. Since the transform extracts the information from one of the two feature modes, while a slice-by-slice operation is conducted along the other feature mode, a proper transform domain can integrate more feature correlations as prior knowledge for the following optimization and UFS. In Table 5, Dir1-EigVector sometimes achieves better results than Dir-I, especially on time-series data. However, there is no absolute conclusion that certain settings of transform domain necessarily outperform others. As we can observe in both Figure 10 and Table 5, there is no setting under which the metrics consistently outperform all other configurations. Identical settings of orientation and transform domain can lead to substantial performance discrepancies across different datasets. In general, the choice of orientation and transform domains requires detailed configurations to achieve the best performance.
(a) JAFFE (different orientations)
(b) USPS (different orientations)
(c) JAFFE (different $\mathbf{M}$)
(d) USPS (different $\mathbf{M}$)
Figure 10: The clustering ACC of STPCA-MP with different data tensor orientations and $\mathbf{M}$. Different data tensor orientations. ’Dir1’ means $D^{o}=\{1,3,2\}$. ’Dir2’ means $D^{o}=\{2,3,1\}$. ’I’ means $\mathbf{M}=\mathbf{I}$. ’DFT’ means that $\mathbf{M}$ is the discrete Fourier matrix (we conduct fast Fourier transform in practice). ’EigVector’ means $\mathbf{M}$ is the eigenvectors decomposed from the self-correlation matrix of the other feature mode. ’Random’ means $\mathbf{M}$ is a random orthogonal matrix. (c) and (d) are obtained by running STPCA-MP-Dir1.
Figure 11: Visualized feature selection results of our proposed method with different data orientations. The experimental condition is the same as in Figure 6.
Table 5: The best ACC and mean POC of our proposed method with different data orientations and transform domains. The best result on each dataset is boldfaced, while the second best one is underlined and italic.
Setting Metric JAFFE BreastMNIST USPS PIE Metric UCIDSA UMCS
Dir1-I ACC 35.29 64.34 63.32 42.84 POC 92.27 41.07
Dir1-DFT ACC 31.35 55.13 64.13 28.20 POC 9.33 38.67
Dir1-EigVector ACC 30.49 57.56 63.7 30.98 POC 92.8 44.80
Dir2-I ACC 30.74 62.44 66.77 36.87 POC 65.07 28.8
Dir2-DFT ACC 28.48 62.29 57.34 38.39 POC 24 16.53
Dir2-EigVector ACC 22.68 65.32 61.55 33.02 POC 45.87 18.13
6 Conclusion
In this paper, we propose a novel Sparse Tensor PCA method named STPCA-MP by combining the encoder-decoder structure of SPCA and the orientation-dependent framework of T-SVDM. We deduce a convex version of STPCA-MP that utilizes the sparse reconstruction tensor to perform UFS. An efficient slice-by-slice algorithm to conduct STPCA-MP is proposed to guarantee fast convergence. The experimental results on real-world data show that our method outperforms state-of-the-art methods on average and can handle different data organization scenarios at a smaller training time cost. In particular, STPCA-MP has the advantage of selecting discriminative features with the greatest BCV on slice-wise data. In addition, we also test the parameter sensitivity, examine the efficient convergence, and discuss the impacts of choosing different data orientations and transform domains. The experimental results show that with a suitable combination of data orientation and transform domain, our method can capture abundant discriminative information and achieve remarkable performance.
There are two valuable research points concerning STPCA-MP in the future. First, the best choice of transform domain is possibly related to the background knowledge of data. More experiments should be carried out for further applications. Second, the extension to higher-order scenarios of STPCA-MP still lacks an effective and efficient optimization algorithm. We will be working on addressing the above two problems in the future.
Appendix A Appendix
[…]
pf3Proof of Theorem 1
{pf3}
To simplify the deduction, we first rewrite the $i$-th problem with compact symbols:
$\min\limits_{\mathbf{U},\mathbf{Q}}$ ${\|\mathbf{X}-\mathbf{U}\mathbf{Q}^{H}\mathbf{X}\|}^{2}_{F}+\lambda{\|\mathbf{Q}\|}_{2,1},$ (39)
$s.t.$ $\mathbf{U}^{H}\mathbf{U}=\mathbf{I}_{k}.$
Thus, we can rewrite the objection function as:
${\rm Tr}\left(\mathbf{X}\mathbf{X}^{H}\right)-\sum\limits^{k}_{j=1}\bigg{[}2Tr\left(\mathbf{u}_{j}^{H}\mathbf{X}\mathbf{X}^{H}\mathbf{q}_{j}\right)-\frac{\lambda}{\|\mathbf{q}_{j}\|_{2}}\mathbf{q}_{j}^{H}\mathbf{q}_{j}-{\rm Tr}\left(\mathbf{q}_{j}^{H}\mathbf{X}\mathbf{X}^{H}\mathbf{q}_{j}\right)\bigg{]}$ (40)
$={\rm Tr}\left(\mathbf{X}\mathbf{X}^{H}\right)-\sum\limits^{k}_{j=1}\bigg{(}2\mathbf{u}_{j}^{H}\mathbf{X}\mathbf{X}^{H}\mathbf{q}_{j}-\frac{\lambda}{\|\mathbf{q}_{j}\|_{2}}\mathbf{q}_{j}^{H}\mathbf{q}_{j}-\mathbf{q}_{j}^{H}\mathbf{X}\mathbf{X}^{H}\mathbf{q}_{j}\bigg{)}$
if we view the above problem as a sum of $k$ subproblems with respect to $\mathbf{u}_{j}$ and $\mathbf{q}_{j}$. Then in each iteration, given a fixed $\mathbf{u}_{j}$, we can have each subproblem minimized at
$\mathbf{q}^{*}_{opt}=\left(\mathbf{X}^{*}_{L_{e}}\mathbf{X}^{T}_{L_{e}}+\frac{\lambda}{\|\mathbf{q}_{j}\|_{2}}\mathbf{I}_{q}\right)^{-1}\mathbf{X}^{*}\mathbf{X}^{T}\mathbf{u}_{j}^{*}.$ (41)
By substituting (41) back to (39) we can obtain
$\mathbf{u}_{opt}=$ (42)
${\underset{\mathbf{u}^{H}\mathbf{u}=1}{\arg\min}}-\mathbf{u}\mathbf{X}\mathbf{X}^{H}\left(\mathbf{X}\mathbf{X}^{H}+\frac{\lambda}{\|\mathbf{q}\|_{2}}\mathbf{I}_{q}\right)^{-1}\mathbf{X}\mathbf{X}^{H}\mathbf{u}^{H}.$
Given the singular decomposition $\mathbf{X}=\mathbf{V}\mathbf{D}\mathbf{M}^{H}$, we have $\mathbf{u}^{opt}_{j}=s_{j}\mathbf{v}_{j}$ where $s_{j}=1$ or $-1$. Then, we obtain $\mathbf{q}^{opt}_{j}=s_{j}d_{jj}/(\|d_{jj}\|^{2}_{2}+{\lambda}/{\|\mathbf{q}_{j}\|_{2}})\mathbf{v}_{j}$.
Finally, we have
$\mathbf{U}_{opt}\mathbf{Q}^{H}_{opt}$ $=\sum\limits^{k}_{j=1}\mathbf{u}_{j}^{opt}(\mathbf{q}_{j}^{opt})^{H}=\mathbf{V}\mathbf{\Sigma}\mathbf{V}^{H}\in H^{q}_{+},$ (43)
[…]
Eventually, after each subproblem of (15) is solved, the optimal solution satisfies $\hat{\mathcal{A}}^{opt}_{::i}=\hat{\mathcal{U}}^{opt}_{::i}(\hat{\mathcal{Q}}^{opt}_{::i})^{H}\in H^{q}_{+}$.
[…]

### v3

@@
 To solve the above problem, we utilize the orientation-dependent tensor-tensor product from Tensor Singular Value Decomposition based on $\star_{\mathbf{M}}$-product (T-SVDM) and extend the one-dimensional Sparse Principal Component Analysis (SPCA) to a tensor form.
-The proposed sparse tensor PCA model can constrain sparsity at the specified mode and yield sparse tensor principal components, enhancing flexibility and accuracy in learning feature relations.
+The proposed sparse tensor PCA model can constrain sparsity at the specified mode and yield sparse tensor principal components, which enhances flexibility and accuracy in learning feature relations.
 To ensure fast convergence and a flexible description of feature correlation, we develop a convex version specially designed for general UFS tasks and propose an efficient slice-by-slice algorithm that performs dual optimization in the transform domain.
-Experimental results on real-world datasets demonstrate the effectiveness and remarkable computational efficiency of the proposed method for tensor data of diverse structures over the state-of-the-arts.
-With a proper combination of data orientation and transform domain, our method is promising for various applications.
-The codes related to our proposed methods and the experiments are available at https://github.com/zjj20212035/STPCA.git.
+Experimental results on real-world datasets demonstrate the effectiveness and remarkable computational efficiency of the proposed method for tensor data of diverse structures over the state-of-the-art.
+When transform axes align with feature distribution patterns, our method is promising for various applications.
+The codes related to our proposed method and the experiments are available at https://github.com/zjj20212035/STPCA.git.
 keywords:
@@
 1 Introduction
-Feature Selection (FS) is a technique of selecting as few features (variables) as possible to meet the demands of a certain task [1, 2].
-It is usually employed as a pre-processing algorithm to relieve the computational burden and enhance interpretability by offering only the necessary features.
-According to the amount of supervised information during training, FS methods can be classified as supervised [3], semi-supervised [4], and unsupervised [5].
-Among them, unsupervised feature selection (UFS) is more challenging due to the unavailability of labels.
-To overcome the difficulty, most UFS methods adopt learning methods including but not limited to principal component analysis (PCA) [38], spectral clustering [40], and graph-based learning [18].
-In this paper, we focus on PCA-based UFS methods.
-In most UFS methods, data samples are vectorized, resulting in the curse of dimensionality [6] as well as data structure loss when dealing with tensors.
-As a natural extension of matrix to store multi-way data samples, tensor is now a common processed object in various communities such as image processing [7], signal processing [8], and economical analysis[9].
-A $n$-th order tensor comprises $n$ modes, each containing several dimensions.
-The tensor structure contains rich information on the relations of modes and dimensions, which is beneficial for analyzing the features.
-In this paper, we focus on leveraging the tensor structure to reduce computational complexity without compromising UFS performance.
-To utilize the structure information, tensor-based UFS methods apply all kinds of Tensor Decomposition (TD) techniques[18, 19, 20].
-Tensor decomposition aims to factorize a tensor into a product of several subtensors, which has been considerably studied for decades since the famous Tucker decomposition [21] was proposed in 1964.
-As different definitions for tensor rank are established, various TD methods have been put forward, such as Higher-order Singular Value Decomposition (HOSVD) [12], CANDECOMP/PARAFAC decomposition (CPD) [22, 23], and Tensor Singular Value Decomposition (T-SVD)[24].
-One purpose of TD is to mine the relations between modes and dimensions, which are naturally related to UFS.
-TD methods preserve the tensor structure of UFS, while UFS helps validate the accuracy and interpretability of the relationships learned by the TD method.
-By minimizing the reconstruction error between a constructed tensor and its decomposition, tensor-based UFS methods successfully preserve the data structure information.
-However, most of these methods do not develop sparse TD and thus use the classifier matrix to perform UFS, which is not beneficial for interpretability and utilization of the learned element relations.
-To address this problem, we start by finding the compact representations of the factor matrices.
-As is widely known, TD methods are also viewed as ways to achieve Tensor Principal Component Analysis (TPCA) [12, 25].
-The principal components have the greatest variance among data samples on a global manifold.
-Their directions describe the feature necessity during the projection onto the manifold.
-Therefore, the key to selecting interpretable and discriminative features could be hidden in TPCA.
-In terms of performing linear projection, the current TPCA methods can be roughly divided into two types: Tuckers-based TPCA [12, 26, 27] and T-SVDM-based TPCA [28, 29, 30].
-Tuckers-based TPCA applies direction-unfolding products defined by Tucker decomposition and its variations (including CPD, which can be viewed as a special case of Tucker’s), while T-SVDM-based TPCA is based on the $\star_{\mathbf{M}}$-product defined by T-SVD and its generalization T-SVDM [31].
-The main differences between the above two types of methods lie in the following two aspects: 1) Tucker-based methods conduct tensor-matrix products through tensor matricization in the original domain, using factor matrices, while T-SVDM-based methods perform tensor-tensor products in the transform domain with a factor tensor.
-By applying factor tensors, T-SVDM-based methods obtain significantly more weights assigned for elements in the data tensor, describing feature relations in a more delicate way.
-2) Tucker-based methods obtain identical results regardless of data organization, while T-SVDM-based methods can achieve various decompositions by applying different transform domains and changing tensor orientation.
-The reorientation works as prior knowledge of mode relations and guides the T-SVDM-based methods to further learn the feature relations in a proper direction, especially in some orientation-specific scenarios such as time series.
-Moreover, the transform operation provides T-SVDM-based methods with greater flexibility, allowing their application to various scenarios.
-In a word, T-SVDM-based methods have a unique advantage in handling complex scenarios.
-We are interested in introducing T-SVDM-based TPCA into UFS.
-To achieve this goal, an extension to a sparse framework is required to ensure the preservation of only the necessary feature relations.
-To the best of our knowledge, this framework is still absent.
-This paper develops a novel sparse T-SVDM-based TPCA framework and deduces a convex version for UFS.
-We explore a way of leveraging the tensor structure to reduce computational complexity while preserving satisfying UFS performance.
+Feature selection (FS) aims to select a minimal subset of features (variables) sufficient for a given task [1, 2].
+Typically employed as a pre-processing step, FS reduces computational complexity and enhances interpretability by providing only essential features.
+FS methods are commonly categorized based on the level of supervision during training: supervised [3], semi-supervised [4], and unsupervised [5].
+Among these, Unsupervised Feature Selection (UFS) presents a greater challenge due to the absence of labels.
+To address this, most UFS methods leverage learning techniques, including but not limited to principal component analysis (PCA) [41], spectral clustering [43], and graph-based learning [21].
+This paper specifically focuses on PCA-based UFS methods.
+A common limitation in many UFS approaches is the vectorization of data samples, which can lead to the curse of dimensionality [6] and loss of inherent data structure, particularly when handling tensors.
+Tensors, as natural multi-way extensions of matrices for representing data, are prevalent in diverse fields like image processing [7], signal processing [8], and econometric analysis [9].
+An $n$-th order tensor consists of $n$ modes, each with multiple dimensions, inherently encoding rich relational information between these modes and dimensions that is valuable for feature analysis.
+This paper focuses on exploiting this tensor structure to reduce computational complexity without compromising UFS performance.
+performance.
+To leverage structural information, tensor-based UFS methods employ various Tensor Decomposition (TD) techniques [21, 22, 23].
+Tensor decomposition factorizes a tensor into a product of several subtensors.
+This area has been extensively studied since the seminal Tucker decomposition [24] was introduced in 1964.
+The establishment of different tensor rank definitions has led to the development of diverse TD methods, such as Higher-Order Singular Value Decomposition (HOSVD) [12], CANDECOMP/PARAFAC decomposition (CPD) [25, 26], and Tensor Singular Value Decomposition (T-SVD) [27].
+A key purpose of TD is to uncover the relationships between modes and dimensions, which are intrinsically linked to UFS.
+TD methods preserve the tensor structure crucial for UFS, while UFS provides a framework to validate the accuracy and interpretability of the relationships learned through TD.
+By minimizing the reconstruction error between the original tensor and its decomposed representation, tensor-based UFS methods effectively preserve data structure information.
+However, most existing methods do not inherently impose sparsity constraints within the TD process itself.
+Consequently, they often rely on a separate classifier matrix to perform feature selection, which can obscure the interpretability and limit the direct utilization of the learned element relationships.
+To address this limitation, we propose seeking compact representations of the factor matrices.
+It is well-established that TD methods can also be interpreted as achieving Tensor Principal Component Analysis (TPCA) [12, 28].
+Principal components capture maximal data variance, and their directions indicate the significance of features during projection onto the underlying manifold.
+Therefore, the key to selecting interpretable and discriminative features likely lies in TPCA.
+Recent studies in domains such as electroencephalogram (EEG) analysis [17] underscore the necessity of orientation-aware approaches.
+[15] demonstrates that spatial-only feature representations, which neglect time-frequency couplings, fail to capture discriminative features in multi-channel EEG signals with temporal variability, resulting in 15%-26% silhouette score degradation for feature selection and clustering tasks.
+Consequently, the authors implement feature selection separately across time, frequency, and time-frequency domains, significantly enhancing clustering accuracy.
+In broader applications, TPCA methods can be categorized into three primary types: Tucker-based TPCA [12, 29, 30], CPD-based TPCA, and T-SVDM-based TPCA [31, 32, 33].
+Key distinctions are summarized in Table 1.
+While Tucker-based and CPD-based methods model potential mode correlations, their reliance on vector outer product principles renders them inherently orientation-insensitive [27].
+This limitation manifests in orientation-specific data (e.g., temporal sequences), where Tucker’s and CPD’s decomposition results remain invariant to data orientation.
+In contrast, T-SVDM-based approaches leverage matrix outer products to adapt to orientation and transform domain variations, yielding direction-sensitive outcomes with broader applicability.
+Furthermore, T-SVDM’s slice-by-slice processing (Section 2.1.2) enables slice-wise sparsity, allowing independent optimization of factor matrices.
+This offers greater flexibility for algorithmic design compared to the rank-constrained CPD and Tucker decomposition that produces dense core tensors.
+[34] further highlights T-SVDM’s reduced parameter sensitivity versus Tucker and CPD in image denoising.
+Although T-SVDM-based methods incur higher computational complexity from factor tensor operations, they provide granular element-wise weighting that meticulously characterizes feature relationships.
+These attributes confer unique advantages for complex scenarios.
+Motivated by these strengths, we propose integrating T-SVDM-based TPCA into UFS.
+Achieving this requires extending the framework with sparsity constraints to preserve only essential feature relations.
+To our knowledge, such a framework remains absent in the literature.
+Table 1: Comparative analysis of UFS methods based on different TD approaches.
+Property Tucker-based CPD-based T-SVDM-based
+Basic product vector outer product vector outer product matrix outer product
+Computational Complexity $O(t(d_{1}d_{2}k_{1}k_{2}+d_{1}d_{3}k_{1}k_{3}+d_{2}d_{3}k_{2}k_{3}))^{\ast}$ $O(t(d_{1}d_{2}+d_{1}d_{3}+d_{2}d_{3})k^{2})$ $O\big{(}d_{3}^{2}d_{1}d_{2}+{}$
+$d_{3}\min(d_{1}d_{2}^{2},d_{1}^{2}d_{2})\big{)}$
+Sensitivity to Orientation Isotropic Permutation-invariant Directionally aware
+Sparsity Handling Dense core Rank-constrained Slice-wise sparsity
+Domain Adaptability Original domain Original domain Transform domain
+$t$: iterations. $d_{i}$: mode-$i$ dimensionality (original tensor). $k_{i}$: mode-$i$ dimensionality (core tensor) ($i=1,2,3$).
+This paper proposes a novel sparse T-SVDM-based TPCA framework with a convex formulation for Unsupervised Feature Selection.
+We establish a methodology that leverages tensor structures to reduce computational complexity while maintaining competitive UFS performance.
 Our contributions are summarized as follows:
@@
 2 Related Work
-In this section, we introduce notations used throughout the paper, and research revolving around tensor decomposition, tensor PCA, and UFS.
+This section introduces mathematical notations employed throughout the paper and reviews relevant research on tensor decomposition, tensor principal component analysis, and unsupervised feature selection.
 2.1 Notations
 2.1.1 Basic Symbols
-In most cases, we use capital letters (e.g.
-$A$) to denote sets, except the common ones like integer set $\mathbb{Z}$, the real-number set $\mathbb{R}$ and the complex-number set $\mathbb{C}$.
-In particular, $H^{d}_{+}$ stands for the HPSD Cone for $d\times{d}$ matrices, and $\Phi$ is the empty set.
-For an integer $n$, $[n]$ represents the set $\left[1,2,3,...,n\right]$.
-#$\#(A)$ stands for the number of elements in the set $A$.
-A set with the superscript "o" has ordered elements; for example, $A^{o}=\{1,2,3\}\neq B^{o}=\{1,3,2\}$.
-We use lowercase letters (e.g.
-$a$) to denote scalars, boldface lowercase letters (e.g.
-$\mathbf{a}$) to denote vectors, boldface capital letters (e.g.
-$\mathbf{A}$) and Euler script letters (e.g.
-$\mathcal{A}$) to denote tensors.
-For a $n^{th}$-order tensor $\mathcal{A}\in\mathbb{C}^{d_{1}\times{d_{2}}\times{\cdots}\times{d_{n}}}$, its $(i_{1},i_{2},\cdots,i_{n})$-th element is denoted by $\mathcal{A}_{i_{1}i_{2}\cdots i_{n}}$.
-For a matrix, the corresponding element is denoted by $a_{i_{1}i_{2}}$.
-For a matrix $\mathbf{A}$, its row vectors are denoted by $\mathbf{A}_{j,:}$ or boldface lowercase letters $\mathbf{a}^{j}$.
-The column vectors are denoted by $\mathbf{A}_{:,j}$ or boldface lowercase letters $\mathbf{a}_{j}$.
-There are two types of common sections for third-order tensors: fibers and slices.
-The vector along the $n$-th mode is referred to as an $n$-mode fiber.
-A third-order tensor $\mathcal{A}\in\mathbb{C}^{d_{1}\times{d_{2}}\times{d_{3}}}$ has column (1-mode) fibers $\mathcal{A}_{:i_{2}i_{3}}$, row (2-mode) fibers $\mathcal{A}_{i_{1}:i_{3}}$, and tube (3-mode) fibers $\mathcal{A}_{i_{1}i_{2}:}$.
-A slice is a matrix obtained by fixing all but two indices in a tensor.
-For a third-order tensor, slices can be divided into frontal slices $\mathcal{A}_{::i_{3}}$, lateral slices $\mathcal{A}_{:i_{2}:}$ and horizontal slices $\mathcal{A}_{i_{1}::}$.
-The transpose, conjugation, and conjugate transpose of a matrix are denoted by $\mathbf{A}^{T}$, $\mathbf{A}^{*}$, and $\mathbf{A}^{H}$, respectively.
-We denote the trace of a square matrix by ${\rm Tr}(\mathbf{A})$.
-We denote Frobenius norm, nuclear norm and $\ell_{2,1}$ norm by ${\|\mathcal{A}\|}_{F}$,${\|\mathbf{A}\|}_{*}$, and ${\|\mathbf{A}\|}_{2,1}$ respectively.
-The nuclear norm of a matrix is defined as the sum of its singular values.
-The $\ell_{2,1}$-norm is obtained by calculating the $\ell_{2}$-norm of each column vector in a matrix and summing the results.
-The Kronecker product is denoted by $\bigotimes$.
+Notations employed throughout the paper are summarized in Table 2.
+Specifically, fibers correspond to vector sections along individual tensor modes, while slices represent matrix sections defined by fixing all but two modes.
+Table 2: Summary of mathematical notations. The following notations use the letter A as an example, and they are equally applicable to other letters.
+Symbol Description
+$A$ Generic set
+$\mathbb{Z},\mathbb{R},\mathbb{C}$ Integer, real, complex number sets
+$H^{d}_{+}$ Hermitian Positive Semidefinite (HPSD) cone ($d\times d$ matrices)
+$\Phi$ Empty set
+$[n]$ Ordered set $\{1,2,\dots,n\}$
+$A^{o}$ Ordered set (element order matters)
+#$\#(A)$ Cardinality of set $A$
+$a$ Scalar
+$\mathbf{a}$ Vector
+$\mathbf{A}$ Matrix
+$\mathcal{A}$ Third-order Tensor
+$\mathcal{A}_{:i_{2}i_{3}}$, $\mathcal{A}_{i_{1}:i_{3}}$, $\mathcal{A}_{i_{1}i_{2}:}$ Column (mode-1), Row (mode-2), and Tube (mode-3) fibers
+$\mathcal{A}_{::i_{3}}$, $\mathcal{A}_{:i_{2}:}$, $\mathcal{A}_{i_{1}::}$ Frontal, Lateral and Horizontal slices
+$\mathcal{A}_{i_{1}\cdots i_{n}}$ Tensor element at $(i_{1},\dots,i_{n})$
+$a_{ij}$ Matrix element at $(i,j)$
+$\mathbf{A}_{j,:},\mathbf{a}^{j}$ $j$-th row vector of matrix
+$\mathbf{A}_{:,j},\mathbf{a}_{j}$ $j$-th column vector of matrix
+$\mathbf{A}^{T},\mathbf{A}^{*},\mathbf{A}^{H}$ Transpose, conjugate, conjugate transpose
+$\mathrm{Tr}(\mathbf{A})$ Matrix trace
+$\bigotimes$ Kronecker product
+$\|\mathcal{A}\|_{F}$ Frobenius norm ($\sqrt{\sum_{i_{1}}\sum_{i_{2}}\sum_{i_{3}}\mathcal{A}^{2}_{i_{1}i_{2}i_{3}}}$)
+$\|\mathbf{A}\|_{*}$ Nuclear norm ($\sum\sigma_{i}^{\ast}$)
+$\|\mathbf{A}\|_{2,1}$ $\ell_{2,1}$ norm ($\sum_{j}\|\mathbf{A}_{:,j}\|_{2}$)
+singular values
 2.1.2 Tensor Operations
-In this section, we introduce several tensor operations.
+In this section, we introduce fundamental tensor operations.
 a) Tensor Unfolding and Tensor-Matrix Product
-We use $n$-mode unfolding to perform tensor matricization.
-A $m$-th order tensor $\mathcal{A}^{d_{1}\times{d_{2}}\dots\times{d_{m}}}$ can be transformed into $\mathbf{X}_{(n)}$ along the $n$-th mode where
+Tensor unfolding (matricization) is performed via $n$-mode unfolding.
+For an $m$-th order tensor $\mathcal{X}\in\mathbb{R}^{d_{1}\times d_{2}\times\cdots\times d_{m}}$, its $n$-mode unfolding $\mathbf{X}_{(n)}$ is defined such that:
 […]
-Using the unfolding technique, the tensor-matrix product is defined as
+Based on this unfolding operation, the tensor-matrix product ($n$-mode product) is defined as:
 $\mathcal{Y}=\mathcal{X}\times_{n}\mathbf{U}\iff\mathbf{Y}_{(n)}=\mathbf{U}\mathbf{X}_{(n)},$ (2)
-which is also known as the $n$-mode product.
 b) $\star_{\mathbf{M}}$-product
-In the T-SVDM framework, third-order tensor products are achieved through the $\star_{\mathbf{M}}$-product.
-Given a third-order tensor $\mathcal{A}\in\mathbb{C}^{d\times{m}\times{n}}$, we first denote a tensor in the transform domain specified by an invertible matrix $\mathbf{M}\in\mathbb{C}^{n\times{n}}$:
+Within the T-SVDM framework, the $\star_{\mathbf{M}}$-product defines tensor operations for third-order tensors.
+Consider $\mathcal{A}\in\mathbb{C}^{d\times{m}\times{n}}$ and an invertible transformation matrix $\mathbf{M}\in\mathbb{C}^{n\times{n}}$.
+We first define the transformed tensor in the $\mathbf{M}$-domain as:
 $\hat{\mathcal{A}}:=\mathcal{A}\times_{3}\mathbf{M}.$ (3)
-Then, given another tensor $\mathcal{B}\in\mathbb{C}^{m\times{l}\times{n}}$, $\mathcal{A}\star_{\mathbf{M}}\mathcal{B}=\mathcal{C}$ is defined as
-$\hat{\mathcal{C}}_{::i}$ $=\hat{\mathcal{A}}_{::i}\hat{\mathcal{B}}_{::i},\quad i=1,2,\cdots,n,$ (4)
+For another tensor $\mathcal{B}\in\mathbb{C}^{m\times{l}\times{n}}$, the $\star_{\mathbf{M}}$-product $\mathcal{C}=\mathcal{A}\star_{\mathbf{M}}\mathcal{B}$ is given by:
+$\hat{\mathcal{C}}_{::i}$ $=\hat{\mathcal{A}}_{::i}\hat{\mathcal{B}}_{::i},\quad\forall i\in\{1,2,\cdots,n\},$ (4)
 $\mathcal{C}$ $=\hat{\mathcal{C}}\times_{3}\mathbf{M}^{-1}\in{\mathbb{C}^{d\times{l}\times{n}}}.$
-Hereafter, we use $\hat{\mathcal{A}}$ to denote the result of $\mathcal{A}\times_{3}\mathbf{M}$.
-The $\star_{\mathbf{M}}$-product is a slice-by-slice operation and is thus orientation-dependent.
-Different orientations of the tensor will lead to different results.
-The identity tensor $\mathcal{I}$ under $\star_{\mathbf{M}}$ satisfies the following condition:
+Throughout this work, $\hat{\mathcal{A}}$ consistently denotes $\mathcal{A}\times_{3}\mathbf{M}$.
+The $\star_{\mathbf{M}}$-product performs slice-by-slice matrix multiplications to achieve tensor-tensor products in the transform domain, making it inherently orientation-sensitive.
+Consequently, different tensor orientations yield distinct results.
+The identity tensor $\mathcal{I}$ under the $\star_{\mathbf{M}}$-product satisfies:
 $\mathcal{A}\star_{\mathbf{M}}\mathcal{I}=\mathcal{I}\star_{\mathbf{M}}\mathcal{A}^{H}=\mathcal{A}.$ (5)
-Each frontal slice of $\hat{\mathcal{I}}$ is an identity matrix $\mathbf{I}$.
+In the transform domain, each frontal slice of $\hat{\mathcal{I}}$ corresponds to an identity matrix $\mathbf{I}$.
 The conjugate transpose of $\mathcal{A}\in\mathbb{C}^{d\times{m}\times{n}}$ under $\star_{\mathbf{M}}$ is defined as:
-$\left(\hat{\mathcal{A}}^{H}\right)_{::i}=\left(\hat{\mathcal{A}}_{::i}\right)^{H},\quad i=1,\cdots,n,$ (6)
-which makes sure that $\mathcal{A}^{H}\star_{\mathbf{M}}\mathcal{B}^{H}=\left(\mathcal{B}\star_{\mathbf{M}}\mathcal{A}\right)^{H}$.
-We can further define a $\star_{\mathbf{M}}$-orthogonal tensor $\mathcal{A}\in{\mathbb{C}^{m\times{m}\times{n}}}$ if
+$\left(\hat{\mathcal{A}}^{H}\right)_{::i}=\left(\hat{\mathcal{A}}_{::i}\right)^{H},\quad\forall i\in\{1,\cdots,n\},$ (6)
+which ensures the fundamental property:
+$\mathcal{A}^{H}\star_{\mathbf{M}}\mathcal{B}^{H}=\left(\mathcal{B}\star_{\mathbf{M}}\mathcal{A}\right)^{H}.$ (7)
+A tensor $\mathcal{A}\in\mathbb{C}^{m\times{m}\times{n}}$ is $\star_{\mathbf{M}}$-orthogonal if it satisfies:
 $\mathcal{A}^{H}\star_{\mathbf{M}}\mathcal{A}=\mathcal{A}\star_{\mathbf{M}}\mathcal{A}^{H}=\mathcal{I}.$ (8)
 2.2 T-SVDM
-Using the $\star_{\mathbf{M}}$ product, T-SVDM [31] decomposes a third-order tensor $\mathcal{A}\in\mathbb{C}^{d_{1}\times{d_{2}}\times{d_{3}}}$ into:
+Within the $\star_{\mathbf{M}}$-product framework, the T-SVDM decomposition [34] factorizes a third-order tensor $\mathcal{A}\in\mathbb{C}^{d_{1}\times d_{2}\times d_{3}}$ as:
 $\mathcal{A}=\mathcal{U}\star_{\mathbf{M}}\mathcal{S}\star_{\mathbf{M}}\mathcal{V}^{H},$ (9)
-where $\mathcal{U}\in\mathbb{C}^{d_{1}\times d_{1}\times d_{3}}$, $\mathcal{V}\in\mathbb{C}^{d_{2}\times d_{2}\times d_{3}}$ are $\star_{\mathbf{M}}-unitary$, and $\mathcal{S}\in\mathbb{C}^{d_{1}\times{d_{2}}\times{d_{3}}}$ is a f-diagonal tensor whose frontal slices are diagonal.
-The decomposition is called T-SVD [24] when $\mathbf{M}$ is the discrete Fourier transform (DFT) matrix and T-SVDM [31] in a usual case.
-[28] was the first to establish the relationship between T-SVD and tensor PCA.
-It applies T-SVD to compress the human-face images with a pre-rotation of the data tensor.
-[31] further extends this idea to T-SVDM.
-[29] highlights the importance of leveraging the orientation-dependence of T-SVDs to extract information from different directions.
-Inspired by these works, we derive a sparse model for tensors combining T-SVDM and SPCA.
+where $\mathcal{U}\in\mathbb{C}^{d_{1}\times d_{1}\times d_{3}}$ and $\mathcal{V}\in\mathbb{C}^{d_{2}\times d_{2}\times d_{3}}$ are $\star_{\mathbf{M}}$-unitary tensors.
+$\mathcal{S}\in\mathbb{C}^{d_{1}\times d_{2}\times d_{3}}$ is an f-diagonal tensor with diagonal frontal slices.
+This decomposition generalizes two important cases: T-SVD [27]: Special case when $\mathbf{M}$ is the DFT matrix; T-SVDM [34]: General case with arbitrary invertible $\mathbf{M}$.
+The connection between T-SVDM and TPCA was first established in [31], which applied T-SVD to facial image compression through data tensor pre-rotation.
+This work was later extended to T-SVDM in [34].
+Notably, [32] demonstrated the critical role of orientation-dependence in T-SVD for multi-directional feature extraction.
+Building upon these foundations, we develop a novel sparse tensor model that integrates T-SVDM with sparse PCA.
 2.3 PCA and Sparse PCA (SPCA)
-Principal Component Analysis (PCA) [47] is a well-known unsupervised feature extraction method.
-It aims to find an orthogonal linear transformation matrix that projects the original data vectors into a lower-dimensional manifold, maximizing the variance of the projected samples.
-The optimization problem of PCA can be formulated as the following self-contained regression form:
+Principal Component Analysis (PCA) [50] is a fundamental unsupervised feature extraction method.
+It seeks an orthogonal linear transformation that projects high-dimensional data onto a lower-dimensional manifold while maximizing the variance of the projected samples.
+The PCA optimization problem can be expressed in the following self-contained regression form:
 $\min\limits_{\mathbf{U}^{T}\mathbf{U}=\mathbf{I}_{k}}{\|{\mathbf{X}-\mathbf{U}\mathbf{U}^{T}\mathbf{X}}\|}^{2}_{F},$ (10)
@@
 $\mathbf{U}\in\mathbb{R}^{d\times{k}}$ ($k\ll d$) is an orthogonal transformation matrix.
-To obtain sparse weights and improve interpretability, Sparse PCA (SPCA) [32] was proposed by adding ridge and lasso constraints to PCA:
+To enhance interpretability through sparse feature weights, Sparse PCA (SPCA) [35] introduces ridge and lasso constraints to the PCA framework:
 $\min\limits_{\mathbf{U},\mathbf{Q}}$ ${\|{\mathbf{X}-\mathbf{U}\mathbf{Q}^{T}\mathbf{X}}\|}^{2}_{F}+\alpha\sum\limits_{j=1}^{k}{\|{\mathbf{q}_{j}}\|}^{2}_{2}+\beta\sum\limits_{j=1}^{k}{\|{\mathbf{q}_{j}}\|}_{1}$ (11)
@@
 where $\alpha,\beta>0$ are regularization parameters.
-The self-contained regression term in SPCA incorporates an encoder-decoder architecture, selectively transforming a sparse subset of input features into encoded latent representations and reconstructing them through an orthogonal decoder, as Figure 2 shows.
-This approach gains flexibility by relaxing the transformation matrix $\mathbf{Q}$.
-When $\mathbf{Q}=\mathbf{U}$, SPCA reduces to PCA.
+The self-contained regression term in SPCA implements an encoder-decoder architecture:
+  • Encoder ($\mathbf{Q}$): Selectively transforms a sparse subset of input features into latent representations
+  • Decoder ($\mathbf{U}$): Reconstructs original data through an orthogonal transformation
+as illustrated in Figure 2.
+This formulation enhances flexibility by decoupling the transformation matrix $\mathbf{Q}$ from $\mathbf{U}$.
+Notably, when $\mathbf{Q}=\mathbf{U}$, SPCA reduces to standard PCA.
 Figure 2: The encoder-decoder principle of one-dimensional SPCA
-By minimizing reconstruction errors, SPCA preserves the global data structure in an unsupervised manner, maintaining essential geometric relationships within the data manifold.
-The lasso constraint enables SPCA to assign weights only to the necessary features.
-The orthogonality constraint ensures that the latent features are complementary and reduces redundancy.
-Therefore, SPCA is well-suited for UFS tasks, particularly in balancing the clustering performance and feature interpretability.
+Through the minimization of reconstruction errors, SPCA preserves global data structures in an unsupervised manner while maintaining essential geometric relationships within the data manifold.
+The lasso constraint enables selective feature weighting by assigning non-zero weights only to necessary features.
+Simultaneously, the orthogonality constraint ensures complementarity of latent features and reduces redundancy.
+These properties make SPCA particularly well-suited for unsupervised feature selection (UFS), effectively balancing clustering performance and feature interpretability.
 2.4 Tensor-based unsupervised feature selection
-Tensor-based unsupervised feature selection remains an emerging research area.
-Most of the existing methods rely on domain knowledge and are limited to certain application scenarios.
-For example, [20] proposes an image-transformation-based feature selection method, aiming at vehicle detection.
-[34] constructs a multi-omics data tensor and exploits feature selection criterion designed using the knowledge in bioinformatics.
-Recently, some methods designed for general cases have emerged.
-GRLTR [35] combines graph regularization with tensor tubal rank constraint to simultaneously preserve global and local information.
-However, it still requires vectorization.
-CPUFS [18] derives a model based on CPD and effectively preserves the tensor structure by minimizing the reconstruction error.
-It represents a promising approach to tensor-based UFS.
-However, CPUFS does not derive sparse factor matrices.
-Instead, it applies spectral clustering and utilizes a sparse linear classifier matrix to perform UFS, leading to insufficient utilization of the information in factor matrices and reduced interpretability of selected features.
-Additionally, the constructed graph in CPUFS needs the number of classes as prior knowledge and can be unreliable due to noise features.
-Recently, developing sparse PCA models for UFS appears to be a new trend [36, 37, 38, 39].
-PCA-like methods do not rely on the construction of graphs.
-By global manifold learning, they can preserve the data structure and achieve superior average performance.
-What’s more, they utilize the factor matrix that maximizes the variance of data samples in the projected space, whose sparse weights demonstrate a clearer relevance of the original features and have better interpretability.
-Unfortunately, existing PCA-like methods are non-tensor-based.
-Motivated by the above considerations, this study develops a sparse TPCA-based methodology tailored for UFS tasks.
+Tensor-based unsupervised feature selection (UFS) remains an emerging research domain where most existing methods rely heavily on domain-specific knowledge, limiting their applicability to particular scenarios.
+For instance, [23] proposes an image-transformation-based feature selection method specialized for vehicle detection, while [37] constructs multi-omics data tensors with feature selection criteria designed using bioinformatics knowledge.
+Although recent methods targeting general applications have emerged—such as GRLTR [38] which combines graph regularization with tensor tubal rank constraints to preserve global and local information—they still require vectorization.
+CPUFS [21] represents a promising tensor-based approach that preserves tensor structures via CP decomposition reconstruction, but it fails to derive sparse factor matrices.
+Instead, CPUFS applies spectral clustering and utilizes a sparse linear classifier matrix for feature selection, leading to insufficient utilization of factor matrix information and reduced interpretability.
+Furthermore, its graph construction requires prior knowledge of class numbers and remains vulnerable to noisy features.
+Recently, sparse PCA models constitute a growing trend in UFS [39, 40, 41, 42] due to their ability to avoid graph construction dependencies, preserve data structures through global manifold learning, deliver superior average performance, and provide interpretable sparse weights via variance-maximizing factor matrices.
+However, existing PCA-based methods lack tensor formulations, limiting their applicability to higher-dimensional data.
+To bridge this gap, we develop a novel sparse TPCA-based methodology specifically designed for tensor UFS tasks.
 3 Methodology
-In this section, we extend the framework of SPCA to tensor form by utilizing techniques from T-SVDM and derive a model called Sparse Tensor PCA based on $\star_{\mathbf{M}}$ product (STPCA-MP).
-Then, we deduce a convex version of STPCA-MP for UFS tasks by merging the encoder and decoder tensors into a single sparse reconstruction tensor.
-We propose an efficient slice-by-slice optimization algorithm that individually solves each dual subproblem in the transform domain.
-Eventually, by considering both the transform matrix $\mathbf{M}$ and the learned reconstruction tensor, we compute a feature score map and perform feature selection.
-The overall framework is illustrated in Figure 1.
+This section extends the SPCA framework to tensor formulation by integrating T-SVDM techniques, establishing a novel Sparse Tensor PCA based on $\star_{\mathbf{M}}$ product (STPCA-MP) model.
+We subsequently derive a convex variant of STPCA-MP for UFS applications by unifying the encoder and decoder tensors into a single sparse reconstruction tensor.
+To optimize this model efficiently, we develop a slice-by-slice algorithm that independently solves each dual subproblem in the transform domain.
+Finally, leveraging both the transformation matrix $\mathbf{M}$ and the learned reconstruction tensor, we compute feature score maps to perform feature selection.
+Figure 1 illustrates the complete framework, with key mathematical symbols summarized in Table 3.
+Table 3: Symbols used throughout the deduction.
+Notation Description
+$\mathcal{X}\in\mathbb{C}^{d_{1}\times d_{2}\times n}$ Data tensor containing $n$ samples
+$\hat{\mathcal{X}}\in\mathbb{C}^{d_{1}\times d_{2}\times n}$ Data tensor in the transform domain
+$\tilde{\mathcal{X}}\in\mathbb{C}^{d_{1}\times d_{2}\times n}$ Reconstructed data tensor
+$D^{o}\in\mathbb{Z}^{3}$ Order set defining the order of modes
+$\mathcal{X}_{D^{o}}\in\mathbb{C}^{q\times n\times p}$ Data tensor permuted by $D^{o}$
+$\mathcal{U}$, $\hat{\mathcal{U}}\in\mathbb{C}^{q\times k\times p}$ Decoder tensor for $\mathcal{X}_{D^{o}}$ in the original and the transform domain
+$\mathcal{Q}$, $\hat{\mathcal{Q}}\in\mathbb{C}^{q\times k\times p}$ Encoder tensor for $\mathcal{X}_{D^{o}}$ in the original and the transform domain
+$\mathcal{Q}_{::i}$ The $i$-th frontal slice of $\mathcal{Q}$
+$\mathbf{Q}_{(1)}$ 1-mode unfolding of $\mathcal{Q}$
+$\left(\mathbf{Q}_{(1)}\right)_{:j}$ The $j$-th column vector in $\mathbf{Q}_{(1)}$
+$\mathcal{A}$, $\hat{\mathcal{A}}\in\mathbb{C}^{q\times q\times p}$ Reconstruction tensor for $\mathcal{X}_{D^{o}}$ in the original and the transform domain
+$\mathcal{I}\in\mathbb{C}^{q\times q\times p}$ Identity tensor under $\star_{\mathbf{M}}$ product
+$\alpha$ Regularization parameter of lasso regression
+$\beta$ Regularization parameter of ridge regression
+$\lambda$ Regularization parameter of $\ell_{2,1}$-norm
+$\eta$ Regularization parameter of trace function
 3.1 Model formulation of STPCA-MP
-We formulate the model in the third-order tensor scenario for the simplicity of deduction and algorithm design.
-The principle can be extended to high-order scenarios in a recursive fashion.
-See Section 4 for details.
-Given a data tensor $\mathcal{X}\in\mathbb{C}^{d_{1}\times d_{2}\times n}$ containing $n$ samples of size $d_{1}\times d_{2}$, we can reorient it with an order set $D^{o}$ such that the second mode represents the sample dimension.
-We denote the rotated data tensor by $\mathcal{X}_{D^{o}}\in\mathbb{C}^{q\times n\times p}$.
-Based on (8), the row dimensionality of each frontal slice in $\mathcal{X}_{D^{o}}$ can be reduced with $\mathcal{U}^{H}\in\mathbb{C}^{k\times q\times p}(k<d)$, followed by reconstruction with $\mathcal{U}$:
+We formulate the model for third-order tensors to simplify deduction and algorithmic development, noting that the principle extends recursively to higher-order scenarios (see Section 4 for generalization details).
+Consider a data tensor $\mathcal{X}\in\mathbb{C}^{d_{1}\times d_{2}\times n}$ containing $n$ samples of dimensions $d_{1}\times d_{2}$.
+Through reorientation via an ordered set $D^{o}$ where the second mode corresponds to the sample dimension, we obtain the rotated tensor $\mathcal{X}_{D^{o}}\in\mathbb{C}^{q\times n\times p}$.
+Note that following rotation, each frontal slice of $\mathcal{X}_{D^{o}}$ undergoes sample-wise mean centering along its second mode (sample dimension).
+Building on the T-SVDM framework (9), we reduce row dimensionality of each frontal slice in $\mathcal{X}_{D^{o}}$ using $\mathcal{U}^{H}\in\mathbb{C}^{k\times q\times p}$ ($k<d$), followed by reconstruction with $\mathcal{U}$:
 $\mathcal{Y}$ $=\mathcal{U}^{H}\star_{\mathbf{M}}\mathcal{X}_{D^{o}},$ (12)
@@
 where $\mathcal{Y}\in\mathbb{C}^{k\times n\times p}$, and $\tilde{\mathcal{X}}_{D^{o}}\in\mathbb{C}^{q\times n\times p}$ is the reconstructed tensor.
-Intuitively, we can first formulate a tensor PCA model following the idea of one-dimensional PCA in (9):
-$\underset{\mathcal{U}}{\min}$ ${\|\mathcal{X}_{D^{o}}-\mathcal{U}\star_{\mathbf{M}}\mathcal{U}^{H}\star_{\mathbf{M}}\mathcal{X}_{D^{o}}\|^{2}_{F}}$ (12)
-$s.t.$ $\quad\mathcal{U}^{H}\star_{\mathbf{M}}\mathcal{U}=\mathcal{I}.$
-Then, inspired by the encoder-decoder structure of SPCA, we use another tensor $\mathcal{Q}\in\mathbb{C}^{d\times k\times p}$ to perform feature extraction instead of $\mathcal{U}$.
-By doing so, we obtain a more flexible reconstruction framework that yields a better latent feature representation:
-$\underset{\mathcal{U}}{\min}$ ${\|\mathcal{X}_{D^{o}}-\mathcal{U}\star_{\mathbf{M}}\mathcal{Q}^{H}\star_{\mathbf{M}}\mathcal{X}_{D^{o}}\|^{2}_{F}}$ (13)
-$s.t.$ $\quad\mathcal{U}^{H}\star_{\mathbf{M}}\mathcal{U}=\mathcal{I}.$
-Specifically, $\mathcal{Q}$ contains the weights that measure the importance of each original feature during the projection into the latent feature space.
-To encode only the necessary original features, $\mathcal{Q}$ is constrained to have sparse loadings.
-Meanwhile, for highly relevant features, we hope the model can treat them equally, instead of preferring certain features, especially when $dm\gg n$.
-Therefore, lasso regression and ridge regression terms are introduced as constraints for $\mathcal{Q}$, resulting in a sparse Tensor PCA model:
-$\underset{\mathcal{U},\mathcal{Q}}{\min}$ ${\|\mathcal{X}_{D^{o}}-\mathcal{U}\star_{\mathbf{M}}\mathcal{Q}^{H}\star_{\mathbf{M}}\mathcal{X}_{D^{o}}\|^{2}_{F}}+\alpha\sum\limits^{pn}_{j=1}{\|\left(\mathbf{Q}_{(1)}\right)_{:j}\|^{2}_{2}}+\beta\sum\limits^{pn}_{j=1}{\|\left(\mathbf{Q}_{(1)}\right)_{:j}\|_{1}}$ (14)
-$s.t.$ $\quad\mathcal{U}^{H}\star_{\mathbf{M}}\mathcal{U}=\mathcal{I}.$
-By obtaining the optimal $\mathcal{Q}$, the sparse tensor principal components can be extracted from the data tensor.
-Discussion: To this end, the proposed model is directly applicable to tasks traditionally addressed by SPCA (e.g., image compression, visualization, and denoising) while preserving the tensor structure.
-However, the encoder tensor alone is suboptimal for UFS tasks.
-In the following section, we reformulate the model to meet UFS-specific requirements.
+Intuitively, we can first formulate a tensor PCA model inheriting the principle of data reconstruction and orthogonality constraint in PCA (Problem (10)):
+$\underset{\mathcal{U}}{\min}\quad{\|\mathcal{X}_{D^{o}}-\mathcal{U}\star_{\mathbf{M}}\mathcal{U}^{H}\star_{\mathbf{M}}\mathcal{X}_{D^{o}}\|^{2}_{F}}$ (13)
+[…]
+which minimizes the data reconstruction error without compromising the tensor structure.
+Building on the encoder-decoder architecture of SPCA [35], we introduce a distinct transformation tensor $\mathcal{Q}\in\mathbb{C}^{q\times k\times p}$ to replace $\mathcal{U}$ for feature extraction.
+This yields a more flexible reconstruction framework that enhances latent feature representation:
+$\underset{\mathcal{U}}{\min}\quad{\|\mathcal{X}_{D^{o}}-\mathcal{U}\star_{\mathbf{M}}\mathcal{Q}^{H}\star_{\mathbf{M}}\mathcal{X}_{D^{o}}\|^{2}_{F}}$ (14)
+[…]
+Similar to SPCA, Problem (14) adopts an encoder-decoder framework:
+  • Tensor encoder ($\mathcal{Q}$): Projects input features into a latent space while preserving the intrinsic tensor structure and orientation relationships.
+  • Tensor decoder ($\mathcal{U}$): Recovers the original tensor through a series of orthogonal transformations, ensuring minimal reconstruction error.
+To enforce feature selection during encoding, we impose sparsity constraints on the transformation tensor $\mathcal{Q}$, ensuring only discriminative features are retained in the latent representation.
+For highly correlated features, particularly in high-dimensional scenarios ($d_{m}\gg n$), we impose equality constraints to prevent model bias toward specific features.
+This is achieved by incorporating lasso and ridge regression terms as $\mathcal{Q}$ constraints, yielding the sparse tensor PCA model:
+$\underset{\mathcal{U},\mathcal{Q}}{\min}\quad{\|\mathcal{X}_{D^{o}}-\mathcal{U}\star_{\mathbf{M}}\mathcal{Q}^{H}\star_{\mathbf{M}}\mathcal{X}_{D^{o}}\|^{2}_{F}}+\alpha\sum\limits^{pn}_{j=1}{\|\left(\mathbf{Q}_{(1)}\right)_{:j}\|^{2}_{2}}+\beta\sum\limits^{pn}_{j=1}{\|\left(\mathbf{Q}_{(1)}\right)_{:j}\|_{1}}$ (15)
+[…]
+which can be viewed as an extension of SPCA to a tensor form. Solving for the optimal $\mathcal{Q}$ enables extraction of sparse tensor principal components from the data tensor.
+Discussion: To this end, the proposed model naturally extends SPCA applications to tensor-structured data, including traditional tasks such as image compression, visualization, and denoising, while preserving intrinsic structural relationships.
+However, the encoder tensor $\mathcal{Q}$ alone proves suboptimal for UFS due to insufficient feature discriminability.
+The following section reformulates the model specifically for UFS requirements.
 3.2 Convex version of STPCA-MP for UFS
-As previously discussed, the self-contained regression term has an encoder-decoder structure.
-Therefore, the importance of each original feature is not only expressed by the weights contained in $\mathcal{Q}$ (encoder) but also described by $\mathcal{U}$ (decoder).
-Comprehensively considering their relationship is essential for adapting the model to UFS.
-Dual Problem: Since the transformation between tensor $\mathcal{A}$ and its counterpart $\hat{\mathcal{A}}$ is uniquely defined by $\hat{\mathcal{A}}=\mathcal{A}\times_{3}\mathbf{M}$, Problem (14) has its dual problem in the transform domain, which means they can be solved simultaneously.
-To facilitate optimization, we can rewrite Problem (14) in the transform domain where matrix algebra can be applied:
-$\underset{\hat{\mathcal{U}},\hat{\mathcal{Q}}}{\min}$ $\sum\limits^{p}_{i=1}{\|\left(\hat{\mathcal{X}}_{D^{o}}\right)_{::i}-\hat{\mathcal{U}}_{::i}\hat{\mathcal{Q}}^{H}_{::i}\left(\hat{\mathcal{X}}_{D^{o}}\right)_{::i}\|^{2}_{F}}+\lambda\sum\limits^{p}_{j=1}{\|\hat{\mathcal{Q}}_{::i}\|_{2,1}}$ (15)
-$s.t.\quad\hat{\mathcal{U}}_{::i}^{H}\left(\hat{\mathcal{U}}_{::i}\right)=\mathbf{I},i\in[p].$
-The $\ell_{2,1}$ norm is employed to simultaneously achieve the goals of lasso regression and ridge regression.
-Variable Merging: Notably, there exists a tensor $\mathcal{A}\in\mathbb{C}^{d\times d}$ whose $i$-th frontal slice in the transform domain $\hat{\mathcal{A}}_{::i}=\hat{\mathcal{U}}_{::i}\hat{\mathcal{Q}}^{H}_{::i}$.
-Each row vector in $\hat{\mathcal{A}}_{::i}$ contains reconstruction weights that combine elements from both $\hat{\mathcal{U}}_{::i}$ and $\hat{\mathcal{Q}}^{H}_{::i}$.
-Therefore, $\mathcal{A}$ is a more suitable feature selection tensor than $\mathcal{Q}$.
-Guided by $D^{o}$, each frontal slice in $\hat{\mathcal{A}}$ reconstructs the fibers of a certain mode in the data tensor, utilizing the structural information hidden in the mode organization.
-Remarkably, it can be proven that the optimal $\hat{\mathcal{A}}$ is a Hermitian positive semidefinite (HPSD) matrix, detailed by Theorem 1.
+As previously established, the self-contained regression term embodies an encoder-decoder architecture.
+Consequently, feature importance is jointly encoded by $\mathcal{Q}$ (encoder) and $\mathcal{U}$ (decoder).
+Comprehensive integration of their relationship is essential for UFS adaptation.
+Dual Problem: Given the bijective mapping $\hat{\mathcal{A}}=\mathcal{A}\times_{3}\mathbf{M}$ between tensors, Problem (15) admits a dual formulation in the transform domain.
+This duality enables simultaneous optimization.
+For computational efficiency, we recast the problem using matrix algebra:
+$\underset{\hat{\mathcal{U}},\hat{\mathcal{Q}}}{\min}\quad\sum\limits^{p}_{i=1}{\|\left(\hat{\mathcal{X}}_{D^{o}}\right)_{::i}-\hat{\mathcal{U}}_{::i}\hat{\mathcal{Q}}^{H}_{::i}\left(\hat{\mathcal{X}}_{D^{o}}\right)_{::i}\|^{2}_{F}}+\lambda\sum\limits^{p}_{j=1}{\|\hat{\mathcal{Q}}_{::i}\|_{2,1}}$ (16)
+[…]
+The $\ell_{2,1}$ norm concurrently satisfies lasso and ridge regression objectives.
+Critically, Problem (16) demonstrates tensor reconstruction via matrix outer products.
+Compared to vector outer products in one-dimensional SPCA, this formulation more effectively incorporates structural priors into sparse weight optimization.
+Variable Merging: As Problem (16) is a non-convex problem requiring alternating update of $\hat{\mathcal{U}}$ and $\hat{\mathcal{Q}}$, a convex relaxation is beneficial for optimization.
+Consequently, we developed a convex formulation strategy through variable merging.
+As illustrated in Figure 3, tensor products in the transform domain are decomposed into multiple outer products of matrices.
+This enables pairwise multiplication of each frontal slice between the encoder and decoder, ultimately generating a novel reconstruction tensor.
+Each row vector in $\hat{\mathcal{A}}_{::i}$ integrates reconstruction weights from both $\hat{\mathcal{U}}_{::i}$ and $\hat{\mathcal{Q}}^{H}_{::i}$, establishing $\mathcal{A}$ as a superior feature selection tensor compared to $\mathcal{Q}$ alone.
+Guided by $D^{o}$, each frontal slice of $\hat{\mathcal{A}}$ reconstructs mode-specific fibers in the data tensor, thereby leveraging structural information inherent in the mode organization.
+Theorem 1 demonstrates that the optimal $\hat{\mathcal{A}}$ constitutes a Hermitian positive semidefinite (HPSD) matrix.
 Theorem 1.
@@
 pfProof of Theorem 1 {pf} See Appendix.
-Norm Simplification: Since $\hat{\mathcal{A}}_{::i}=\hat{\mathcal{U}}_{::i}\hat{\mathcal{Q}}^{H}_{::i}$ and $d\gg k$, $\hat{\mathcal{A}}_{::i}$ is also a low-rank matrix.
-It is widely known that nuclear norm $\|\cdot\|_{*}$ is often used as the convex approximation to the low-rank matrix constraint.
-However, its high computational complexity poses a challenge.
-Fortunately, in our case, $\|\hat{\mathcal{A}}_{::i}\|_{*}={\rm Tr}\left(\hat{\mathcal{A}}_{::i}\right)$ due to the HPSD constraint, significantly reducing computational complexity.
-Meanwhile, provided that $\hat{\mathcal{U}}_{::i}$ is orthogonal, it is easy to prove that
+Figure 3: A schematic diagram of merging the encoder and decoder tensors.
+Norm Simplification: Given $\hat{\mathcal{A}}_{::i}=\hat{\mathcal{U}}_{::i}\hat{\mathcal{Q}}^{H}_{::i}$ and $d\gg k$, each $\hat{\mathcal{A}}_{::i}$ is inherently low-rank.
+While the nuclear norm $\|\cdot\|_{*}$ serves as the standard convex approximation for low-rank constraints, its computational demands present significant challenges.
+Fortunately, the HPSD constraint enables the equivalence $\|\hat{\mathcal{A}}_{::i}\|_{*}={\rm Tr}\left(\hat{\mathcal{A}}_{::i}\right)$, dramatically reducing computational complexity.
+Furthermore, leveraging the orthogonality of $\hat{\mathcal{U}}_{::i}$, we establish:
 ${\|\hat{\mathcal{Q}}_{::i}\|}_{2,1}={\|\hat{\mathcal{U}}_{::i}\hat{\mathcal{Q}}^{H}_{::i}\|}_{2,1}={\|\hat{\mathcal{A}}_{::i}\|}_{2,1}={\|\hat{\mathcal{A}}^{H}_{::i}\|}_{2,1}.$ (17)
-Therefore we can use ${\|\hat{\mathcal{A}}_{::i}\|}_{2,1}$ to simplify optimization.
-Moreover, minimizing ${\|\hat{\mathcal{A}}_{::i}\|}_{2,1}$ and ${\rm Tr}\left(\hat{\mathcal{A}}_{::i}\right)$ is equivalent to minimizing ${\|\mathcal{A}_{::i}\|}_{2,1}$ and ${\rm Tr}\left(\mathcal{A}_{::i}\right)$ when $\mathbf{M}$ is orthogonal.
-This implies that the elements in a column vector of $\mathcal{A}_{::i}$ (representing different weights for the same feature) are minimized toward zero for redundant features, thereby fulfilling the feature selection task.
+This enables substitution with ${\|\hat{\mathcal{A}}_{::i}\|}_{2,1}$ to streamline optimization.
+Moreover, when $\mathbf{M}$ is orthogonal, minimizing ${\|\hat{\mathcal{A}}_{::i}\|}_{2,1}$ and ${\rm Tr}\left(\hat{\mathcal{A}}_{::i}\right)$ is equivalent to minimizing ${\|\mathcal{A}_{::i}\|}_{2,1}$ and ${\rm Tr}\left(\mathcal{A}_{::i}\right)$.
+Consequently, column vectors in $\mathcal{A}_{::i}$ (representing weights for individual features) are collectively minimized toward zero for redundant features, thereby fulfilling feature selection.
 Finally, a convex version of STPCA-MP for UFS is formulated as follows:
@@
 s.t. $\hat{\mathcal{A}}_{::i}\in H^{q}_{+},i\in[p].$
-Discussion: The convex version of STPCA-MP exhibits a compact model expression.
-By merging the sparse encoder and orthogonal decoder, the reconstruction tensor simultaneously assigns weights to original features and their latent representations, comprehensively measuring feature importance.
+The HPSD constraint, while emerging naturally from variable merging, plays a dual role: it statistically guarantees positive contributions from all projection directions to the reconstruction process.
+This property intrinsically suppresses random noise features, leading to enhanced weight interpretability and ultimately more effective feature selection.
+Discussion: The convex STPCA-MP formulation achieves model compactness through variable merging.
+By unifying the sparse encoder and orthogonal decoder, the reconstruction tensor $\mathcal{A}$ simultaneously weights original features and their latent representations, providing a comprehensive feature importance measure.
 3.3 Optimization Algorithm
-The key to solving Problem (17) lies in leveraging the tensor structure to handle the HPSD constraint with reduced computational complexity.
-Multiple Dual Subproblems: The proposed model in the original domain (Problem (17)) has a dual problem in the transform domain, since their solutions are uniquely connected via $\mathbf{M}$.
-As demonstrated below, the model is significantly easier to solve in the transform domain.
-First, we rewrite Problem (17) in the transform domain:
+Solving Problem (18) efficiently requires leveraging the tensor structure to handle HPSD constraints with reduced computational complexity.
+Multiple Dual Subproblems: Due to the bijective mapping $\hat{\mathcal{A}}=\mathcal{A}\times_{3}\mathbf{M}$, Problem (18) admits a dual formulation in the transform domain.
+This dual problem proves computationally tractable as shown below.
+We first recast it as:
 $\underset{\hat{\mathcal{A}}}{\min}$ $\sum\limits^{p}_{i=1}\left({\|\left(\hat{\mathcal{X}}_{D^{o}}\right)_{::i}-\hat{\mathcal{A}}_{::i}\left(\hat{\mathcal{X}}_{D^{o}}\right)_{::i}\|^{2}_{F}}+\lambda{\|\hat{\mathcal{A}}_{::i}\|}_{2,1}+\eta{\rm Tr}\left(\hat{\mathcal{A}}_{::i}\right)\right)$ (19)
-$s.t.$ $\hat{\mathcal{A}}_{::i}\in H^{q}_{+},i\in[p],$
-which can be decomposed into $p$ independent subproblems. We can divide and conquer each subproblem and gather the results to construct $\mathcal{A}$.
-Derivative: For the $i$-th subproblem, we calculate the derivative concerning $\hat{\mathcal{A}}_{::i}$ and set the value to 0, yielding:
-$\hat{\mathcal{A}}_{::i}=\left(\mathbf{S}_{i}-\frac{\eta_{i}}{2}\mathbf{I}_{q}\right)\left(\mathbf{S}_{i}+\lambda_{i}{\mathbf{W}_{i}}+\epsilon_{2}\mathbf{I}_{q}\right)^{-1},$ (19)
-where $\mathbf{S_{i}}=\left(\hat{\mathcal{X}}_{D^{o}}\right)_{::i}\left(\hat{\mathcal{X}}_{D^{o}}\right)^{T}_{::i}$ and $\mathbf{W}_{i}\in{\mathbb{R}^{q\times q}}$ is a diagonal matrix whose $j$-th diagonal element is calculated by:
-$\mathbf{w}_{jj}=\left(1/\left(2\sqrt{\hat{\mathbf{a}}_{j}^{H}\hat{\mathbf{a}}_{j}+\epsilon_{1}}\right)\right).$ (20)
-HPSD Projection: The next step is to project $\hat{\mathcal{A}}_{::i}$ onto the HPSD cone.
-This is achieved using the HPSD projection operator $P_{H^{d}_{+}}(\cdot)$ [55] which, for an arbitrary square matrix $\mathbf{A}\in\mathbb{C}^{b\times b}$, is defined as:
+s.t. $\hat{\mathcal{A}}_{::i}\in H^{q}_{+},\quad i\in[p],$
+which decomposes into $p$ independent subproblems. This enables a divide-and-conquer strategy: solve each frontal slice optimization separately, then reconstruct $\mathcal{A}$ post-optimization.
+Derivative Analysis: For the $i$-th subproblem, setting the gradient with respect to $\hat{\mathcal{A}}_{::i}$ to zero yields:
+$\hat{\mathcal{A}}_{::i}=\left(\mathbf{S}_{i}-\frac{\eta_{i}}{2}\mathbf{I}_{q}\right)\left(\mathbf{S}_{i}+\lambda_{i}{\mathbf{W}_{i}}+\epsilon_{1}\mathbf{I}_{q}\right)^{-1},$ (20)
+where $\mathbf{S}_{i}=\left(\hat{\mathcal{X}}_{D^{o}}\right)_{::i}\left(\hat{\mathcal{X}}_{D^{o}}\right)^{T}_{::i}$, $\epsilon_{1}>0$ is a small constant added for inversion stability, and $\mathbf{W}_{i}\in{\mathbb{R}^{q\times q}}$ denotes a diagonal matrix with $j$-th diagonal element:
+$\mathbf{w}_{jj}=\left(1/\left(2\sqrt{\hat{\mathbf{a}}_{j}^{H}\hat{\mathbf{a}}_{j}+\epsilon_{2}}\right)\right).$ (21)
+$\epsilon_{2}>0$ is a small constant to prevent the denominator from approaching zero.
+HPSD Projection: Subsequently, $\hat{\mathcal{A}}_{::i}$ is projected onto the HPSD cone using the operator $P_{H^{d}_{+}}(\cdot)$ [58].
+For an arbitrary square matrix $\mathbf{A}\in\mathbb{C}^{b\times b}$, this projection is defined through eigenvalue decomposition:
 $P_{H^{b}_{+}}(\mathbf{A})$ $=\Pi_{H^{b}_{+}}\left(\Pi_{H^{b}}(\mathbf{A})\right)$ (22)
 $=\Pi_{H^{b}_{+}}\left(\frac{1}{2}(\mathbf{A}+\mathbf{A}^{H})\right)$
-$=\sum\limits_{i=1}^{b}max\left\{\sigma_{i},0\right\}\mathbf{u}_{i}^{H}\mathbf{u}_{i},$
-where $\sum\nolimits_{i=1}^{b}\sigma_{i}\mathbf{u}_{i}^{H}\mathbf{u}_{i}$ is the eigenvalue decomposition of $\frac{1}{2}(\mathbf{A}+\mathbf{A}^{H})$.
-Based on the above techniques, we solve the $i$-th subproblem of (18) with the following two-stage algorithm:
-Step 1: Fix $\mathbf{W}_{i}$ and Update $\hat{\mathcal{A}}_{::i}$.
-When $\mathbf{W}_{i}$ is fixed, the optimal $\hat{\mathcal{A}}_{::i}$ is obtained at:
+$=\sum\limits_{i=1}^{b}\max\left\{\sigma_{i},0\right\}\mathbf{u}_{i}\mathbf{u}_{i}^{H},$
+where $\frac{1}{2}(\mathbf{A}+\mathbf{A}^{H})=\sum\nolimits_{i=1}^{b}\sigma_{i}\mathbf{u}_{i}\mathbf{u}_{i}^{H}$ is its eigenvalue decomposition.
+Building on these techniques, we solve the $i$-th subproblem of (19) using a two-stage iterative algorithm:
+Step 1: Update $\hat{\mathcal{A}}_{::i}$ with $\mathbf{W}_{i}$ fixed. For fixed $\mathbf{W}_{i}$, the optimal $\hat{\mathcal{A}}_{::i}$ is given by:
 $\hat{\mathcal{A}}_{::i}=P_{H^{q}_{+}}\left(\left(\mathbf{S}-\frac{\eta_{i}}{2}\mathbf{I}_{q}\right)\left(\mathbf{S}+\lambda_{i}{\mathbf{W}_{i}}+\epsilon_{2}\mathbf{I}_{q}\right)^{-1}\right).$ (23)
-Step 2: Fix $\hat{\mathcal{A}}_{::i}$ and Update $\mathbf{W}_{i}$.
-When $\hat{\mathcal{A}}_{::i}$ is fixed, $\mathbf{W}_{i}$ can computed directly by (20).
-Discussion: Formulating a series of convex subproblems simplifies the optimization algorithm design.
-Although the HPSD projection requires eigenvalue decomposition with a computational complexity of $O(x^{3})$, the exploitation of tensor structures ensures that $x\ll\#{\rm Feature}$.
-The efficiency of the proposed method is demonstrated in the experimental results.
-3.4 Feature selection
-We assess the importance of an original feature by its contribution to the reconstruction.
-Given the presence of the transform domain, the influence of the transform matrix $\mathbf{M}$ must be considered.
+Step 2: Update $\mathbf{W}_{i}$ with $\hat{\mathcal{A}}_{::i}$ fixed. For fixed $\hat{\mathcal{A}}_{::i}$, $\mathbf{W}_{i}$ is directly computed via (21).
+Discussion: This convex subproblem formulation significantly simplifies optimization design.
+While HPSD projection requires eigenvalue decomposition with $O(x^{3})$ complexity, tensor structure exploitation ensures $x\ll|\mathcal{F}|$ where $\mathcal{F}$ denotes the feature set.
+Experimental results demonstrate the method’s computational efficiency.
+3.4 Feature selection via score map
+We quantify feature importance by its contribution to tensor reconstruction, accounting for the transform domain induced by $\mathbf{M}$.
 Suppose we reorient the data tensor to obtain $\mathcal{X}_{D^{o}}\in\mathbb{C}^{p\times n\times q}$.
@@
 $\tilde{f}_{ij}$ $=\sum\limits^{p}_{l=1}\hat{a}_{ilj}\hat{f}_{lj}=\sum\limits^{p}_{l=1}\left(\sum\limits^{p}_{k=1}m_{ik}a_{ilk}\right)\left(\sum\limits^{q}_{h=1}m_{jh}f_{lh}\right)=\sum\limits^{p}_{l=1}\sum\limits^{q}_{h=1}\left(\sum\limits^{p}_{k=1}m_{jh}m_{ik}a_{ilj}\right)f_{lh},$
-which indicates that the contribution of $f_{lh}$ to the reconstruction of $f_{ij}$ can be described by $\sum^{p}_{k=1}m_{jh}m_{ik}a_{ilj}$.
-Therefore, the contribution matrix of $f_{lh}$ is defined as:
-$\left[Contribution\right]_{ij}=\sum\limits^{p}_{l=1}\sum\limits^{q}_{h=1}\left(\sum\limits^{p}_{k=1}m_{jh}m_{ik}a_{ilj}\right).$ (24)
-Finally, the score for $f_{lh}$ is computed as $\|Contribution\|^{2}_{F}$, which comprehensively captures its influence on the reconstruction of all features and leverages the global information.
-We summarize the whole algorithm in Algorithm 1.
-Discussion: Equation (24) shows that an appropriate choice of $\mathbf{M}$ integrates feature correlation information from other modes into the mode targeted for slice-by-slice optimization.
-In our proposed method, each feature is scored by globally modeling its dependencies with other features without requiring sample vectorization, thus preserving the data organization and reducing computational complexity.
+The contribution of feature $f_{lh}$ to the reconstruction of all features is quantified by the matrix $\mathbf{C}^{lh}\in\mathbb{C}^{p\times q}$:
+$\mathbf{C}^{lh}_{ij}=\sum_{k=1}^{p}m_{jh}m_{ik}a_{ilj}.$ (25)
+The feature importance score is then defined as:
+$score(f_{lh})=\|\mathbf{C}^{lh}\|_{F}^{2}=\sum_{i=1}^{p}\sum_{j=1}^{q}\left(\sum_{k=1}^{p}m_{jh}m_{ik}a_{ilj}\right)^{2},$ (26)
+which comprehensively captures $f_{lh}$’s global influence on feature reconstruction. The complete procedure is summarized in Algorithm 1.
+Discussion: Equation (25) demonstrates that an appropriately chosen $\mathbf{M}$ effectively integrates cross-mode feature correlations into the slice-by-slice optimization process.
+Our scoring mechanism globally quantifies feature dependencies through tensor operations, eliminating the need for sample vectorization.
+This approach provides dual advantages:
+  • Preservation of intrinsic data organization structures
+  • Significant reduction in computational complexity
+by maintaining the native tensor representation throughout the feature evaluation process.
+3.5 Implementation details
+The proposed optimization framework decomposes into multiple independent subproblems amenable to parallel computation. We establish a unified convergence criterion for all subproblems:
+$|\text{obj}^{(h)}-\text{obj}^{(h-1)}|<10^{-5},\quad h=1,2,3,\cdots$ (27)
+where $\text{obj}^{(h)}$ denotes the objective function value at iteration $h$.
+For regularization parameters $\lambda$ and $\eta$, we optimized them via grid search per dataset in the following comparison experiments.
+In practice, $\lambda$ and $\eta$ can be set to $10$ for a rapid deployment.
+The stabilization constants are fixed at $\epsilon_{1}=10^{-3}$ and $\epsilon_{2}=10^{-2}$ throughout all experiments.
+For data preprocessing, we apply centralization to $\left(\mathcal{X}_{D^{o}}\right)_{::i}$ for each subproblem individually to capture inter-dimension variance more effectively.
+To narrow down the grid search range, we normalize $\mathcal{X}$ by dividing each element by the maximum absolute value, scaling all values to the interval [-1, 1].
 Algorithm 1 The optimization algorithm for STPCA-MP
-0:  A third-order centralized data tensor $\mathcal{X}\in\mathbb{C}^{d_{1}\times d_{2}\times n}$, the order set $D^{o}$ determining $q$, an invertible matrices $\mathbf{M}\in\mathbb{C}^{q\times q}$, regularization parameters $\lambda,\eta$.
+0:  A third-order centralized data tensor $\mathcal{X}\in\mathbb{C}^{d_{1}\times d_{2}\times n}$, the order set $D^{o}$ determining $q$, an invertible matrices $\mathbf{M}\in\mathbb{C}^{q\times q}$, regularization parameters $\lambda,\eta$, and $\epsilon_{1}=10^{-3}$, $\epsilon_{2}=10^{-2}$.
 0:  $h$ selected features.
-1:  Randomly initialize $\left(\tilde{\mathcal{A}}\right)_{::i}\in{H^{q}_{+}}$, transform $\mathcal{X}$ into $\tilde{\mathcal{X}}$.
+1:  Randomly initialize $\left(\hat{\mathcal{A}}\right)_{::i}\in{H^{q}_{+}}$, transform $\mathcal{X}$ into $\hat{\mathcal{X}}$.
 2:  for $i=1,2,\cdots,q$ do
-3:     revolve $\tilde{\mathcal{X}}$ with $D^{o}$ and obtain $\tilde{\mathcal{X}}_{D^{o}}$.
-4:     input $\left(\tilde{\mathcal{X}}_{D^{o}}\right)_{::i}$ and alternatively update $\hat{\mathcal{A}}_{::i}$ and $\mathbf{W}_{i}$ according to (22) until converge.
+3:     revolve $\hat{\mathcal{X}}$ with $D^{o}$ and obtain $\hat{\mathcal{X}}_{D^{o}}$.
+4:     input centralized $\left(\hat{\mathcal{X}}_{D^{o}}\right)_{::i}$ and alternatively update $\hat{\mathcal{A}}_{::i}$ and $\mathbf{W}_{i}$ according to (23) until $|\text{obj}^{(h)}-\text{obj}^{(h-1)}|<10^{-5}$.
 5:  end for
-6:  Compute $\mathcal{A}=\hat{\mathcal{A}}\times_{3}\mathbf{M}^{-1}$ and $\|Contribution\|^{2}_{F}$ for each score according to (24).
+6:  Compute $\mathcal{A}=\hat{\mathcal{A}}\times_{3}\mathbf{M}^{-1}$ and $\|\mathbf{C}^{lh}\|_{F}^{2}$ for each score according to (25).
 7:  Output a feature score map $score$.
-8:  Sort $Score$ in descending order, and select features of the $h$ greatest scores. For slice-wise tensor, the score for each dimension will be obtained by summation.
+8:  Sort $score$ in descending order, and select features of the $h$ greatest scores. For a slice-wise tensor, the score for each dimension will be obtained by summation.
 4 Discussion
@@
 4.1 Convergence analysis
-First, we prove the convergence of the proposed algorithm. Let us introduce a lemma [40]:
+First, we prove the convergence of the proposed algorithm. The proof leverages the following fundamental lemma [43]:
 Lemma 1.
@@
 When the solution in the transform domain is obtained, the dual problem in the original domain is solved as well.
-4.2 Computational complexity
-Suppose $\mathcal{X}\in\mathbb{C}^{d_{1}\times d_{2}\times n}$ is rotated into $\mathcal{X}_{D^{o}}\in\mathbb{C}^{q\times n\times p}$ ($n$ is the number of samples, $(p,q)\in\{(d_{1},d_{2}),(d_{2},d_{1})\}$).
-For each subproblem, it takes $O\left(q^{2}\right)$ to convert data into the transform domain and initialize the data covariance matrix, and $O\left(q^{3}\right)$ to update the reconstruction matrix in each iteration (including matrix inversion and HPSD projection).
-Therefore we have $p$ subproblems each solved at the computational cost of $O\left(q^{3}t_{i}+q^{2}n\right)$ ($t_{i}$ is the number of iterations for the $i$-th subproblem).
-The final feature score computation takes $O(p^{2}q)$.
-Eventually, the total computational complexity is $O\left(\sum\limits^{p}_{i=1}\left(q^{2}n+q^{3}t_{i}\right)+p^{2}q\right)$.
-The number of features ($q\times p$) is usually much greater than the number of dimensions in a single mode ($q$ or $p$).
-Due to the exploitation of tensor structures with separate operations on each slice, STPCA-MP can handle data tensors at a remarkably smaller computation cost than non-tensor-based methods.
-For example, for image data of size $32\times 32$, non-tensor-based methods requiring matrix inversion or eigenvalue decomposition (such as SOGFS [40] and SPCAFS [38]; see Table 1) have a computational complexity of $O(1024^{3})$, whereas STPCA-MP requires only $O(32\times(32^{3}))$, reducing the complexity by $1024$ times.
-In addition, the computational complexity is linear to the number of samples, which means the proposed method will still be efficient when facing data tensors with numerous samples.
-Table 1 shows the computational complexity of competitive methods in the following experiments.
-We can see that tensor-based methods reduce the computational burden from $\prod_{i}d_{i}^{x}$ to $d_{i}^{x}$ in each iteration by utilizing the tensor structure.
-Among tensor-based methods, CPUFS and CPUFSnn are the most recently proposed, which have a computational complexity linear to $d_{i}$ at the price of squared to $n$.
-They are non-convex models and usually take hundreds of iterations to converge, making it a challenge when facing large-scale datasets.
+4.2 Computational complexity analysis
+Consider a data tensor $\mathcal{X}\in\mathbb{C}^{d_{1}\times d_{2}\times n}$ reoriented as $\mathcal{X}_{D^{o}}\in\mathbb{C}^{q\times n\times p}$, where $n$ denotes the number of samples and $(p,q)\in\{(d_{1},d_{2}),(d_{2},d_{1})\}$.
+The computational cost comprises three main components:
+  • Transform domain conversion: $O(q^{2})$ operations for data transformation and covariance matrix initialization per frontal slice.
+  • Iterative optimization: $O(q^{3})$ per iteration for reconstruction matrix updates (including matrix inversion and HPSD projection).
+  • Feature scoring: $O(p^{2}q)$ for final importance evaluation.
+For $p$ independent subproblems, each requiring $t_{i}$ iterations, the total complexity is:
+$O\left(\sum_{i=1}^{p}(q^{2}n+q^{3}t_{i})+p^{2}q\right)$ (38)
+Computational Efficiency Advantages: Table 4 summarizes the computational complexity of comparable methods in subsequent experiments.
+The proposed STPCA-MP framework achieves significant computational efficiency through three core mechanisms: First, dimensionality reduction via tensor decomposition leverages the inherent structure by exploiting the inequality $q\times p\gg\max(q,p)$ through slice-by-slice operations, systematically replacing high-dimensional matrix computations with mode-decomposed equivalents and reducing per-iteration complexity from $\prod_{i}d_{i}^{x}$ to $d_{i}^{x}$.
+For $32\times 32$ image data, this reduces complexity from $O(1024^{3})$ (as in vectorized methods SOGFS [43] and SPCAFS [41]) to $O(32\times(32^{3}))$, representing a 1024-fold reduction.
+Second, linear sample scalability ensures complexity grows as $O(n)$, maintaining large-scale dataset efficiency, in contrast to the quadratic sample complexity $O(n^{2})$ incurred by graph-updating methods like CPUFS [21].
+Third, the convex formulation with HPSD projection enables stable, rapid convergence, reducing iteration counts and further conserving runtime.
+Crucially, STPCA-MP uniquely maintains all advantages simultaneously, preserving the original tensor structure while achieving linear scaling in both feature dimensions ($d_{i}$) and sample size ($n$)—a dual efficiency unattained by approaches like CPUFS and CPUFSnn that sacrifice sample efficiency for dimensionality reduction.
 Table 4: Computational complexity of competitive methods on third-order tensors.
@@
 Method Input Computational complexity
-SOGFS [40] matrix $O\left(dn^{2}+t(d^{3}t_{1}+n^{2}c)\right)$
-RNE [50] matrix $O\left(dn^{2}+t(t_{1}(n+s)d^{2}+t_{2}n))\right)$
-SPCAFS [39] matrix $O\left(d^{2}n+d^{3}t\right)$
-NLGMS [56] matrix $O\left(t\left(d^{3}+n^{3}+nd+n^{2}d+d\log{d}\right)\right)$
-CPUFS [18] tensor $O\left(d_{1}nc^{2}+d_{2}nc^{2}+d_{1}d_{2}nc+n^{2}c)t\right)$
-CPUFSnn [18] tensor $O\left((d_{1}nc^{2}+d_{2}nc^{2}+d_{1}d_{2}nc+(n_{1}c)^{2}+(n_{2}c)^{2}+n^{2}c)t\right)$
-MSPCA [33] tensor $O\left(t_{2}\sum_{i=1}^{2}(t_{1}+n)d_{i}^{3}\right)$
-STPCA-MP tensor $O\left(\sum^{p}_{i=1}\left(q^{2}n+q^{3}t_{i}\right)+p^{2}q\right)(q\in\{d_{1},d_{2}\})$
+SOGFS [^1] Structured Optimal Graph for Unsupervised Feature Selection [43] matrix $O\left(dn^{2}+t(d^{3}t_{1}+n^{2}c)\right)$
+RNE [^2] Robust Neighborhood Embedding for unsupervised feature selection [53] matrix $O\left(dn^{2}+t(t_{1}(n+s)d^{2}+t_{2}n))\right)$
+SPCAFS [^3] Sparse PCA via $\ell_{2,p}$-Norm regularization for unsupervised Feature Selection [42] matrix $O\left(d^{2}n+d^{3}t\right)$
+NLGMS [^4] Unsupervised group feature selection approach via non-convex regularized graph embedding and self-representation [59] matrix $O\left(t\left(d^{3}+n^{3}+nd+n^{2}d+d\log{d}\right)\right)$
+CPUFS [^5] CP decomposition based Unsupervised Feature Selection [21] tensor $O\left(d_{1}nc^{2}+d_{2}nc^{2}+d_{1}d_{2}nc+n^{2}c)t\right)$
+CPUFSnn [^6] CP decomposition based Unsupervised Feature Selection with non-negativity [21] tensor $O\left((d_{1}nc^{2}+d_{2}nc^{2}+d_{1}d_{2}nc+(n_{1}c)^{2}+(n_{2}c)^{2}+n^{2}c)t\right)$
+MSPCA [^7] Multilinear Sparse Principal Component Analysis [36] tensor $O\left(t_{2}\sum_{i=1}^{2}(t_{1}+n)d_{i}^{3}\right)$
+STPCA-MP (ours) tensor $O\left(\sum^{p}_{i=1}\left(q^{2}n+q^{3}t_{i}\right)+p^{2}q\right)(q\in\{d_{1},d_{2}\})$
 4.3 Extension to higher-order tensor data
@@
 5 Experiments
-In this section, we conduct experiments on real-world data to demonstrate our proposed method’s effectiveness.
-The codes related to our proposed methods and the experiments are available at https://github.com/zjj20212035/STPCA.git.
+In this section, we conduct experiments on real-world datasets to empirically validate the effectiveness of our proposed method.
+Implementation codes for STPCA-MP and all experiments are publicly hosted at: https://github.com/zjj20212035/STPCA.git.
 5.1 Experimental Settings
 5.1.1 Experimental environment
-All the experiments are conducted on a personal computer with a 2.20 GHz CPU and 64GB main memory under the environment of the Windows 11 operation system.
+All experiments are executed on a standard workstation equipped with a 2.20 GHz CPU and 64 GB RAM, utilizing the Windows 11 operating system environment.
 5.1.2 Comparative methods
-We compare our proposed methods with some state-of-the-art UFS methods as follows:
+We evaluate our method against representative state-of-the-art UFS baselines:
   1. [•]
   2. 1.
-    CPUFS and CPUFSnn [18] (2023): CPUFS is a tensor-based UFS method combining CP decomposition and spectral clustering.
-It utilizes a sparse classifier to perform UFS.
-CPUFSnn is a variation with nonnegative constraints on the classifier matrix.
+    CPUFS and CPUFSnn [21] (2023): CPUFS is a tensor-based unsupervised feature selection method that integrates CP decomposition with spectral clustering.
+The method employs a sparse classifier to perform feature selection.
+CPUFSnn extends this framework by imposing non-negative constraints on the classifier matrix, enhancing interpretability through element-wise positivity requirements.
   3. 2.
-    MSPCA [33] (2014): MSPCA is a sparse tensor PCA method that utilizes Tucker decomposition to make use of tensor structure.
-It was originally proposed as a feature extraction method.
-We use the Kronecker product of its factor matrices to perform UFS, as an example of Tucker-based UFS methods.
+    MSPCA [36] (2014): MSPCA represents a sparse tensor PCA approach employing Tucker decomposition to leverage inherent tensor structures.
+Originally developed as a feature extraction technique, we adapt it for unsupervised feature selection (UFS) by utilizing the Kronecker product of its factor matrices, thereby demonstrating a Tucker-based UFS implementation.
   4. 3.
-    NLGMS [56] (2024): NLGMS is a non-tensor-based method that introduces a self-representation technique into the graph-based framework to preserve global and local data structure.
-The self-representation term reconstructs each sample with the linear combination of other samples.
+    NLGMS [59] (2024): NLGMS constitutes a non-tensor-based approach that incorporates a self-representation mechanism within a graph-based framework to simultaneously maintain both global and local data structures.
+The method’s self-representation component reconstructs individual samples through linear combinations of other data instances.
   5. 4.
-    SPCAFS [38] (2021): SPCAFS is a non-tensor-based method that introduces $\ell_{2,p}$-norm into the one-dimensional PCA model. It forms a simple yet effective model.
+    SPCAFS [41] (2021): SPCAFS is a non-tensor-based approach that incorporates the $\ell_{2,p}$-norm regularization into the standard one-dimensional PCA framework.
+This formulation yields an efficient and computationally tractable model.
   6. 5.
-    RNE [50] (2020): RNE is a non-tensor-based method that makes use of the locally linear embedding algorithm to obtain an optimal projection matrix.
-It keeps the samples close to their neighbors after being projected onto a local manifold.
+    RNE [53] (2020): RNE represents a non-tensor-based approach that employs the locally linear embedding (LLE) algorithm to derive an optimal projection matrix.
+The method maintains neighborhood relationships by preserving proximity between samples and their local neighbors when projected onto the manifold space.
   7. 6.
-    SOGFS [40] (2019): SOGFS is a non-tensor-based method that optimizes an adaptive graph to preserve the local geometrical structure of data.
-It updates the similarity matrix during each iteration to obtain a more flexible graph.
-We also perform clustering with all features as a baseline method. All the comparative methods are implemented using the codes provided by their authors.
+    SOGFS [43] (2019): SOGFS is a non-tensor-based approach that optimizes an adaptive graph structure to maintain the local geometric relationships within the data.
+The method iteratively refines the similarity matrix to achieve progressively more flexible graph representations.
+For baseline comparison, we additionally conduct clustering using the complete set of original features. All competing methods are implemented through their respective authors’ original code releases.
 5.1.3 Parameter settings
-For all methods, we tune the regularization parameters via grid search over $\{10^{-2},10^{-1},1,10,10^{2}\}$.
-For graph-based methods, we set the number of k-nearest neighbors to 5.
-For methods utilizing $\ell_{2,p}$-norm, we set $p=1$.
-For STPCA-MP, we set $\mathbf{M}=\mathbf{I}$ in most experiments except the one that discusses the choice of $\mathbf{M}$.
-The direction set $D^{o}$ is also tuned by a grid-search strategy in the range $\{\{1,3,2\},\{2,3,1\}\}$.
-The other parameters remain the default in the open code provided by the authors.
-The best results of all methods are presented.
+All methods employ consistent experimental protocols: (1) Regularization parameters are determined through grid search over $\{10^{-2},10^{-1},1,10,10^{2}\}$; (2) Uniform feature dimensionality is maintained across comparative methods.
+Specific configurations include: (i) Graph-based methods use $k=5$ nearest neighbors; (ii) $\ell_{2,p}$-norm methods fix $p=1$; (iii) STPCA-MP initializes $\mathbf{M}=\mathbf{I}$ for standard experiments (excluding the dedicated $\mathbf{M}$ analysis); (iv) Direction set $D^{o}$ undergoes grid search in $\{{1,3,2\},\{2,3,1\}}$.
+Remaining parameters follow default settings from authors’ original implementations.
+All reported results reflect optimal parameter configurations.
 5.1.4 Evaluation methodology and metrics
-In clustering experiments, after a UFS method has finished selecting features, we evaluate its performance by first conducting K-means clustering and then mapping the obtained pseudo labels to the real labels by adopting the Kuhn-Munkres algorithm [41].
-Due to the K-means’ dependence on initialization, we repeat the clustering 30 times.
-The clustering metrics we use are as follows:
+Following feature selection by each UFS method, we assess clustering performance through a standardized evaluation pipeline:
+  1. [1)]
+  2. 1.
+    Cluster Generation: Apply $K$-means clustering to the selected features.
+  3. 2.
+    Label Alignment: Establish correspondence between pseudo-labels and ground truth using the Kuhn-Munkres algorithm [44].
+  4. 3.
+    Stability Control: Repeat the clustering process 30 independent times and average the result to account for $K$-means’ initialization sensitivity.
+The following clustering metrics are employed for quantitative evaluation:
   1. [•]
   2. 1.
-    Accuracy (ACC): ACC refers to the clustering accuracy and is computed according to (35) and (36), where $n$ is the total number of samples, $map(i)$ denotes the clustering label of the $i$-th sample after mapping to the real label, and $label(i)$ represents the ground truth label.
-A larger ACC indicates better performance.
-    $\text{ACC}=\frac{1}{n}\sum\limits_{i=1}^{n}\delta\left(map(i),label(i)\right).$ (35)
-    $\delta(p,q)=\left\{\begin{aligned} &1\quad\text{if}\;p=q,\\ &0\quad\text{otherwise}.\end{aligned}\right.$ (36)
+    Accuracy (ACC): The clustering accuracy (ACC) measures the agreement between cluster assignments and ground truth labels, computed as:
+    $\mathrm{ACC}=\frac{1}{n}\sum_{i=1}^{n}\delta(\mathrm{map}(i),\mathrm{label}(i))$ (39)
+    where $\delta(p,q)$ is the indicator function defined in (40), $n$ denotes the total sample size, $\mathrm{map}(i)$ represents the mapped cluster label of the $i$-th sample, and $\mathrm{label}(i)$ indicates the ground truth label.
+Higher ACC values correspond to better clustering performance.
+The indicator function is given by:
+    $\delta(p,q)=\begin{cases}1&\text{if }p=q\\ 0&\text{otherwise}\end{cases}$ (40)
   3. 2.
-    Normalized Mutual Information (NMI) [42]: NMI is utilized to describe the mutual dependence between clustering results (after being mapped into the real labels) and ground truth labels.
+    Normalized Mutual Information (NMI) [45]: NMI is utilized to describe the mutual dependence between mapped clustering labels and ground truth labels.
 Given the clustering results $\mathbf{m}$ and the ground truth labels $\mathbf{l}$ (both including all samples), $I(\mathbf{m},\mathbf{l})$ denotes the mutual information between $\mathbf{m}$ and $\mathbf{l}$.
 $H(\mathbf{m})$ and $H(\mathbf{l})$ are the entropy of $\mathbf{m}$ and $\mathbf{l}$ respectively.
-The same as ACC, a larger NMI means better clustering performance.
+A larger NMI means better clustering performance.
     $\text{NMI}(\mathbf{m},\mathbf{l})=\frac{I(\mathbf{m},\mathbf{l})}{\sqrt{H(\mathbf{m})H(\mathbf{l})}}.$ (41)
   4. 3.
+    Silhouette score (SS) [16]: SS is a widely-used metric for evaluating clustering quality that measures both intra-cluster compactness and inter-cluster separation.
+For a given sample $i$, the silhouette score $s(i)$ is computed as:
+    $s(i)=\frac{b(i)-a(i)}{\max\{a(i),b(i)\}},$ (42)
+    where $a(i)$ is the average distance between sample $i$ and all other points in the same cluster (intra-cluster distance), and $b(i)$ is the smallest average distance between $i$ and points in any other cluster (neighboring cluster distance).
+The global silhouette score SS for the entire dataset ranges from $-1$ to $1$, with values closer to $1$ indicating better-defined clusters:
+    $\text{SS}=\frac{1}{N}\sum_{i=1}^{N}s(i).$ (43)
+  5. 4.
     Training time: For an algorithm, training time starts from the initialization and ends when the algorithm outputs the selected features.
-For orientation-specific data such as multi-channel time series, the temporal dependency of the elements within the same channel makes the clustering metrics unreliable due to the feature independence hypothesis in K-means.
-Therefore, we define a metric as follows:
+For orientation-specific data (e.g., multi-channel time series), the temporal dependencies within channels violate the feature independence assumption underlying $K$-means clustering.
+This fundamental mismatch renders standard clustering metrics unreliable.
+We therefore propose the Proportion of Correctly selected features (POC) as an alternative evaluation metric:
   1. [•]
@@
     Proportion of correctly selected features (POC): We compute POC as:
-    $\text{POC}=c/h,$ (38)
-    where $c$ is the number of correctly selected features, and $h$ is the total number of selected features at each execution.
-POC reflects a UFS method’s ability to select discriminative features.
-In the experiments, we select features with maximal between-class variance (BCV) as the correct ones.
-We compute POC on slice-wise data instead of ACC and NMI.
+    $\text{POC}=\frac{c}{h},$ (44)
+    where $c$ denotes the number of features selected by both the unsupervised feature selection (UFS) method and the maximal between-class variance (BCV) criterion, and $h$ represents the total features selected per execution.
+Suppose data has been centralized, the BCV for the $(i,j)$-th element in a frontal slice is defined as:
+    $\text{BCV}(i,j)=\frac{1}{K}\sum_{k=1}^{K}\left(\mu^{(k)}_{ij}-\mu_{ij}\right)^{2},\quad i=1,\dots,d_{1};\ j=1,\dots,d_{2}$ (45)
+    where $\mu_{ij}$ is the sample mean of the $(i,j)$-th element across all samples, $\mu^{(k)}_{ij}$ is the sample mean within class $k$, and $K$ is the number of classes.
+Here, $d_{1}$ and $d_{2}$ denote the dimensionalities of the first and second modes, respectively.
+For slice-wise data where each channel is treated as a feature (See Table 5 and Section 5.2), the BCV for the $f$-th channel is obtained by aggregating along the temporal dimension:
+    $\text{BCV}_{f}=\sum_{j=1}^{d_{2}}\text{BCV}(i,j)\quad\text{for fixed channel index }i=f.$ (46)
+    This formulation quantifies the cumulative class-discriminative power of the entire time series in channel $f$.
+The POC metric thereby directly evaluates a method’s capability to identify discriminative features where conventional clustering metrics fail.
 5.2 Datasets
-Before detailing specific datasets, we first categorize tensor datasets based on their structural organization.
-As shown in Figure 2, third-order tensor datasets can be divided into two scenarios: tube-wise and slice-wise, according to their organization.
-Slice-wise data is orientation-specific and exhibits feature relations along the 1-mode, while tube-wise data exhibits feature relations along both the 1-mode and 2-mode..
-This distinction provides a basis for evaluating the ability of UFS methods to utilize tensor structure information.
-In our evaluation, we compare clustering ACC and NMI on tube-wise data and POC on slice-wise data.
-Experimental results demonstrate the effectiveness of the proposed method in both scenarios.
+Before delving into specific datasets, we first categorize tensor datasets based on their structural organization.
+As illustrated in Table 5, third-order tensor datasets can be classified into two scenarios: tube-wise and slice-wise, according to their organizational patterns.
+Slice-wise data is orientation-specific, exhibiting feature correlations along the 1-mode, whereas tube-wise data demonstrates feature correlations along both the 1-mode and 2-mode.
+This distinction serves as a foundation for evaluating the capability of UFS methods to leverage tensor structural information.
+In our evaluation, we compare ACC, NMI, and SS for tube-wise data, and use the POC for slice-wise data.
+Experimental results validate the effectiveness of the proposed method in both scenarios.
 Table 5: Two scenarios for third-order data tensors.
-Assume each frontal slice is a sample.
-Different colors refer to different variables.
-This paper regards variables as features.
-Tube-wise scenario: elements within one fiber belong to one variable.
-Slice-wise scenario: elements within one horizontal slice belong to one variable.
+Each frontal slice corresponds to a distinct sample.
+Variables (represented by distinct colors) are treated as features.
+This study adopts two feature interpretation scenarios.
+Tube-wise: Elements within a single fiber constitute one feature variable.
+Slice-wise: Elements within a horizontal slice comprise one feature variable.
 Scenario Visualization Example
@@
   2. 1.
-    Orbit: As an example of slice-wise data, Orbit comprises 3-dimensional orbit samples with two distinct radii.
+    Orbit (Figure 4(a)): As an example of slice-wise data, Orbit comprises 3-dimensional orbit samples with two distinct radii.
 Each sample includes three channels representing orbit coordinates, with the remaining channels containing random Gaussian noise.
   3. 2.
-    Array: As an example of tube-wise data, Array is constructed using a $10\times 10$ synthetic radar unit array that receives a series of echoes from a stationary target.
+    Array (Figure 4(f)): As an example of tube-wise data, Array is constructed using a $10\times 10$ synthetic radar unit array that receives a series of echoes from a stationary target.
 Each echo is represented as a matrix of unit sampling values, treated as a sample.
 Random measurement errors are introduced to half of the echo matrices to create two distinct classes.
-We evaluate our method on nine real-world datasets, including: 1) six tube-wise datasets: COIL20 [54], USPS [52], BreastMNIST [44], Imm40 [51], PIE [13] and JAFFE [43]; 2) four slice-wise datasets: UCIHAR [45] and UCIDSA [46], UMCS [^1] https://www.kaggle.com/c/seizure-detection/data and ASCH [53].
+We evaluate our method on nine real-world datasets, including: 1) six tube-wise datasets: COIL20 [57], USPS [55], BreastMNIST [47], Imm40 [54], PIE [13] and JAFFE [46]; 2) four slice-wise datasets: UCIHAR [48] and UCIDSA [49], UMCS [^8] https://www.kaggle.com/c/seizure-detection/data, and ASCH [56].
 The detailed information of each dataset is shown in Table 6.
 PIE, Imm40, and JAFFE contain frontal facial images, labeled according to individuals.
+COIL20 comprises images of various objects captured from multiple viewing angles.
 BreastMNIST is a medical dataset containing breast ultrasound images, from which we construct the data tensor using the test set.
@@
 Due to the large volume of chosen slice-wise data, we retrieve a subset of samples to construct datasets.
-For UCIDSA, we extract the first $125$ sample points from all $45$ channels covering $45$ activities.
+For UCIDSA, we extract the first $125$ sample points from all $45$ channels covering $19$ activities.
 For UCIHAR, we select $9$ individuals with all $561$ features sampled during $6$ activities.
@@
 All datasets are normalized and bounded within $[-1,1]$ for consistency.
+All the above data preprocessing is consistent across all comparative methods, including ours.
 Table 6: Statistics of datasets.
@@
 UCIHAR [48] $561\times 9$ 273 6
-UMCS 1 $65\times 125$ 174 2
+UMCS 8 $65\times 125$ 174 2
 ASCH [56] $16\times 640$ 84 2
 5.3 Synthetic experiment
-In this experiment, we conduct four typical UFS methods on synthetic datasets with the grid search strategy.
-The dataset formulation and the best experimental results of comparative methods are illustrated in Figure 3.
-On Orbit, we retrieve three top-ranked features in each method’s result and plot their samples in a subset of data.
-If a method selects the correct features (i.e.
-the three orbit coordinates), then the scatter plot will contain rings of two different radii.
-Otherwise, it will be difficult to distinguish the two classes.
-On Array, we retrieve the feature score map corresponding to the best POC for each method, analyzing the match degree between the top-ranked feature positions and the units with measurement error.
-We conclude the observations as follows.
+In this experiment, we apply four typical UFS methods to synthetic datasets using the grid search strategy.
+The dataset construction and the optimal experimental results of the comparative methods are presented in Figure 4.
+For the Orbit dataset (Figure 4(a)), we retrieve the three top-ranked features from the results of each method and plot their corresponding samples within a data subset.
+If a method successfully selects the correct features (i.e., the three orbit coordinates), the scatter plot will display rings with two distinct radii.
+Conversely, when the wrong features are chosen, distinguishing between the two classes becomes challenging.
+Regarding the Array dataset (Figure 4(f)), we retrieve the feature score map associated with the best POC for each method.
+Then, we analyze the degree of match between the positions of the top-ranked features and the units with measurement errors.
+Considering the experimental results on both synthetic datasets, the key observations are summarized as follows:
   1. [•]
   2. 1.
-    Tensor-based methods are generally more capable of analyzing the data structure than non-tensor-based methods.
-Both CPUFS and STPCA-MP select the orbit coordinates on Orbit.
-Although CPUFS fails to precisely locate the units with error on Array, its score map clearly shows a grid-like characteristic that connects the rows and columns to analyze feature relations, capturing a roughly correct attention region.
-A similar pattern also appears in STPCA-MP, where the column-wise relations are learned.
-In contrast, RNE and NLGMS isolate the features, losing the tensor structure information.
-Especially in RNE, the scores on Array seem to be randomly distributed.
+    Tensor-based methods generally outperform non-tensor-based methods in analyzing data structure.
+Both CPUFS and STPCA-MP correctly identify the orbit coordinates on the Orbit dataset (Figure 4(d)(e)).
+Although CPUFS fails to precisely pinpoint the error units on the Array dataset, its score map exhibits a distinct grid-like pattern (Figure 4(i)), which captures the row-column feature relationships and identifies an approximately correct region of interest.
+Similarly, STPCA-MP learns the column-wise relationships (Figure 4(j)).
+In contrast, RNE and NLGMS treat features in isolation, thus losing the valuable tensor structure information.
+Notably, for RNE in Figure 4(g), the scores are seemingly randomly distributed.
   3. 2.
-    Our proposed method performs the best in selecting discriminative features on the two synthetic datasets.
-STPCA-MP selects the correct features on both datasets.
-In particular, the score map on Array is relatively sparse in rows and columns.
-By taking advantage of data orientation, STPCA-MP can more easily mine the latent feature relations and locate discriminative features.
+    Our proposed method excels in selecting discriminative features on both synthetic datasets.
+STPCA-MP selects the correct features on both datasets (Figure 4(e)(j)).
+Specifically, the score map for the Array dataset (Figure 4(j)) shows relatively sparse rows and columns.
+By leveraging the data’s orientation, STPCA-MP can more effectively uncover latent feature relationships and identify discriminative features.
 Figure 4: Synthetic experiments on Orbit and Array.
@@
 The best results and the corresponding training time of all methods are given in Table 7.
-The detailed ACC and NMI curves are illustrated in Figure 4.
-By analyzing the results, we come out with the following conclusions:
+The detailed ACC, NMI, and SS curves are illustrated in Figure 5.
+From these results, we draw three main conclusions:
   1. [•]
   2. 1.
-    Features selection improves data compactness.
-On each dataset, one or more methods are capable of achieving performance that approaches or even surpasses the baseline while utilizing only a subset of features.
-In particular, the proposed STPCA-MP achieves significantly higher clustering metrics than all features on Jaffe, BreastMNIST, and Imm40.
+    Features selection enhances data compactness.
+Multiple methods achieve performance comparable to or superior to the baseline while utilizing only a subset of features.
+Notably, STPCA-MP significantly outperforms All Features on PIE (Figure 5(a)(g)(m)), Jaffe (Figure 5(c)(i)(o)), BreastMNIST (Figure 5(d)(j)(p)), and Imm40 (Figure 5(e)(k)(q)).
   3. 2.
-    STPCA-MP delivers reliable and outstanding performance in clustering tasks.
-Firstly, compared to graph-based methods like SOGFS, STPCA-MP obtains more global data information by exploiting the reconstruction term, which guarantees the basic clustering performance.
-We can see the curves of STPCA-MP, on the whole, have a smaller oscillation in amplitude and share better consistency in trend between ACC and NMI.
-Besides, STPCA-MP reflects a more natural decreasing (or increasing) progression in variation pattern.
-Secondly, the T-SVDM-based sparse framework is beneficial for extracting tensor structure information, enabling STPCA-MP to describe the feature correlation in a more delicate manner, compared to CPD-based tensor methods (CPUFS and CPUFSnn), Tucker-based tensor method (MSPCA), and non-tensor-based methods (SPCAFS, RNE).
-Therefore, STPCA-MP demonstrates superior performance, consistently ranking at the forefront on all datasets.
+    STPCA-MP demonstrates reliable and superior clustering performance.
+The proposed method captures more global data information through its reconstruction term, outperforming graph-based approaches like SOGFS, CPUFS, and CPUFSnn.
+This is evidenced by: (i) smoother performance curves with smaller amplitude oscillations (e.g.
+Figure 5(e)(k)(q)), and (ii) more consistent trends across ACC, NMI, and SS metrics (Figure 5(c)(i)(o)).
+Furthermore, STPCA-MP’s T-SVDM-based framework effectively extracts tensor structural information, enabling more precise feature correlation modeling than: (a) CPD-based methods (CPUFS, CPUFSnn), (b) Tucker-based MSPCA, and (c) non-tensor methods (SPCAFS, RNE) on datasets with relatively strict feature spatial distributions (PIE, JAFFE, and Imm40).
+On datasets with irregular feature space distributions (COIL20 and USPS), STPCA-MP does not demonstrate a significant advantage in average performance, yet maintains top-ranked results across all tested feature selection dimensions (Figure 5(b)(h)(n)) and ultimately attains the optimal results (Table 7).
   4. 3.
-    STPCA-MP achieves the best performance at a smaller training time cost.
-On all datasets in Table 4, STPCA-MP takes less than $1$ second to complete training and feature selection, due to its slice-by-slice operation.
-Note that this is accomplished with reasonable prior knowledge of data organization.
-As a comparison, NLGMS, also showing remarkable average performance, brings much more computational cost as it overlooks the tensor structure by vectorization operation.
-Table 4: Metrics of the comparative methods on real-world datasets.
-The best ACC, best NMI, and mean POC regarding grid-search range are shown.
-The training time corresponding to the best result is recorded in seconds.
-The best result on each dataset is boldfaced, while the second best one is underlined and italic.
+    STPCA-MP combines high performance with computational efficiency.
+As shown in Table 7, STPCA-MP completes training and feature selection in under 1 second per dataset through its slice-by-slice operations.
+This efficiency is achieved while maintaining proper data structure awareness.
+In contrast, NLGMS and SPCAFS incur substantially higher computational costs due to their vectorization approach, despite showing competitive accuracy.
+[…]
 Method PIE COIL20 JAFFE BreastMNIST Imm40 USPS UCIDSA UCIHAR UMCS ASCH
@@
 NMI 51.19 NMI 75.74 NMI 4.51 NMI 2.41 NMI 74.21 NMI 61.11
+SS 0.0127 SS 0.2105 SS -0.0518 SS 0.0220 SS 0.0994 SS 0.1924
 RNE[53] ACC 18.67 ACC 57.83 ACC 26.96 ACC 60.26 ACC 58.54 ACC 55.51 POC TIME 40.00 83.57 POC TIME 0 170.22 POC TIME 40.00 193.22 POC TIME 66.67 205.33
 NMI 42.64 NMI 73.81 NMI 12.89 NMI 4.34 NMI 81.11 NMI 50.74
+SS -0.1828 SS 0.1304 SS -0.1284 SS 0.0279 SS 0.3168 SS 0.1211
 TIME 5.56 TIME 18.17 TIME 63.42 TIME 8.16 TIME 15.44 TIME 1.60
@@
 NMI 53.01 NMI 72.55 NMI 10.01 NMI 4.55 NMI 79.60 NMI 52.21
+SS 0.0263 SS 0.1661 SS -0.0366 SS 0.0456 SS 0.2411 SS 0.0916
 TIME 34.28 TIME 55.56 TIME 1794.41 TIME 37.12 TIME 83.77 TIME 391.60
@@
 NMI 65.04 NMI 74.02 NMI 12.07 NMI 5.14 NMI 84.09 NMI 61.13
+SS 0.2091 SS 0.2303 SS -0.0392 SS 0.0845 SS 0.3413 SS 0.2275
 TIME 18.37 TIME 3.11 TIME 59.36 TIME 0.61 TIME 15.76 TIME 0.34
@@
 NMI 62.69 NMI 76.74 NMI 18.34 NMI 4.29 NMI 82.03 NMI 61.45
+SS 0.1439 SS 0.2449 SS -0.0137 SS 0.0493 SS 0.3245 SS 0.2035
 TIME 25.99 TIME 25.71 TIME 498.12 TIME 2.98 TIME 19.08 TIME 511.68
@@
 NMI 56.56 NMI 75.48 NMI 13.42 NMI 3.48 NMI 81.41 NMI 59.79
+SS 0.0364 SS 0.2159 SS -0.0374 SS 0.0629 SS 0.2825 SS 0.1999
 TIME 290.69 TIME 69.89 TIME 11.20 TIME 6.87 TIME 14.01 TIME 1559.53
@@
 NMI 32.51 NMI 75.85 NMI 11.86 NMI 4.03 NMI 79.18 NMI 58.37
+SS 0.0279 SS 0.2380 SS -0.0563 SS 0.0644 SS 0.2150 SS 0.1879
 TIME 101.76 TIME 20.72 TIME 11.68 TIME 7.11 TIME 13.35 TIME 485.50
@@
 NMI 53.81 NMI 75.84 NMI 10.47 NMI 3.46 NMI 83.40 NMI 59.86
+SS 0.0284 SS 0.2244 SS -0.0784 SS 0.0442 SS 0.3139 SS 0.1853
 TIME 0.33 TIME 0.02 TIME 0.01 TIME 0.01 TIME 0.07 TIME 0.05
@@
 NMI 66.80 NMI 76.29 NMI 23.42 NMI 5.20 NMI 87.81 NMI 61.75
+SS 0.1812 SS 0.2494 SS -0.0244 SS 0.0811 SS 0.4400 SS 0.2141
 TIME 0.03 TIME 0.04 TIME 0.05 TIME 0.02 TIME 0.02 TIME 0.06
-Figure 4: The ACC and NMI curves on tube-wise datasets. The upper row shows the ACC curves, while the bottom row displays the NMI curves.
+Figure 5: The ACC, NMI, and SS curves on tube-wise datasets. The upper, middle, and bottom row displays the ACC, NMI, and SS curves, respectively.
 5.5 POC comparison experiment on slice-wise data
-For slice-wise datasets, we view each channel (row vector) as a feature.
-Therefore, we perform this experiment in the following steps: First, we implement a UFS algorithm and obtain the scores for each element.
-Second, we sum row-wise scores to evaluate features, using POC to measure the agreement between top-ranked features and those with maximal between-class variance (BCV).
-We compute BCV after the data are zero-centered.
-The BCV of selected features on four slice-wise datasets is shown in Figure 5, where the results corresponding to the greatest POC of typical comparative methods are given.
-The mean POC of all methods is illustrated in Table 4.
+For slice-wise datasets where each channel (row vector) is treated as a feature, the experiment proceeds through two key phases: (i) Feature Scoring: executing the UFS algorithm to generate element-wise importance scores, followed by (ii) Feature Evaluation: aggregating scores row-wise for channel-level rankings, computing BCV on zero-centered data, and computing POC by comparing top-ranked features with maximal BCV features.
+The number of selected features is set to $15$, $100$, $15$, and $6$ on UCIDSA, UCIHAR, UMCS, and ASCH, respectively.
+The BCV distribution of selected features across four slice-wise datasets is visualized in Figure 6, displaying results corresponding to each method’s maximal POC performance.
+Mean POC values are quantitatively compared in Table 7.
 We have the following observations:
@@
   2. 1.
-    For each dataset, there exists a small set of features with prominent BCV, judging by the peaks on the BCV curves in Figure 5. Greater BCV usually implies better inter-class discriminability.
+    Each dataset exhibits a subset of features with significantly higher BCV, as evidenced by distinct peaks in Figure 6(a)(g)(m)(s).
+These peaks correlate strongly with inter-class discriminative power.
   3. 2.
-    SPCA-based methods are generally more capable of selecting the features of the greatest BCVs.
-SPCAFS, MSPCA, and STPCA-MP have a stabler performance in terms of POC, owing to the encoder-decoder framework that minimizes the reconstruction error as well as maximizes the variance among samples on the global manifold.
-Attributed to the sparse tensor decomposition, MSPCA also achieves relatively stable performance on slice-wise data, referring to its mean POC in Table 4.
+    SPCA-based methods demonstrate superior BCV feature selection capability.
+SPCAFS, MSPCA, and STPCA-MP maintain stable POC performance through their shared encoder-decoder framework, which simultaneously minimizes reconstruction error and maximizes global sample variance.
+Notably, MSPCA’s sparse tensor decomposition contributes to its competitive performance on slice-wise data, as confirmed by its mean POC in Table 7.
   4. 3.
-    STPCA-MP has the advantage of handling slice-wise data over comparative methods.
-Due to the exploitation of data orientation, STPCA-MP can better learn the intrinsic relationship between different channels in a slice-wise tensor, selecting a considerable amount of BCV peaks in Figure 5.
-Although the best result of NLGMS is compatible to our method on some datasets, STPCA-MP demonstrates enhanced stability in selecting discriminative features, according to the mean POC regarding grid-search range in Table 4.
-On average, STPCA-MP is more effective toward slice-wise data, compared to the state-of-the-arts.
+    STPCA-MP outperforms competitors in slice-wise data processing.
+By explicitly modeling data orientation, STPCA-MP effectively captures cross-channel relationships within slice-wise tensors, consistently selecting BCV peaks, as shown in Figure 6(b)(h)(n)(t).
+While NLGMS achieves comparable peak performance on certain datasets, STPCA-MP shows superior stability in discriminative feature selection across the full parameter range, as illustrated in Table 7, establishing it as the most effective method for slice-wise data analysis.
 Figure 6: The BCV of selected features on slice-wise datasets.
 5.6 Visualization
-In this experiment, we perform a visualization to empirically compare the interpretability of selected features.
-The best results of three tensor-based methods and one non-tensor-based method in terms of NMI on two image datasets: JAFFE and BreastMNIST, are visualized in Figure 6, where we highlight the selected features to see the attention region of the UFS methods.
+In this experiment, we conduct feature visualization to empirically compare the interpretability of selected features.
+Figure 7 displays the best NMI results from three tensor-based methods (STPCA-MP, CPUFS, MSPCA) and one non-tensor method (NLGMS) on three image datasets: JAFFE, BreastMNIST, and PIE, with selected features highlighted to reveal method attention patterns.
 We have the following observations:
@@
   2. 1.
-    STPCA-MP pays more attention to discriminative features.
-The exploitation of data orientation enables STPCA-MP to mine essential information with different spatial distributions.
-On JAFFE, STPCA-MP selects more horizontal features that cover the regions of eyes, eyebrows, mouth, and hair, etc.
-beneficial for distinguishing facial expressions, while CPUFS and MSPCA are misguided by background pixels.
-NLGMS, though it selects reasonable features as well, has a more scattered region and fails to select correlated features.
-On BreastMNIST, the features selected STPCA-MP tend to be grouped at an abnormal shadowed region composed of both horizontally and vertically distributed features, while the results of CPUFS, MSPCA and NLGMS are diverse.
-We can also observe that compared to CPUFS, MSPCA selects more rows (or columns) containing discriminative features and is less affected by the background, which we believe is attributed to the direct utilization of sparse factor matrices (CPUFS doesn’t apply a sparse tensor decomposition).
+    STPCA-MP demonstrates superior discriminative feature selection.
+By leveraging data orientation, STPCA-MP effectively captures essential spatial patterns: (i) On PIE and JAFFE, it consistently selects horizontal features corresponding to facial components (eyes, eyebrows, mouth, hair) crucial for facial and expression recognition (Figure 7(b)(k)), while CPUFS, NLGMS and MSPCA misallocate attention to background pixels (Figure 7(c)(l)(d)(m)); (ii) On BreastMNIST, it concentrates on relevant shadowed regions with both horizontal and vertical features (Figure 7(f)), unlike the scattered selections of CPUFS, MSPCA, and NLGMS (Figure 7(g)(h)(I)).
+Notably, MSPCA’s sparse tensor decomposition provides moderate improvement over CPUFS by reducing background interference (Figure 7(d)(h)(m)).
   3. 2.
-    The score map of STPCA-MP exhibits significantly clearer and more distinct patterns than comparative methods.
-Thanks to the slice-by-slice operation, STPCA-MP describes feature correlations in a more delicate manner.
-On JAFFE, the score map of STPCA-MP approximates the shape of a human face.
-On BreastMNIST, the score map of STPCA-MP demonstrates a smooth score gradient transition radiating from the center of the shadowed region.
-In contrast, the score maps of CPUFS and MSPCA are less interpretable, due to the vague description of mode relations rooted in CPD and Tucker decomposition.
-As for NLGMS, its score map adds significant weights to the redundant features in the background.
-This phenomenon is possibly evoked by the sensitivity to noise features of the constructed graph.
-Figure 6: Visualization on JAFFE and BreastMNIST.
+    STPCA-MP produces more interpretable score maps.
+The slice-by-slice operation yields precise feature correlation modeling on: (i) Facial datasets, where STPCA-MP exhibits anatomically plausible attention maps resembling face contours (Figure 7(b)(k)); (ii) BreastMNIST, where STPCA-MP shows smooth score gradients radiating from pathological regions (Figure 7(f)).
+In contrast, CPUFS and MSPCA generate less coherent patterns due to inherent limitations of CPD/Tucker decompositions in modeling mode relations (Figure 7(c)(d)(g)(h)(l)(m)), while NLGMS overweight background noise (Figure 7(e)(i)), likely due to graph construction sensitivity.
+Figure 7: Visualization on PIE, JAFFE, and BreastMNIST.
 We show the score maps and highlight the positions of the top-ranked features in the samples.
-We set the number of selected features to $300$ on JAFFE, and $50$ on BreastMNIST.
+We set the number of selected features to $300$ on JAFFE, $50$ on BreastMNIST, and $100$ on PIE.
 5.7 Training time comparison
-In this experiment, we compare the computational cost of comparative methods.
-The training time corresponding to the best clustering results is given in Table 4, while the average training time in terms of regularization parameter combinations is illustrated in Figure 7.
+In this experiment, we analyze the computational efficiency of comparative methods.
+The training time corresponding to optimal clustering performance is reported in Table 7, with average training times across parameter combinations shown in Figure 8.
 We have the following observations:
@@
   2. 1.
-    Tensor-based methods are generally more efficient on high-dimensional data.
-By leveraging tensor-based techniques, the computational complexity is reduced from quadratic or cubic in the full feature space to quadratic or cubic within a single mode dimension.
-Therefore CPUFS, CPUFSnn, MSPCA, and our proposed method averagely cost less training time on JAFFE, UMCS, and ASCH compared to other methods.
+    Tensor-based methods demonstrate superior efficiency on high-dimensional data.
+By exploiting tensor operations, these methods reduce computational complexity from quadratic/cubic in full feature space to quadratic/cubic within single-mode dimensions.
+This explains why CPUFS, CPUFSnn, MSPCA, and our method consistently achieve lower training times on JAFFE, UMCS, and ASCH datasets compared to non-tensor alternatives (Figure 8(a)(b)(c)).
   3. 2.
-    STPCA-MP is efficient on both high-dimensional and large-scale data.
-CPUFS and CPUFSnn usually take hundreds of iterations to converge because of the non-convexity.
-Moreover, their computational complexity scales quadratically with #$\#(Sample)$, whereas STPCA-MP exhibits linear complexity relative to #$\#(Sample)$.
-Therefore, STPCA-MP takes considerably less training time than CPUFS and CPUFSnn on all datasets, especially on large-scale datasets.
-Although MSPCA is on average more efficient than STPCA-MP due to a smaller weight parameter size requiring computation in Tucker decomposition, it gains no advantage over the proposed method in both clustering tasks and the POC metric.
-Moreover, the training time associated with optimal performance shows only a marginal discrepancy between MSPCA and STPCA-MP (See Table 4), which is in an acceptable range.
-On the whole, the proposed method is both effective and efficient.
+    STPCA-MP maintains efficiency across both high-dimensional and large-scale datasets.While CPUFS and CPUFSnn require hundreds of iterations (due to non-convexity) with quadratic sample complexity $\mathcal{O}(n^{2})$, STPCA-MP achieves linear complexity $\mathcal{O}(n)$.
+This efficiency advantage becomes particularly pronounced on large-scale datasets like USPS (Figure 8(d)).
+Although MSPCA’s Tucker decomposition with smaller weight parameters yields marginally faster average training times, it fails to outperform STPCA-MP in either clustering metrics or POC (Table 7).
+Notably, the runtime difference between MSPCA and our method at optimal performance is negligible (Table 7), confirming STPCA-MP’s balanced effectiveness-efficiency trade-off.
 (a) JAFFE (#Feature=$4096$, #Sample$=213$)
@@
 5.8 Convergence analysis
-In this experiment, we study the convergence of the proposed method.
-Figure 8 illustrates the converge curves of objective function value on four real-world datasets.
-Utilizing the HPSD projector, the proposed two-step algorithm converges fast into the solution space within dozens of iterations.
+In this experiment, we investigate the convergence properties of the proposed method through a representative subproblem analysis due to space constraints.
+Figure 9 demonstrates the convergence behavior by plotting both the objective function values and runtime per iteration across four real-world datasets.
+The two-step algorithm achieves rapid convergence to the solution space within dozens of iterations, facilitated by the HPSD projector that ensures stable optimization.
+The slice-by-slice operation significantly reduces computational complexity, enabling efficient execution of each iteration while maintaining solution quality.
 (a) PIE
@@
 (d) UCIDSA
-Figure 8: Converge curves of STPCA-MP on JAFFE and BreastMNIST. All regularization parameters are set to 1.
+Figure 9: Convergence curves and runtime per iteration of STPCA-MP on four real-world datasets. All regularization parameters are set to 10.
 5.9 Parameter sensitivity analysis
-In this section, we analyze the impact of the parameters $\lambda$ and $\eta$ in STPCA-MP on the performance of feature selection.
-We vary the parameters in the range of $\{10^{-2},10^{-1},1,10^{1},10^{2}\}$, set the selected features to 100, and record the ACC and NMI of STPCA-MP with all possible parameter combinations.
-The results on JAFFE and USPS is illustrated in Figure 9.
-We can see that
+In this section, we evaluate the impact of regularization parameters $\lambda$ and $\eta$ on the feature selection performance of STPCA-MP.
+By varying both parameters within $\{10^{-2},10^{-1},1,10^{1},10^{2}\}$ while fixing the selected features at 100, we record the ACC and NMI values across all parameter combinations.
+Figure 10 presents the results on JAFFE and USPS datasets, revealing two key findings:
   1. [•]
   2. 1.
-    On JAFFE, the performance of STPCA-MP is relatively stable when $\lambda$ and $\eta$ vary in a certain range, and degenerates when the parameters become too large.
-This tendency is consistent between ACC and NMI.
+    For JAFFE, STPCA-MP maintains stable performance when $\lambda$ and $\eta$ remain within a moderate range, with noticeable degradation occurring only at excessively large parameter values - a trend consistently observed in both ACC and NMI metrics.
   3. 2.
-    On USPS, STPCA-MP is generally not sensitive to the choice of parameters, achieving relatively stable performance with respect to both ACC and NMI.
+    For USPS, the method demonstrates robust performance across the tested parameter range, showing minimal sensitivity to $\lambda$ and $\eta$ variations in terms of both ACC and NMI measurements.
 (a) JAFFE (ACC)
@@
 Figure 10: Parameter sensitivity analysis of STPCA-MP on JAFFE and USPS.
-5.10 Discussion on different choices on data orientation and transform domain.
-In this experiment, we discuss the influence of data tensor orientation and transform domain.
-We conduct STPCA-MP with different data tensor rotations and transform matrices $\mathbf{M}$ on several datasets.
-The detailed ACC curves on JAFFE and USPS are illustrated in Figure 10.
-We further provide the best ACC and the mean POC in Table 5.
-The visualized result with $\mathbf{M}=\mathbf{I}$ on PIE and JAFFE is given in Figure 11.
+5.10 Impact of Data Orientation and Transform Domain
+In this experiment, investigate how data tensor orientation and transform matrix $\mathbf{M}$ affect STPCA-MP performance.
+Experiments evaluate different tensor rotations and transform matrices across multiple datasets, with detailed ACC curves for JAFFE and USPS shown in Figure 11.
+Table 8 reports the optimal ACC and mean POC values, while Figure 12 visualizes results using $\mathbf{M}=\mathbf{I}$ on PIE and JAFFE.
 We have the following observations:
@@
   2. 1.
-    Different data orientations can significantly affect STPCA-MP’s clustering performance.
-We can see significant gaps among certain ACC curves in Figure 10.
-Like looking at an identical object from different views, decomposing a data tensor in different directions naturally leads to discrepant information.
-The visualized result in Figure 11 implies that STPCA-MP can capture features with different spatial distributions given different data orientations.
-A proper choice of data orientation is beneficial for mining feature correlations.
-The spatial distribution and latent correlation structure of discriminative features are heterogeneous among data tensors.
-On UCIDSA and UMCS, the overall performance of the Dir1s is better than the Dir2s, since the feature relations in multi-channel time-series data exist along 1-mode.
-In conclusion, performing slice-by-slice operations along a certain direction benefits UFS at multiple levels on different datasets.
+    Data orientation critically influences clustering performance.
+Significant ACC variations in Figure 11(a) and (b) demonstrate that, similar to viewing objects from different perspectives, tensor decomposition along distinct directions captures varying information.
+Figure 12 confirms STPCA-MP’s ability to identify orientation-dependent spatial feature distributions.
+Optimal orientation selection enhances feature correlation mining, particularly for multi-channel time-series data like UCIDSA and UMCS, where 1-mode operations outperform due to inherent feature relationships along this dimension.
   3. 2.
-    A suitable transform domain can benefit the performance.
-Since the transform extracts the information from one of the two feature modes, while a slice-by-slice operation is conducted along the other feature mode, a proper transform domain can integrate more feature correlations as prior knowledge for the following optimization and UFS.
-In Table 5, Dir1-EigVector sometimes achieves better results than Dir-I, especially on time-series data.
-However, there is no absolute conclusion that certain settings of transform domain necessarily outperform others.
-As we can observe in both Figure 10 and Table 5, there is no setting under which the metrics consistently outperform all other configurations.
-Identical settings of orientation and transform domain can lead to substantial performance discrepancies across different datasets.
-In general, the choice of orientation and transform domains requires detailed configurations to achieve the best performance.
+    Transform domain selection impacts results.
+A well-selected transform matrix extracts information from one feature mode while slice-by-slice operations process the other, enabling prior knowledge incorporation for optimization.
+Table 8 shows Dir1-EigVector occasionally surpassing Dir-I, especially for time-series data.
+However, no transform configuration universally dominates - both Figure 11(a)(b)(c)(d) and Table 8 demonstrate performance varies substantially across datasets.
+Therefore, optimal orientation and transform choices require dataset-specific tuning.
 (a) JAFFE (different orientations)
@@
 An efficient slice-by-slice algorithm to conduct STPCA-MP is proposed to guarantee fast convergence.
-The experimental results on real-world data show that our method outperforms state-of-the-art methods on average and can handle different data organization scenarios at a smaller training time cost.
+The experimental results on real-world data show that our method outperforms state-of-the-art methods on average in terms of clustering metrics (ACC, NMI, and SS) in different data organization scenarios at a smaller training time cost.
 In particular, STPCA-MP has the advantage of selecting discriminative features with the greatest BCV on slice-wise data.
 In addition, we also test the parameter sensitivity, examine the efficient convergence, and discuss the impacts of choosing different data orientations and transform domains.
-The experimental results show that with a suitable combination of data orientation and transform domain, our method can capture abundant discriminative information and achieve remarkable performance.
-There are two valuable research points concerning STPCA-MP in the future.
-First, the best choice of transform domain is possibly related to the background knowledge of data.
+The experimental results show that when the transform axes align with the feature distribution pattern, our method can capture abundant discriminative information and achieve remarkable performance.
+There are two valuable research points concerning STPCA-MP for the future.
+First, the best choice of transform domain is possibly related to the data background knowledge.
 More experiments should be carried out for further applications.
@@
 We will be working on addressing the above two problems in the future.
-Appendix A Appendix
+Appendix A Proof of Theorem 1
 […]
-pf3Proof of Theorem 1
+pf3Proof
 {pf3}
@@
 $s.t.$ $\mathbf{U}^{H}\mathbf{U}=\mathbf{I}_{k}.$
-Thus, we can rewrite the objection function as:
+Thus, we can rewrite the objective function as:
 ${\rm Tr}\left(\mathbf{X}\mathbf{X}^{H}\right)-\sum\limits^{k}_{j=1}\bigg{[}2Tr\left(\mathbf{u}_{j}^{H}\mathbf{X}\mathbf{X}^{H}\mathbf{q}_{j}\right)-\frac{\lambda}{\|\mathbf{q}_{j}\|_{2}}\mathbf{q}_{j}^{H}\mathbf{q}_{j}-{\rm Tr}\left(\mathbf{q}_{j}^{H}\mathbf{X}\mathbf{X}^{H}\mathbf{q}_{j}\right)\bigg{]}$ (48)
@@
 Eventually, after each subproblem of (16) is solved, the optimal solution satisfies $\hat{\mathcal{A}}^{opt}_{::i}=\hat{\mathcal{U}}^{opt}_{::i}(\hat{\mathcal{Q}}^{opt}_{::i})^{H}\in H^{q}_{+}$.
+Appendix B Reproducibility Statement
+To facilitate the reproducibility of our research, we provide the following resources:
+Code Implementation
+  • Code Repository: Available at https://github.com/zjj20212035/STPCA.git under the main branch (git checkout v1.1).
+  • Dependencies:
+    • MATLAB R2023b
+    • Tensor Toolbox v3.5 (included in utils/)
+  • Parallel Computing: All experiments utilize MATLAB’s Parallel Computing Toolbox.
+Datasets
+  • All datasets are provided in data/ folder as .mat files, each is unified to contain the following common attributes:
+    • data: $d_{1}\times d_{2}\times n$ tensor ($n$=samples, $d_{i}$=$i$-mode dimensionality)
+    • label: $n\times 1$ class vector
+    • tensor_type: tube-wise or slice-wise
+    • tensor_size: dimensionality of the feature space
+  • No manual data preprocessing needed - all preprocessing is handled automatically by:
+    • config_Table7.m
+    • config_Figure7.m
+    where each dataset shares the same normalization that is carried out by dividing each element by the maximum absolute value, scaling all values to the interval $[-1,1]$.
+Experiment Reproduction
+We have streamlined the code to allow one-click reproduction of the paper’s key results. Open the project in MATLAB and
+  • To replicate Table 7:
+        cd code
+        run(’exp_Table7_clustering_POC.m’)
+  • To replicate Figure 7:
+        cd code
+        run(’exp_Figure7_visualization.m’)
+Hyperparameters are specified in config_Table7.m and config_Figure7.m, where:
+  • Regularization parameters $\lambda$ and $\eta$ are tuned by grid search in the range of $\{10^{-2},10^{-1},1,10,10^{2}\}$,
+  • Orientation is tuned from $D^{o}=\{\{1,3,2\},\{2,3,1\}\}$ and the best result is shown,
+  • The stabilization constants are fixed at $\epsilon_{1}=10^{-3}$ and $\epsilon_{2}=10^{-2}$ throughout all experiments.
 […]